### v1

# True Zero-Shot Inference of Dynamical Systems Preserving Long-Term Statistics

Abstract
Complex, temporally evolving phenomena, from climate to brain activity, are governed by dynamical systems (DS). DS reconstruction (DSR) seeks to infer generative surrogate models of these from observed data, reproducing their long-term behavior. Existing DSR approaches require purpose-training for any new system observed, lacking the zero-shot and in-context inference capabilities known from LLMs. Here we introduce DynaMix, a novel multivariate ALRNN-based mixture-of-experts architecture pre-trained for DSR, the first DSR model able to generalize zero-shot to out-of-domain DS. Just from a provided context signal, without any re-training, DynaMix faithfully forecasts the long-term evolution of novel DS where existing time series (TS) foundation models, like Chronos, fail – at a fraction of the number of parameters and orders of magnitude faster inference times. DynaMix outperforms TS foundation models in terms of long-term statistics, and often also short-term forecasts, even on real-world time series, like traffic or weather data, typically used for training and evaluating TS models, but not at all part of DynaMix’ training corpus. We illustrate some of the failure modes of TS models for DSR problems, and conclude that models built on DS principles may bear a huge potential also for advancing the TS prediction field.
1 Introduction
Figure 1: DynaMix achieves zero-shot DSR of attractor geometry and long-term temporal properties (power spectrum) from a short context signal while Chronos [2] fails.
Most real-world processes, from atmospheric phenomena and stock markets to brain activity or ecological networks, can be described as dynamical systems (DS) [54, 15, 82, 57]. Reconstructing these from observational data, called dynamical systems reconstruction (DSR), has been a long-standing challenge in scientific modeling [92, 23, 31]. DSR goes beyond conventional time series (TS) modeling, as we wish to have a generative model of the underlying process which exhibits long-term behavior with the same temporal and geometrical signatures (Fig. 1), i.e. the same invariant or ‘climate’ statistics [62, 66], as important in scientific applications. Achieving this usually requires special control-theoretic training techniques [56, 38] or loss objectives [65, 66, 40, 72] which accentuate the system’s long-term dynamics. Numerous deep learning approaches for DSR, based on recurrent neural networks (RNNs; [81, 22, 83, 16, 6, 69, 38, 7, 66]), neural ODEs [17, 43, 1, 45], Koopman operators [14, 53, 60, 3, 58, 28, 86], or library-based methods [12, 52, 41, 19, 55], have been advanced over the years. However, all of these require purpose-training on the specific system observed and struggle to generalize beyond their training distribution [33].
Inspired by the strong in-context and zero-shot generalization abilities of LLMs [10, 27, 21, 18], there has been a push recently to develop models with likewise properties for the time series domain. Time series foundation models like Chronos [2] or Mamba4Cast [5] are pre-trained on a huge database of real-world and artificial time series, and then tasked to forecast novel time series from which snippets are presented in-context, without any parameter fine-tuning. They are not built for DSR, however, and – as we show here – typically fail to properly capture a system’s long-term behavior and the structure of its attractors (Fig. 1; [93]).
To address this gap, we introduce DynaMix, a novel mixture-of-experts model designed for DSR and zero-shot forecasting. By training across a diverse range of DS, DynaMix learns transferable representations, enabling it to forecast previously unseen systems without retraining, including their underlying attractor geometries and invariant statistics (Fig. 1). The core features of DynaMix are:
  • Accurate zero-shot DSR: DynaMix achieves strong generalization across diverse DS, eliminating the need for extensive fine-tuning while maintaining accuracy in reproducing attractor geometry and long-term statistics.
  • Multivariate information transfer: Due to its multivariate architecture, the model efficiently captures dependencies among multiple system dimensions, enabling accurate reconstruction of their coupled dynamics. Further, it is neither bound to a specific dimensionality, but can flexibly adapt to other dimensions through specific embeddings, nor to a specific context length.
  • Computational and parameter efficiency: The model achieves high performance with a very lightweight architecture, enabling orders of magnitude faster inference than other foundation models.
  • Interpretable dynamics composition: The model provides insight into the dynamical composition of reconstructed systems, elucidating similarities between different DS.
  • Time series forecasting: Beyond DSR, our model is applicable to general time series forecasting where we employ different embedding techniques to obtain data representations accessible to DSR.
We evaluate our model on multiple benchmark DS and real-world time series, demonstrating superior zero-shot generalization on DSR problems compared to current TS foundation models.
2 Related work
Dynamical systems reconstruction (DSR)
In DSR we seek to learn generative models from time series data that represent the underlying system dynamics with all its topological, geometrical, and temporal properties [23, 33, 66, 30, 13, 11], in that sense providing an approximation to the underlying system’s governing equations. By definition, such a model should not only render viable short-term forecasts, but also reproduce the long-term evolution of the DS it has been trained on, both in state space and in the time domain. Methods approaching this goal have been founded on predefined function libraries, such as Sparse Identification of Nonlinear Dynamics (SINDy; [12, 52, 19, 41, 55]), on reservoir computers [63, 65, 66], neural ODEs [17, 43, 1, 45], Koopman operators [14, 53, 60, 3, 58, 28, 86], or different types of RNNs [81, 22, 83, 16, 6, 69, 38, 7]. More important than the architecture itself seems to be their proper training in order to ensure long-term (invariant) statistics of the underlying system are met: Methods like sparse [56, 6] or generalized [38] teacher forcing allow RNN-generated trajectories to ‘explore the future’ whilst training, yet keep loss gradients in check. Alternatively, long-term statistics based on the observed system’s Lyapunov spectrum, fractal geometry, or invariant measures may be added to regularize the loss [65, 66, 40, 72], but require to compute these properties from the data first. Despite these advances, out-of-domain generalization remains a key challenge in DSR [33]. Meta- and hierarchical learning models, trained across many DS simultaneously, have recently been designed as steps toward foundation models for DSR [9, 59, 44, 91], but still require parameter fine-tuning and lack in-context inference capabilities.
Time series foundation models
Large language models (LLMs) exhibit an impressive ability to infer patterns from prompts (context) and generalize to novel situations without retraining [10, 27, 21, 18], although the underlying reasons for this are still a matter of debate [49, 85, 90]. This inspired the development of general-purpose time series (TS) foundation models which could accurately forecast time series from a short segment provided ‘in-context’ [2, 26], without the need of task-specific fine-tuning [20]. One idea is to simply use pretrained LLMs directly as TS forecasters [34, 67, 75, 88, 79], but the inherent differences between textual and many temporal data, such as continuity, present significant challenges [78]. To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos [2] or Moirai [51], were specifically pretrained on a large corpus of time series data. Promising zero-shot forecasting capabilities have also been achieved with alternative architectural designs, such as Tiny Time Mixers [26] or state-space models like Mamba [5].
The success of TS foundation models raised hope these could also be utilized for zero-shot DSR, but the – to our knowledge – so far only previous study on this, based on Chronos, had mixed outcomes and fell short of a full DSR evaluation [93]: Successful DSR, if present at all, heavily depended on the initial conditions, and mechanisms such as context parroting visually gave the wrong illusion that features of the dynamics had been captured. In fact, here we show that existing TS foundation models are generally not capable of producing valid DSRs (Fig. 1), and highlight some of their failure modes. This is in contrast to the zero-shot DSR foundation architecture we develop here, based on models [7] and training algorithms [56] successful in DSR.
3 Methods
3.1 Model architecture
To enable zero-shot reconstruction of novel DS, here we develop a specific mixture-of-experts (MoE) architecture that can be pretrained across many diverse DS (Fig. 9), with different experts possibly specializing in different dynamical regimes. As a base model for the experts, we leverage a recent parameter-friendly RNN designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; [7]):
$\bm{z}_{t}=\bm{A}\bm{z}_{t-1}+\bm{W}\Phi^{*}(\bm{z}_{t-1})+\bm{h}\;.$ (1)
The model describes the evolution of an $M$-dimensional latent process $\bm{z}_{t}\in\mathbb{R}^{M}$, with linear self-connections $\bm{A}\in\text{diag}(\mathbb{R}^{M})$, weight matrix $\bm{W}\in\mathbb{R}^{M\times M}$, bias term $\bm{h}\in\mathbb{R}^{M}$, and $\Phi^{*}(\bm{z}_{t})$ defined as
$\Phi^{*}(\bm{z}_{t}):=\left[z_{1,t},\cdots,z_{M-P,t},\max(0,z_{M-P+1,t}),\cdots,\max(0,z_{M,t})\right]^{T},$ (2)
i.e. with a ReLU nonlinearity on only $P<<M$ out of the $M$ AL-RNN units. The first $N$ units are interpreted as the network’s readouts and provide the predicted observations, $\bm{\hat{X}}=\{\bm{\hat{x}}_{t}=\bm{z}_{1:N,t}\}\in\mathbb{R}^{N\times T}$, where $N$ is the observation dimension.
The selection of AL-RNN experts is achieved through a gating network (Fig. 2), which receives as inputs the (generally multivariate) context time series $\bm{C}=\{\bm{c}_{t}\}\in\mathbb{R}^{N\times T}$ as well as the current latent state $\bm{z}_{t}$. Both are passed into a state attention mechanism defined as
$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T}^{\top}\right|^{\top}\bm{1}_{N}}{T_{\text{att}}}\right)\in\mathbb{R}^{T}\;,$ (3)
where $\bm{D}\in\mathbb{R}^{N\times M}$ is a learnable matrix which maps the latent state into observation space, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{\Sigma})$ exploration noise with a learnable covariance $\bm{\Sigma}$, $T_{\text{att}}$ a learnable temperature parameter, $\bm{1}_{\{T,N\}}$ column vectors of ones of length $T,N$, respectively, and $\sigma(\cdot)$ the softmax returning normalized weights. Hence, this mechanism computes attention weights based on some distance between projected latent states and actual context observations. At the same time, the context signal $\bm{C}$ is processed by a CNN, yielding temporal features $\tilde{\bm{C}}\in\mathbb{R}^{N\times T}$. These features are then weighted with the attention weights $\bm{w}^{att}_{t}$, and together with the current state $\bm{z}_{t}$ further processed by a multi-layer perceptron (MLP) followed by a softmax, yielding a set of expert weights at any time $t$:
$\bm{w}^{exp}_{t}=\sigma\left(\frac{\text{MLP}(\tilde{\bm{C}}{\bm{w}^{att}_{t}},\bm{z}_{t})}{T_{\text{exp}}}\right)\in\mathbb{R}^{J}\;,$ (4)
where $T_{\text{exp}}$ is another learnable temperature parameter. The forward-iterated latent states $\bm{z}_{t+1}^{j}$ of the individual experts $j\in\{1,...,J\}$ are then weighted by $\bm{w}^{exp}_{t}$ to yield the next time step prediction $\bm{z}_{t+1}=\sum_{j=1}^{J}w^{exp}_{j,t}\cdot\bm{z}_{t+1}^{j}$. Fig. 2 illustrates the whole approach. Note that one key advantage of this compared to other TS foundation architectures is that the context length is flexible by design, i.e. the CNN and the attention mechanism may take as input a context signal of arbitrary length.
Figure 2: Illustration of the DynaMix architecture. At test time, only a context signal $\bm{C}$ is provided and guides the selection of experts to yield arbitrarily long forward predictions of the dynamics.
3.2 Model training
For training, a collection of just $34$ different chaotic and cyclic DS is used [29], see Appx. A.2 for full specification. From each, multivariate time series $\bm{X}\in\mathbb{R}^{N\times T}$ are simulated ($\approx 1.6$ million in total) and then standardized (dimension-wise), of which the first $T_{C}<T$ column entries are defined as the context signal $\bm{C}=\bm{X}_{1:T_{C}}$. The experts are initialized at $t_{0}=T_{C}-\Delta t+1$ and forward-iterated until time $T$, such that they overlap for a period $\Delta t$ with the context. This is to ensure that the model learns to optimally utilize the context for generalization. A MSE loss is then computed across ground truth, $\bm{X}_{T_{C}-\Delta t+1:T}$, and respective model-generated time series $\hat{\bm{X}}$. The model is trained by sparse teacher forcing (STF; [6, 56]), a control-theoretic technique specifically designed for DSR. STF replaces part of the forward-iterated latent states $\bm{z}_{t}$ by data-inferred states $\hat{\bm{z}}_{t}$ at optimally chosen intervals $\tau$, to avoid exploding gradients even for chaotic systems while enabling the DSR model “to explore the future” (see [56] for detailed theoretical motivation). As shown in Fig. 24, STF with optimal $\tau$ is essential for achieving good zero-shot DSR results. Note that STF is only used for training and turned off at test time.
3.3 DSR evaluation & time series forecasting
We used two established measures to assess DSR quality [46, 6, 38, 89, 93, 66, 56, 61, 30]: First, to quantify agreement in state space (attractor) geometry, we employed a Kullback-Leibler divergence defined across space, $D_{stsp}$ (see Appx. A.1 for details). For assessing agreement in long-term temporal properties, we used the Hellinger distance $D_{H}$ defined across power spectra of the true and model-generated trajectories. In addition, a short-term, $n$-step-ahead mean absolute prediction error was evaluated. As has been emphasized repeatedly in the statistical and DSR literature [89, 46, 6, 56, 64, 30], prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig. 22).
Empirically observed time series are often just one-dimensional ($N^{*}=1$), but come from an inherently much higher-dimensional underlying DS. Our approach can in principle handle context signals of dimension $N^{*}<N$ smaller than those used in training by, e.g., just filling up $N-N^{*}$ dimensions in $\bm{C}$ with zeros. However, to work efficiently in a DSR setting, the context signal should represent the underlying DS as good as possible, for which it commonly needs to be lifted into a higher-dimensional space where sets explored by the dynamics ideally become diffeomorphic to those in the system’s true state space. Temporal delay embedding is the most popular technique to achieve this [76, 70]. For a $1d$ time series $\{x_{t}\}$, a $d$-dimensional embedding can be defined as [47]
$\bm{x}^{emb}_{t}=(x_{t},x_{t-\tau_{1}},...,x_{t-\tau_{d-1}})\;,$ (5)
where the $\tau_{i}$ are time lags, commonly estimated from the autocorrelation of the time series. Alternatively, if the delay embedding would become too large ($d>N$), we augment the empirically observed time series by a variant of the positional encoding common in transformers, defined here as
$\bm{x}^{emb}_{t}=\left(x_{t},\sin{\left(\frac{2\pi t}{\tau}+\phi_{1}\right)},...,\sin{\left(\frac{2\pi t}{\tau}+\phi_{N-1}\right)}\right)$ (6)
where $\tau:=\operatorname*{arg\,max}_{\tau>\tau_{min}}\mathbb{E}[x_{t}x_{t+\tau}]$ is given – provided it exceeds a threshold $\tau_{min}$ – by the maximal autocorrelation, and a random phase $\phi_{i}\in\left[0,\frac{\pi}{2}\right]$ is assigned to each dimension.
4 Results
4.1 Zero-shot DSR
Figure 3: a) DynaMix zero-shot DSR (red) compared to ground truth (blue) using a $500$-step context (gray) for the Sprott M system. b) Zero-shot forecasts (red) for the Selkov DS (true vector field in blue) from different initial conditions outside the context range (gray). c) DSR quality as a function of context length. d) DSR quality as a function of the temporal resolution $\Delta t$ of the context signal.
Figs. 1 (top) & 3a present example zero-shot DSRs of new chaotic DS, which were not part of DynaMix’ training domain, from just a provided context signal $\bm{C}$ (see Figs. 10-14 for further examples). Despite the inherent challenge of predicting chaotic systems, where small differences in initial conditions lead to rapidly diverging trajectories (Fig. 22), DynaMix successfully captures the system’s temporal & geometrical behavior, i.e. accurately reconstructs the underlying state space and faithfully recovers the true system’s power spectrum. Fig. 3b highlights further aspects of DynaMix’ reconstructions: Even though not trained on any $2d$ systems, the model successfully recreates the limit cycle from the $2d$ Selkov system [73]. Notably, the model even generalizes to new initial conditions outside the scope of the context data, i.e. correctly infers properties of the state space beyond the short context trajectory. As further shown in Fig. 3c, the context length necessary to achieve good DSR quickly converges, with as few as $500$ time steps often sufficient. Fig. 3d illustrates that our approach also works across a surprisingly wide range (about an order of magnitude) of temporal resolutions at which the underlying DS was sampled, confirming its robustness w.r.t. sampling frequency.
We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos [2], Mamba4Cast [5], and Tiny Time Mixers [26], using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see sect. A.1). As seen in Fig. 4a-b, DynaMix clearly and significantly outperforms all of these. In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the (strongest in our batch) Chronos-t5 models (t-tests, all $p>0.37$), as shown in Fig. 4c. Table LABEL:tab:performance_DSR assembles further results on performance comparisons, while Figs. 15-20 provide specific examples. Moreover, a significant limitation of particularly the large Chronos models with millions of parameters is their extreme inefficiency in both inference time and computational costs, as compared in Fig. 5. In fact, their inference time can even exceed that of a custom-trained model. In contrast, DynaMix provides a highly efficient alternative, offering superior DSR performance while using much less parameters and computation time.
Figure 4: Zero-shot DSR performance across all 54 test set DS for DynaMix and various TS foundation models for context length $T_{C}=2000$ (see Fig. 21 for results with $T_{C}=512$). Median±$\pm$±MAD of a) $D_{stsp}$ (geometrical disagreement), b) $D_{H}$ (temporal disagreement), and c) short-term prediction error.
Figure 5: Inference times for zero-shot generation of Lorenz-63 DS with $10000$ forecasting steps. All models were provided the exact same context and run on the same hardware. Note the log-scale on the $x$-axis. For comparison, also the time needed for training and inference of a custom-trained AL-RNN is shown in red.
4.2 Reasons for the failure of TS foundation models on DSR
Why do TS foundation models perform so poorly on DSR problems? Essentially, they face three key issues: First, most TS foundation models cannot efficiently deal with multivariate time series, but treat dimensions independently. However, in a nonlinear DS all variables are usually coupled, and their joint evolution is actually defining for the system’s (long-term) dynamics and attractor states. Thus, ignoring this, as illustrated in Fig. 6a, results in poor long-term accuracy (see also Figs. 15-20). This highlights a fundamental limitation: Univariate approaches are inherently insufficient for reconstructing multivariate, interacting DS. Second, TS foundation models are trained for short-term prediction, but not for DSR, providing another reason why they fail to reconstruct the correct long-term behavior. As illustrated in Figs. 1 & 6b, in the longer run they often converge to fixed points or simple cycles where the true dynamics is chaotic, not reflecting the context anymore. Third, context parroting, a phenomenon originally identified in LLMs [4], has also been observed in TS foundation models [93], where they tend to repetitively reproduce the exact input pattern rather than dynamically adapting to the system’s true evolution. Thus, TS foundation models simply produce oscillations, where the true dynamics is chaotic and therefore by definition irregular and non-repeating (Fig. 6c). This is illustrated in more detail in Fig. 6d for Chronos-t5-base, the best of the TS foundation models in our batch. It is particularly evident in the power spectrum (Fig. 6d right), where Chronos’ sharp, narrow peaks wrongly suggest periodicity, while the true spectrum is usually rather broad and smeared out for chaotic systems, as the one exhibited by DynaMix. We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent [56], a measure for the exponential divergence rate of trajectories (see Appx. A.1). For Chronos, the value $\lambda_{Chronos}\approx 0.06$ close to zero confirms the nearly periodic behavior, while DynaMix produces exponentially diverging trajectories with $\lambda_{DynaMix}\approx 1.48$ close to the ground truth value of $\lambda_{Lorenz63}\approx 1.37$. Hence, unlike DynaMix, Chronos does not reconstruct the underlying dynamics but just repeats the context. On the other hand, if the underlying dynamics truly is cyclic, DynaMix settles into a cyclic pattern as well, exhibiting no intrinsic bias, as illustrated in Fig. 6e.
Figure 6: Common problems in TS foundation models (red: forecast, gray: context, blue: ground truth). a) Dimensions are dynamically decoupled. b) Long-term forecasts often converge to simple fixed points or cycles unrelated to the true dynamics. c) Context parroting covers up the true dynamics. d) Example forecasts for the chaotic Lorenz-63 system with DynaMix and Chronos-t5-base. Chronos’ context parroting results in cyclic repetition of a fixed pattern, thus inherently missing the true aperiodic, chaotic dynamics. This is evident in the close-to-0 max. Lyapunov exponent (center) and peaked power spectrum (right). e) For comparison, if the underlying behavior is truly cyclic with prominent peaks in the power spectrum, both Chronos and DynaMix are able to capture it.
4.3 Dynamical similarity
DynaMix’ mixture-of-experts design produces a natural similarity measure for comparing the dynamics of different systems, by evaluating the relative contribution of each of the experts to the forecast dynamics. As shown in Fig. 7a (bottom), different experts specialize in specific dynamical patterns or regimes. Hence, we can utilize the average expert usage $\bm{e}=\frac{1}{T}\sum_{t=1}^{T}\bm{w}^{exp}_{t}$ across the generated time series to construct a similarity metric
$S=\frac{1}{S_{max}}\left(||\bm{e}_{1}-\bm{e}_{2}||_{1}\right)^{-1}\;,$ (7)
defined here as the inverse $L_{1}$ norm of the $\bm{e}$ vectors between any two systems, rescaled to the range $[0,1]$.
Figure 7: a) Top: Vector fields of different experts which in combination determine the attractor. Bottom plot illustrates the dominant expert in each time window. b) Similarity matrix among all $84$ types of attractor dynamics in our dataset, determined by computing $S$ from the reconstructed time series. c) Examples of two pairs of attractors classified as similar (top), and one dissimilar pair.
4.4 Time series forecasting from a DS perspective
Figure 8: Time series forecasts for different data domains (gray: context, blue: ground truth, other colors: various model predictions): Forecasts of a) partially ($1$d) observed Lorenz-63 DS, b) electricity transformer temperature data, c) hourly car traffic data with weekly cycle, d) Huawei cloud request data, e) soil temperature development, f) air pressure data, g) human functional magnetic resonance imaging (fMRI) data, h) human electroencephalogram (EEG) data.
Table 1: Performance comparison on empirical time series in terms of geometrical divergence ($D_{\text{stsp}}$), long-term temporal distance ($D_{H}$), and forecast error (MAE). Best in red, second-best in blue.
System DynaMix Chronos-t5-base Mamba4Cast TTM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Partially obs. DS 0.11 0.25 0.32 0.11 0.32 0.39 8.14 0.35 0.76 6.54 0.46 0.89
ETTh1 0.8 0.08 5.97 0.33 0.09 3.22 0.94 0.1 3.63 6.15 0.09 4.64
Traffic 0.68 0.21 8.42 0.35 0.1 3.16 1.4 0.34 7.68 1.77 0.25 5.25
Cloud Requests 0.17 0.14 147.28 1.57 0.3 557.35 7.91 0.16 282.95 1.39 0.26 147.54
Weather (Temp.) 0.59 0.09 2.24 4.72 0.22 4.14 6.96 0.58 9.39 5.39 0.3 2.76
Weather (Press.) 0.41 0.19 8.85 6.62 0.23 9.05 7.03 0.52 8.6 7.27 0.25 10.03
Human fMRI 0.63 0.09 0.56 5.77 0.12 0.85 7.99 0.11 1.35 6.66 0.21 0.9
Human EEG 0.78 0.23 1.09 10.17 0.24 1.1 6.0 0.62 1.12 2.48 0.37 0.92
We next tested and compared DynaMix on a variety of real-world datasets often used to probe time series models. As discussed in sect. 3.3, for DynaMix we used a multivariate embedding of the typically $1d$ empirical time series, to better reflect the underlying DS properties. Fig. 8a confirms for a $1d$ time series from the $3d$ chaotic Lorenz-63 system that this works well and DynaMix, in contrast to all other models, is still able to reconstruct the chaotic behavior. When applied to real-world time series, such as traffic, weather, or cloud request data, where previous TS foundation models have shown inconsistent results [80], our model produces accurate forecasts which reflect the dynamics well (Fig. 8b-f, see Appx. A.2 for further details on the datasets). Our model seems to particularly excel on human physiological signals, such as functional magnetic resonance imaging (fMRI) or electroencephalographic (EEG) recordings, where all other models essentially produce meaningless dynamics, at best capturing the mean trend (Fig. 8g,h).
Table 1 quantitatively confirms that DynaMix mostly outperforms the other models tested on the DSR measures. Surprisingly, it often even outperforms them in terms of the forecasting error MAE (mean-absolute error), although its training corpus consisted purely of the simulated DS shown in Fig. 9, i.e. did not include any empirical data at all, let alone traffic, cloud usage, electricity, or temperature data! This is in stark contrast to the other TS foundation models tested here, which mostly had examples of such empirical data in their training corpus and hence might be expected to have an edge over DynaMix. Many real-world data, like weather, human fMRI, or EEG data, are known to bear signatures of deterministic chaos [56, 48], potentially at least partly explaining DynaMix’ strongly competitive forecasts, although it has never seen such type of data in training.
5 Conclusions
Here we introduce the first DSR foundation architecture that achieves zero-shot reconstructions of novel DS, accurately reproducing their long-term statistics, just from a provided context signal without any retraining or fine-tuning. It hugely outperforms current TS foundation models like Chronos in terms of DSR quality, although much more lightweight with only a fraction (0.1%) of the number of parameters. Consequently, DynaMix features inference times about an order of magnitude faster than the closest competitor. Most surprisingly, it even outperforms current TS foundation models on empirical data of types it has – in contrast to its competitors – never seen in training, often producing also better short-term forecasts. All of this is achieved with a fairly narrow training corpus consisting mainly of $3d$ synthetic data from chaotic systems, which suggests there may be considerable room for further improvement by extending DynaMix’ training to include, e.g., various empirical time series.
One important take-home from this work therefore is that foundation models built based on principles of DS theory may be able to profoundly improve performance of current zero-shot TS forecasters. Most (if not all) empirically observed time series come from some underlying DS, and acknowledging this fact in model training and construction may help to advance the field. Besides DynaMix’ specific DSR architecture, control-theoretically motivated training techniques, and embedding principles, also the fact that it has seen many chaotic systems in training may play a role. Most complex DS in nature [15, 82, 57], engineering [74], and society [54, 87] are likely chaotic, featuring an infinite spectrum of potential temporal patterns (unstable periodic orbits; [36]).
Limitations
For now, we have mainly focused on rather stationary synthetic and empirical time series (but see Fig. 8b). Changes in statistical properties over time, discontinuities, or widely differing time scales in the data observed, impose severe difficulties for zero-shot forecasting, as illustrated in Fig. 23a,b, which also plague other TS foundation models. However, this may partly be amended by explicitly including non-stationary DS or DS evolving on many different time scales, featuring sharp transitions, etc., in the training corpus. Incorrect embeddings, which may not be immediately obvious, can also lead to forecast failures (Fig. 23c).
Acknowledgements
This work was funded by the German Research Foundation (DFG) within Germany’s Excellence Strategy EXC 2181/1 – 390900948 (STRUCTURES), and through DFG individual grant Du 354/15-1 to DD.
Appendix A Appendix
A.1 Methodological details
Training method
For training our model, we used a variant of sparse teacher forcing (STF), a control-theoretic method for DSR, with the theoretical background and rationale developed in [56]. STF and related training methods [6, 8, 38, 56, 24] circumvent the exploding gradient problem when training on chaotic systems, while still enabling the DSR model to sufficiently forward-iterate trajectories into the future to capture long-term properties (see [56] for details), and produces state-of-the-art results for DSR [7, 9]. In STF, current latent states states $\bm{z}_{t}$ are replaced with states inferred from the data at fixed intervals $\tau$ by (pseudo-)inversion of the decoder model, $\tilde{\bm{z}}_{t}=g^{-1}(\bm{x}_{t})$, thus recalibrating the trajectory at times ideally chosen based on the system’s Lyapunov spectrum [56] (or simply treating $\tau$ as a hyper-parameter). In our case of an identity mapping from a latent subspace to the observations, $\hat{\bm{x}}_{t}=\mathcal{I}_{(N\times M)}\bm{z}_{t}$, this inversion becomes trivial, where $N$ is the number of observations (and thus readout neurons) and $\mathcal{I}_{(N\times M)}$ an $N\times M$ matrix with $\mathcal{I}_{(N\times M),rr}=1$ for $r\leq N$ and zeros else. Thus, during training, we force
$\bm{z}_{t+1}=\begin{cases}F_{\bm{\theta}}(\tilde{\bm{z}}_{t},\bm{C})&\text{if }t\in\mathcal{T}=\{l\tau+1\}_{l\in\mathbb{N}_{0}}\\ F_{\bm{\theta}}(\bm{z}_{t},\bm{C})&\text{otherwise, }\end{cases}$ (8)
where $\tilde{\bm{z}}_{t}=(\bm{x}_{t},\bm{z}_{N+1:M,t})^{T}$, $F_{\theta}$ is the mixture of experts using the context $\bm{C}$, and $\tau=10$ here (see Fig. 24). Importantly, STF is only used for training the model, applied after calculating the loss, and is turned off at test time.
For the loss, we simply use the standard mean squared error (MSE) between model predictions $\hat{\bm{X}}$ and ground truth observations $\bm{X}$,
$\mathcal{L}_{MSE}(\hat{\bm{X}},\bm{X})=\frac{1}{N\cdot(T-T_{C}+\Delta t)}\sum_{t=T_{C}-\Delta t+1}^{T}\left\lVert\hat{\bm{x}}_{t}-\bm{x}_{t}\right\rVert_{2}^{2}.$ (9)
To this we add a regularization term meant to encourage the model to explore a wider region of state space based on the context, by enhancing the variance $\bm{\Sigma}$ of the exploration noise in the state attention:
$\mathcal{L}_{reg}=\lambda\left(1+\sum_{i}\Sigma_{ii}\right)^{-1}\;,$ (10)
where we chose $\lambda=0.01$. Rectified adaptive moment estimation (RADAM) [50] was employed as the optimizer, with $L=50$ batches of $S_{B}=16$ sequences per epoch, each of length $T=550$, and $2000$ epochs in total. We used a learning rate exponentially decaying from $\eta_{\text{start}}=2\cdot 10^{-3}$ to $\eta_{\text{end}}=10^{-5}$. The context length used in training was set to $T_{C}=500$, and the window of overlap with the model-generated time series to $\Delta t=50$, see sect. 3.2. Training was performed on a single CPU (18-Core Xeon Gold 6254), and a single epoch took $60$ seconds depending on sequence length and model size. At test time, no retraining or fine-tuning is performed, but utilizing the context $\bm{C}$ the model just forward-iterates from the last context time step. Hyperparameters for training were partly selected according to previous results in [7, 6], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.01,0.02\}$, $\Delta t=\{0,50,100\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
Model details
We use $J=20$ AL-RNN experts for our model. Each expert has a latent dimension of $M=20$, of which $P=7$ are rectified-linear units (ReLUs). As suggested in [7], we only use the first $N$ of the $M-P$ linear units for the readout. We followed the initialization protocol in [6, 77], drawing $\bm{W}$ from a Gaussian with mean $\bm{0}$ and $\bm{\sigma}=0.01\mathds{1}$, setting $\bm{h}=\mathbf{0}$, and $\bm{A}$ to the diagonal of a normalized positive-definite random matrix. The initial latent state was estimated as
$\bm{z}_{1}=\begin{bmatrix}\bm{x}_{1}\\ \bm{Lx}_{1}\end{bmatrix},$ (11)
where $\bm{L}\in\mathbb{R}^{(M-N)\times N}$ is jointly learned with other model parameters.
[…]
Performance measures
As in [46, 56, 38, 8, 61, 30, 93], we assess the geometric (dis)similarity between true and model-generated attractors using a Kullback-Leibler divergence ($D_{\text{stsp}}$) defined across state space. Specifically, $D_{\text{stsp}}$ determines the overlap between the distributions of true trajectory points, $p_{true}(\bm{x})$, and of model-generated trajectories, $p_{gen}(\bm{x}|\bm{z})$, given by
$D_{\text{stsp}}(p_{true}(\bm{x})\lVert p_{gen}(\bm{x}|\bm{z}))=\int p_{true}(\bm{x})\log\frac{p_{true}(\bm{x})}{p_{gen}(\bm{x}|\bm{z})}d\bm{x}.$ (12)
While especially for higher-dimensional state spaces these distributions may be approximated by Gaussian mixture models [46], here a discretized version was sufficient, with the state space parcellated into $K=m^{N}$ bins, where $m$ is the number of bins per dimension and $N$ the dimensionality of the system. We estimate the occupation probabilities $p_{i}$ of each bin via the relative frequencies $\hat{p}_{i}$ of trajectory visits and approximate $D_{\text{stsp}}$ as
$D_{\text{stsp}}\approx\sum_{i=1}^{K}\hat{p}_{true;i}\log\frac{\hat{p}_{true;i}}{\hat{p}_{gen;i}}.$ (13)
We set $m=30$ bins per dimension, following [37]. To ensure a steady-state distribution is reached, long trajectories of $T=10,000$ time steps are sampled.
To evaluate the agreement in long-term temporal dynamics, we compute the Hellinger distance $D_{H}$ between the power spectra of the true and model-generated time series [56, 38], defined as
[…]
where $F(\omega)$ and $G(\omega)$ are the power spectra of the true and generated time series, respectively. Power spectra are obtained via the dimension-wise Fast Fourier Transform (FFT), smoothed using a Gaussian kernel (with $\sigma=20$ for DS and $\sigma=2$ for the short empirical time series), and normalized to ensure comparability. High-frequency tails dominated by noise are truncated as described in [38]. The aggregated Hellinger distance $D_{H}$ is then computed as the average across all dimension-wise spectral comparisons.
The $n$-step ahead prediction error to assess short-term forecast quality (with $n=10$ in our evaluation) is given by
$\text{PE}(n)=\left\|\bm{x}_{T_{C}+n}-\hat{\bm{x}}_{T_{C}+n}\right\|_{1}.$ (15)
The MAE for evaluating short term predictions on the empirical time series [39] is given by
$\text{MAE}=\frac{1}{n}\sum_{t=T_{C}+1}^{T_{C}+n}\left|x_{t}-\hat{x}_{t}\right|\;,$ (16)
where we chose $n$ according to the data’s temporal scale and resolution ({ETTh1,fMRI}: $n=40$,{traffic, cloud requests, partially obs. DS}: $n=80$, {human EEG, weather air pressure}: $n=120$, {weather temperature}: $n=200$).
Lyapunov exponent
For numerically estimating the maximum Lyapunov exponent $\lambda_{\text{max}}$ from a one-dimensional empirical time series $\{x_{t}\},t=1\dots T$, one common approach is the Rosenstein algorithm [68]. First, the time series is embedded into a state space using time-delay embedding as in eq. 5 (here using the PECUZAL algorithm [47]). For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity [68]) and $l_{s}=0.2$. The local divergence between initially nearby trajectories is then tracked over a range of time steps $k=0\dots k_{max}$,
$d_{i}(k)=\|\bm{x}_{i+k}-\bm{x}_{j(i)+k}\|_{2}\;.$ (17)
The maximum Lyapunov exponent is given by
$d_{i}(k)\approx d_{i}(0)e^{\lambda_{\text{max}}k\Delta t}\;,$ (18)
where $\Delta t$ is the temporal resolution of the series. Taking the logarithm and averaging over all pairs leads to
$\langle\ln d(k)\rangle\approx\lambda_{\max}k\Delta t+C.$ (19)
In chaotic systems, for a suitable choice of $k_{max}$ (here: $50$), this quantity initially grows approximately linearly with time [42], with the slope given by $\lambda_{\max}>0$. In contrast, periodic (cyclic) systems have $\lambda_{\max}=0$.
A.2 Datasets
Training data
DynaMix is trained on about $1.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in [29]. Gaussian noise of $5\%$ of the data standard deviation was added to all dimensions, and time series were standardized to ensure comparable scaling of all data. Attractor dynamics were chosen to reflect different types of behavior, see Fig. 9.
Figure 9: Training data (blue) and their reconstruction in red.
Test data
Our test set for DS consists of simulated time series of length $10^{5}$ sampled from $54$ different $3d$ DS collected in [29], which are not part of the training set. Furthermore, we evaluated the Selkov 2d DS [73] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
$\frac{dx}{dt}$ $=-x+ay+x^{2}y,$ (20)
$\frac{dy}{dt}$ $=b-ay-x^{2}y,$
where we chose $a=0.1$ and $b=0.5$.
For evaluation of time series forecasting, different real-world data were used. The traffic data are hourly recordings of the number of cars passing road junctions (https://www.kaggle.com/datasets/fedesoriano/traffic-prediction-dataset/data). The cloud data, also used in [80] to evaluate TS foundation models, are publicly available from Huawei Cloud https://github.com/sir-lab/data-release. It consists of function requests from Huawei’s serverless cloud platform. The weather data consists of daily sampled soil temperature and air pressure measured in the city of Mannheim, Germany (source: Deutscher Wetterdienst [German weather service]), and can be accessed via https://www.dwd.de/EN/ourservices/cdc/cdc_ueberblick-klimadaten_en.html. The functional magnetic resonance imaging (fMRI) data comes from human subjects performing cognitive tasks and is publicly available on GitHub [48]. We followed Kramer et al. [48] and selected the first principal component of BOLD activity in each of the $20$ brain regions. Kramer et al. [48] report a positive maximum Lyapunov exponent for models reconstructed from these time series, indicating their chaotic nature (see also [84]). The ETTh1 dataset is part of the Electricity Transformer Temperature (ETT) benchmark, which is widely used for evaluating TS forecasting models. It contains hourly data collected from a power transformer station [94] and can be accessed at https://github.com/zhouhaoyi/ETDataset. Electroencephalogram (EEG) data were taken from a study by Schalk et al. [71], comprising 64-channel data collected from human subjects performing various motor and imagery tasks. Following the approach of Brenner et al. [6], the signals were smoothed using a Hann window of length $15$. As for the training data, all test data were standardized for processing by the models.
A.3 TS foundation models
Chronos
Chronos is a recent TS foundation framework that adapts a transformer-based LLM architecture for probabilistic time series forecasting [2]. Key to this approach is the tokenization of real-valued time series observations through scaling and quantization, transforming them into token sequences accessible to language models. The model is pretrained on extensive datasets, including both synthetic data generated via Gaussian processes, as well as an extensive batch of empirical time series from [2, 32], including traffic, weather/ climate, electricity and web data, enabling it to achieve strong zero-shot performance across diverse datasets without task-specific fine-tuning. For our evaluation we used the standard pipeline as described in https://github.com/amazon-science/chronos-forecasting.
Mamba4Cast
Mamba4Cast is a zero-shot time series forecasting model based on the Mamba architecture [35], a type of linear (‘state-space’) RNN with nonlinear input & output gating, and inspired by Prior-data Fitted Networks (PFNs) [5]. Trained exclusively on synthetic data, Mamba4Cast can generate zero-shot forecasts when provided with time series context information. For evaluation we follow https://github.com/automl/Mamba4Cast.
Tiny Time Mixers
Tiny Time Mixers is another time series foundation model using a non-transformer architecture [26]. It builds upon the TSMixer backbone [25] to which it adds several key innovations such as adaptive patching, diverse resolution sampling, and multi-resolution prefix tuning to improve generalization across datasets with varying temporal resolutions. Evaluation here is performed as in https://github.com/glehet/TTM1.
For comparability, all models were evaluated on the exact same CPU (18-Core Xeon Gold 6254) and GPU (Nvidia RTX 2080 Ti) using 512GB of RAM.
A.4 Further results
Figure 10: Zero-shot DSR (red) from a $2000$-step context of unseen DS not contained in the training corpus (ground truth in blue).
Figure 11: Zero-shot DSR of chaotic Lorenz-63 system (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
Figure 12: Zero-shot DSR of Lorenz-63 system in cyclic regime (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
Figure 13: Zero-shot DSR of chaotic finance system (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
Figure 14: Zero-shot DSR of chaotic Genesio Tesi system (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
Figure 15: Comparison of zero-shot forecasting of finance system for DynaMix vs. different TS foundation models.
Figure 16: Comparison of zero-shot forecasting of Sprott D system for DynaMix vs. different TS foundation models.
Figure 17: Comparison of zero-shot forecasting of Sprott M system for DynaMix vs. different TS foundation models.
Figure 18: Comparison of zero-shot forecasting of Vallise El Nino system for DynaMix vs. different TS foundation models.
Figure 19: Comparison of zero-shot forecasting of Sprott C system for DynaMix vs. different TS foundation models.
Figure 20: Comparison of zero-shot forecasting of Sprott A system for DynaMix vs. different TS foundation models.
Figure 21: Zero-shot DSR performance across all 54 test set DS for DynaMix and various TS foundation models as in Fig. 4, but for context length $T_{C}=512$. Detailed values in Table LABEL:tab:performance_DSR_CL512.
Figure 22: Exponential divergence of two initially nearby trajectories (in blue and red) illustrated for the chaotic Lorenz-63 system. The prediction error is still sensible on a short time scale ($\text{MAE}=0.09$), but then rapidly increases and breaks down as a suitable metric in the longer-term ($\text{MAE}=1.12$), although both trajectories were drawn from the exact same system with the very same parameters.
Figure 23: Failure of proper zero-shot DSR. a) DynaMix struggles with sharp peaks in an otherwise much more slowly evolving time series. b) As quantified in Fig. 3, selecting a too broad temporal resolution of the context time series can lead to improper reconstructions with, in this case, temporal disalignment w.r.t. the true signal. However, this is a general problem for any TS foundation model, and Chronos fares even much worse in this example. c) Zero-shot DSR may fail (red) as well if the context time series is not properly embedded. In this example, the required embedding dimension exceeded the model capacity. However, as shown in orange, this can be amended by using the positional encoding, eq. 6.
Figure 24: DSR performance for different teacher forcing intervals $\tau$, illustrating an optimal value is essential for successful training, see [56, 6].
Table 2: Performance comparison for zero-shot inference across all DS in the dataset from [29] using $T_{C}=2000$. Best value in red, second best in blue.
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Atmospheric Regime 9.7 0.6 0.3 9.2 0.4 0.2 10.5 0.4 0.2 12.0 0.5 0.1 13.3 0.5 0.0 13.7 0.5 0.3 11.6 0.9 0.3 12.3 0.8 0.2 12.3 0.9 0.3 11.7 0.7 0.3 11.2 0.8 0.6 13.9 0.4 1.3
BeerRNN 8.2 0.4 3.6 12.1 0.5 0.8 9.3 0.4 1.3 13.3 0.4 1.6 10.9 0.3 1.3 13.1 0.5 1.5 12.4 0.9 1.7 12.7 0.8 2.1 11.9 0.9 2.0 11.7 0.8 1.9 11.4 0.8 1.1 14.2 0.4 2.5
Bouali 4.7 0.1 0.7 7.0 0.3 0.5 12.1 0.2 0.6 11.6 0.1 0.6 13.1 0.2 0.3 13.5 0.4 1.5 11.2 0.7 2.2 12.0 0.6 3.4 12.3 0.5 2.2 9.8 0.5 1.5 12.9 0.7 2.6 14.0 0.3 2.5
Burke Shaw 3.4 0.1 0.2 8.0 0.1 0.3 5.6 0.1 0.2 5.1 0.2 0.1 11.1 0.2 0.4 13.2 0.6 1.0 11.0 0.6 0.9 11.0 0.7 0.5 11.7 0.8 0.2 10.0 0.7 1.1 10.2 0.7 0.6 12.9 0.4 0.8
Ca Two Plus Quasip. 3.3 0.1 0.3 9.6 0.5 1.2 9.5 0.3 0.3 12.6 0.2 0.9 12.7 0.4 0.6 13.2 0.4 1.6 9.4 0.7 0.8 9.4 0.6 1.5 9.6 0.7 0.8 11.5 0.5 0.9 11.1 0.6 2.0 13.8 0.2 5.2
Chen Lee 5.1 0.4 0.1 4.2 0.2 0.0 3.6 0.2 0.0 3.9 0.2 0.0 5.3 0.3 0.0 12.0 0.8 0.3 4.5 0.8 0.3 8.8 0.7 1.1 7.7 0.7 0.5 7.2 0.7 0.5 11.3 0.8 0.2 8.9 0.4 0.6
Chua 3.0 0.1 0.4 12.9 0.2 0.4 12.4 0.4 0.4 13.3 0.2 0.1 13.5 0.4 0.4 13.5 0.3 0.9 11.9 0.6 1.0 11.7 0.4 0.6 11.7 0.4 0.6 12.0 0.4 0.7 12.1 0.4 0.2 13.4 0.3 1.0
Coullet 0.8 0.0 0.3 0.6 0.0 0.1 6.3 0.0 0.1 3.9 0.1 0.1 13.7 0.3 0.1 14.6 0.3 0.6 11.4 0.7 0.6 12.8 0.7 0.4 11.8 0.6 0.7 10.7 0.6 1.1 13.3 0.6 0.7 15.0 0.2 1.8
Dadras 3.0 0.2 13.9 10.0 0.5 9.2 9.2 0.5 4.2 11.8 0.4 7.5 11.4 0.3 4.7 12.9 0.8 9.0 10.5 0.8 8.9 9.4 0.7 9.0 11.0 0.6 8.8 8.8 0.5 8.9 13.1 0.8 7.9 12.2 0.4 9.2
Dequan Li 10.0 0.5 0.1 12.6 0.5 0.1 13.2 0.4 0.1 13.1 0.3 0.1 12.9 0.4 0.5 13.1 0.3 0.4 12.4 0.4 0.4 12.7 0.4 0.3 12.2 0.4 0.3 12.6 0.4 0.6 12.3 0.4 1.1 13.0 0.4 0.4
Finance 3.0 0.2 0.6 8.5 0.4 0.2 9.4 0.6 0.3 13.4 0.3 0.5 12.2 0.4 0.3 13.6 0.5 0.3 12.0 0.7 1.3 11.7 0.6 2.3 12.4 0.7 1.3 11.8 0.6 1.4 12.6 0.8 2.3 14.0 0.3 2.1
Genesio Tesi 10.5 0.2 0.3 8.9 0.1 0.1 8.7 0.1 0.0 7.7 0.1 0.0 11.8 0.4 0.2 12.3 0.3 0.3 10.7 0.3 0.2 11.4 0.9 0.8 10.6 0.6 1.1 10.0 0.4 1.6 10.4 0.8 0.4 12.3 0.2 0.7
Guckenheimer Holmes 7.9 0.0 0.2 10.6 0.4 0.1 11.7 0.2 0.0 11.3 0.4 0.1 13.2 0.7 0.1 12.5 0.2 0.1 10.7 0.6 0.1 10.6 0.6 0.2 9.2 0.6 0.1 10.3 0.6 0.1 10.4 0.5 0.5 11.6 0.3 0.6
Hadley 8.4 0.3 0.2 10.8 0.5 0.2 9.3 0.6 0.2 11.9 0.4 0.2 12.5 0.4 0.2 12.5 0.3 0.4 10.9 0.7 0.4 10.7 0.7 1.0 9.5 0.6 0.3 10.2 0.4 0.4 11.7 0.7 0.8 12.5 0.3 1.6
Halvorsen 2.8 0.1 0.2 6.0 0.1 0.2 8.4 0.1 0.2 5.9 0.1 0.3 12.6 0.2 0.3 13.3 0.3 1.3 11.8 0.8 1.1 9.5 0.7 0.8 8.9 0.8 0.4 10.4 0.5 0.4 12.2 0.6 1.3 13.2 0.3 1.5
Hindmarsh Rose 6.8 0.5 0.6 9.7 0.3 0.1 6.7 0.5 0.3 8.4 0.1 0.1 11.5 0.4 0.7 13.8 0.6 1.0 12.7 0.7 1.3 12.5 0.7 1.5 13.9 0.6 1.1 13.4 0.7 1.1 13.9 0.6 0.7 14.0 0.2 1.6
Isothermal Chemical 2.5 0.2 0.1 3.7 0.1 0.0 7.9 0.3 0.0 9.7 0.3 0.0 9.0 0.3 0.0 13.8 0.5 0.2 7.9 0.8 0.3 5.9 0.8 0.2 8.4 0.8 0.2 8.5 0.8 0.2 12.8 0.8 0.5 15.6 0.2 2.8
Itik Banks tumor 9.1 0.2 1.2 10.0 0.6 0.5 8.4 0.5 0.3 10.5 0.4 0.3 12.8 0.3 0.2 12.1 0.7 0.4 9.2 0.7 1.2 9.5 0.7 1.2 9.7 0.8 1.2 9.6 0.7 1.0 13.3 0.9 0.5 13.2 0.3 1.7
Jerk Circuit 2.7 0.3 0.4 10.1 0.4 1.0 9.7 0.4 0.4 7.5 0.3 0.4 11.1 0.4 0.4 12.3 0.3 0.2 10.8 0.8 1.6 10.9 0.5 2.0 10.5 0.9 1.9 10.1 0.9 1.8 11.1 0.9 1.8 13.9 0.4 2.1
Laser 6.5 0.3 0.1 11.1 0.4 0.1 10.8 0.2 0.1 12.8 0.2 0.2 13.8 0.5 0.4 14.1 0.5 0.3 11.8 0.7 0.3 11.5 0.7 0.4 11.0 0.7 0.5 11.9 0.7 1.6 13.3 0.6 0.7 13.8 0.2 1.2
Lorenz 1.1 0.1 0.7 6.6 0.2 0.3 6.0 0.2 0.3 10.0 0.2 0.2 12.0 0.4 0.3 13.0 0.5 0.3 10.9 0.7 0.3 11.3 0.5 0.6 12.7 0.6 0.9 12.8 0.3 1.8 11.4 0.7 0.2 13.0 0.3 1.9
Lorenz 84 6.9 0.3 4.9 9.5 0.4 5.2 11.0 0.4 4.8 11.5 0.4 3.2 12.2 0.4 2.5 12.9 0.5 4.4 12.9 0.8 2.4 12.9 0.8 2.5 11.5 0.9 2.4 12.3 0.7 2.4 13.4 0.8 3.3 13.4 0.5 2.4
Lu Chen 3.4 0.5 0.8 6.3 0.4 0.2 5.3 0.3 0.3 8.0 0.1 0.2 7.1 0.1 0.2 4.9 0.2 0.1 11.6 0.9 1.3 11.9 1.0 1.2 11.3 1.0 1.9 11.4 1.0 1.4 9.7 0.8 0.7 13.7 0.4 3.4
Moore Spiegel 5.9 0.4 1.0 10.4 0.6 1.0 10.0 0.5 0.0 12.5 0.4 0.2 9.9 0.5 0.6 12.1 0.3 0.9 10.5 0.9 2.8 11.4 0.9 1.8 10.1 0.9 2.5 9.8 0.7 3.4 9.4 0.8 4.6 12.5 0.3 3.6
Multi Chua 7.8 0.3 0.8 9.3 0.5 0.0 9.3 0.2 0.2 7.1 0.4 0.1 13.3 0.2 0.4 14.0 0.3 1.2 11.5 0.6 1.8 12.2 0.4 3.1 9.5 0.5 2.6 10.1 0.4 2.4 10.6 0.4 1.2 14.0 0.3 2.8
Nose Hoover 10.0 0.3 0.3 9.9 0.5 0.1 9.6 0.6 0.1 11.5 0.3 0.5 12.5 0.7 0.5 13.0 0.5 0.5 10.6 0.9 0.8 10.6 0.9 0.7 10.1 0.7 0.7 10.2 0.5 0.4 10.2 0.7 0.5 12.9 0.4 0.4
Pehlivan Wei 4.0 0.2 0.3 8.5 0.4 0.4 10.1 0.4 0.3 11.3 0.3 0.3 10.8 0.4 0.3 12.6 0.5 0.7 9.2 0.7 1.0 10.1 0.8 1.0 9.2 0.7 0.6 10.6 0.8 1.3 9.3 0.7 1.0 12.2 0.6 1.9
Sakarya 1.5 0.2 0.3 5.6 0.5 0.2 5.0 0.3 0.2 7.5 0.2 0.3 11.1 0.4 0.2 12.5 0.4 0.4 10.2 0.9 0.4 10.0 0.6 1.4 7.7 0.9 0.9 9.1 0.4 0.8 13.1 0.8 0.8 12.6 0.4 0.7
San Um Srisuchinwong 4.1 0.2 0.2 12.9 0.4 1.0 12.7 0.5 1.6 11.4 0.5 1.0 13.3 0.5 1.2 13.0 0.6 1.1 12.7 0.7 0.9 10.4 0.7 0.9 13.4 0.6 1.2 11.7 0.7 1.5 10.7 0.7 1.6 13.1 0.4 2.4
Sprott A 2.7 0.1 0.3 8.0 0.2 0.3 8.9 0.3 0.3 6.6 0.2 0.2 11.8 0.3 0.3 12.3 0.3 1.1 10.5 0.8 1.0 8.1 0.5 1.6 10.0 0.7 0.7 10.4 0.7 1.0 9.9 0.7 0.8 12.0 0.3 3.0
Sprott B 2.4 0.2 0.3 4.3 0.2 0.8 5.6 0.4 1.6 11.7 0.5 0.2 12.7 0.4 0.6 13.2 0.7 1.7 11.8 0.8 0.8 11.5 0.6 0.6 12.3 0.8 1.2 10.6 0.7 1.4 12.5 0.8 1.8 12.6 0.4 1.5
Sprott C 2.3 0.1 0.6 11.6 0.4 1.2 12.4 0.4 0.7 11.9 0.4 1.2 13.3 0.3 0.3 13.1 0.4 2.4 10.0 0.5 4.7 9.2 0.5 3.9 10.7 0.5 5.3 9.3 0.5 4.3 12.8 0.6 2.9 13.3 0.2 2.4
Sprott D 4.0 0.2 0.3 9.8 0.6 0.1 10.3 0.4 0.1 12.9 0.3 0.2 13.4 0.5 0.4 13.4 0.4 0.1 12.2 0.7 0.2 10.9 0.6 0.2 9.9 0.5 0.6 11.5 0.6 0.6 12.9 0.7 0.2 14.0 0.2 1.4
Sprott F 3.0 0.2 0.3 9.6 0.5 0.3 9.2 0.5 0.6 12.3 0.3 0.4 12.6 0.4 0.7 12.7 0.4 0.6 9.9 0.7 1.5 10.4 0.7 1.3 10.4 0.6 1.0 12.1 0.5 1.1 11.4 0.8 1.1 13.3 0.3 0.9
Sprott G 2.3 0.1 0.7 7.0 0.6 0.1 7.7 0.5 0.1 10.0 0.3 0.3 7.2 0.1 0.4 13.0 0.3 0.6 10.9 0.9 1.2 11.4 0.6 1.3 9.2 0.8 1.4 10.8 0.8 1.7 14.1 0.8 0.7 13.1 0.2 1.3
Sprott I 2.7 0.1 0.5 9.7 0.5 0.4 11.3 0.2 0.4 9.0 0.1 0.4 11.9 0.3 0.5 13.7 0.2 1.1 10.2 0.9 2.1 9.8 0.9 3.4 8.4 0.6 2.9 9.1 0.6 2.8 12.2 0.7 1.9 14.4 0.2 6.2
Sprott L 3.4 0.1 0.4 8.0 0.5 0.3 6.8 0.1 0.3 8.2 0.2 0.1 11.6 0.3 0.2 13.0 0.2 0.1 10.5 0.9 1.2 10.4 0.8 1.8 11.4 0.9 1.2 12.1 0.7 1.0 10.8 0.8 1.3 14.2 0.3 3.1
Sprott M 3.8 0.1 0.4 9.8 0.4 0.0 10.5 0.4 0.1 12.4 0.1 0.1 12.9 0.2 0.2 14.0 0.6 0.6 9.8 0.8 1.4 10.3 0.6 1.6 9.9 0.8 1.5 11.8 0.6 1.6 11.6 0.7 1.5 14.3 0.2 3.2
Sprott O 1.3 0.1 0.9 3.6 0.1 0.3 4.6 0.0 0.5 9.7 0.1 0.4 12.4 0.4 0.4 12.9 0.4 0.9 10.8 1.0 1.3 12.6 0.9 1.3 10.4 0.8 1.6 11.5 0.9 1.1 10.7 0.8 0.6 14.0 0.4 2.0
Sprott P 5.5 0.2 0.6 10.8 0.3 2.2 11.9 0.3 1.3 12.7 0.3 1.1 12.7 0.3 1.2 13.3 0.3 2.0 7.3 0.5 3.1 10.6 0.6 4.5 8.3 0.6 4.2 10.0 0.6 5.0 11.1 0.7 2.7 12.2 0.3 5.7
Sprott Q 4.5 0.1 0.5 10.1 0.2 0.2 7.6 0.1 0.2 7.3 0.1 0.3 11.6 0.2 0.2 13.2 0.3 0.8 11.8 0.8 2.5 11.6 0.7 1.1 11.9 0.6 0.5 11.2 0.6 2.8 10.9 0.7 0.7 13.1 0.4 4.3
Sprott R 3.5 0.3 0.4 9.3 0.8 0.1 6.7 0.5 0.2 11.0 0.5 0.4 12.6 0.5 0.6 13.1 0.4 0.3 10.6 0.8 0.4 10.4 0.5 0.3 9.6 0.7 0.8 10.5 0.7 0.8 11.8 0.8 0.8 13.8 0.3 1.3
Sprott S 5.3 0.1 0.6 10.1 0.9 1.7 9.9 0.6 0.7 12.3 0.2 0.9 12.6 0.4 0.4 13.1 0.3 1.1 10.6 0.7 0.6 9.7 0.8 3.5 11.0 0.6 1.8 10.4 0.4 3.7 11.8 0.8 1.8 11.6 0.3 4.6
Sprott Torus 3.2 0.2 0.1 5.4 0.7 0.4 10.4 0.7 0.1 14.3 0.6 0.2 14.5 0.6 0.4 13.8 0.7 0.3 6.2 0.8 2.9 8.3 0.6 5.2 5.3 0.8 5.4 5.8 0.7 4.1 13.6 0.8 1.0 15.5 0.3 2.1
Thomas 1.4 0.3 2.1 10.6 0.5 3.8 10.1 0.8 3.9 9.9 0.3 3.7 13.9 0.5 2.6 13.9 0.7 2.8 12.6 0.8 1.9 12.6 0.8 2.0 12.8 0.8 2.0 12.9 0.8 2.0 10.7 0.5 3.0 13.3 0.2 2.6
Thomas Labyrinth 4.4 0.1 1.8 12.5 0.2 2.8 12.5 0.3 2.9 12.4 0.4 3.1 12.5 0.4 2.9 12.5 0.3 3.8 11.4 0.5 2.7 12.5 0.5 2.2 11.9 0.5 2.5 11.7 0.5 2.1 12.3 0.5 3.1 12.4 0.2 1.8
Torus 11.5 0.9 2.7 9.8 0.3 0.1 9.4 0.2 0.2 10.9 0.3 0.2 12.3 0.3 0.3 12.4 0.5 0.3 12.1 0.9 1.9 10.5 1.0 2.0 11.6 0.9 1.8 11.9 0.5 1.9 9.1 0.8 1.9 12.6 0.5 3.2
Tsucs 2 5.7 0.6 0.4 11.6 0.3 0.1 10.3 0.5 0.2 10.6 0.3 0.3 11.6 0.6 0.2 13.8 0.7 0.9 10.6 0.7 0.4 11.2 0.8 0.4 8.2 0.7 0.4 10.5 0.8 0.4 11.1 0.7 0.4 8.4 0.4 0.5
Vallise El Nino 1.2 0.1 0.3 8.4 0.3 0.6 11.2 0.5 1.2 12.5 0.3 1.3 12.7 0.5 1.9 13.5 0.6 3.2 12.3 0.7 2.5 13.4 0.5 2.7 12.3 0.7 3.3 12.6 0.5 1.2 13.9 0.8 2.9 13.7 0.3 3.7
Wang Sun 2.8 0.2 0.1 13.2 0.5 0.5 12.2 0.5 0.2 11.8 0.5 0.5 11.1 0.4 0.5 14.6 0.7 1.2 13.9 0.7 0.9 10.9 0.7 0.9 10.8 0.6 1.1 10.9 0.7 0.8 13.3 0.7 0.6 11.1 0.4 0.8
Winsmi Reduced 6.5 0.4 0.2 11.4 0.4 0.1 10.7 0.5 0.1 13.5 0.6 1.1 13.7 0.6 0.1 14.0 0.4 1.3 11.9 0.9 3.7 12.2 0.8 4.0 10.5 0.6 3.6 9.4 0.3 2.3 12.0 0.8 3.2 14.2 0.2 3.3
Yu Wang 5.5 0.2 0.6 9.8 0.7 0.8 8.4 0.3 0.3 12.4 0.3 0.2 12.4 0.5 0.7 13.0 0.6 0.9 12.4 0.9 3.6 12.5 0.7 3.2 12.8 0.8 3.3 11.9 0.9 3.9 12.5 0.8 2.3 13.5 0.4 2.9
Yu Wang 2 3.8 0.0 0.1 9.6 0.5 0.1 10.0 0.6 0.1 12.4 0.3 0.2 12.2 0.3 0.1 14.1 0.7 0.2 9.9 0.2 0.2 9.5 0.7 0.4 9.9 0.6 0.6 10.4 0.4 0.3 10.6 0.7 0.6 12.6 0.2 0.4
Zhou Chen 4.9 0.3 0.2 11.1 0.3 0.5 5.9 0.4 0.2 11.6 0.3 0.9 14.2 0.4 0.7 10.3 0.3 0.8 12.8 0.4 0.7 12.7 0.5 0.9 13.5 0.4 1.0 12.9 0.5 1.2 13.1 0.5 0.6 13.2 0.2 2.1
Table 3: Performance comparison for zero-shot inference across all DS in the dataset from [29] using $T_{C}=512$. Best value in red, second best in blue.
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Atmospheric Regime 8.4 0.6 2.8 9.7 0.5 0.4 11.1 0.4 0.6 10.5 0.6 0.6 10.1 0.5 1.3 12.0 0.4 0.5 11.9 0.9 3.6 12.0 0.9 3.8 13.5 0.9 2.7 12.7 0.8 4.9 11.3 0.9 1.2 14.2 0.6 3.6
BeerRNN 11.7 0.8 4.7 8.0 0.5 0.5 3.3 0.2 0.4 9.5 0.2 0.2 7.3 0.4 0.2 11.9 0.4 0.3 12.3 0.9 1.4 12.5 0.9 2.2 12.1 0.9 1.8 12.6 0.8 1.4 13.0 0.9 1.9 14.1 0.6 4.1
Bouali 4.1 0.1 0.2 8.3 0.5 0.0 7.3 0.3 0.1 8.6 0.1 0.0 11.1 0.3 0.1 13.2 0.5 0.1 10.6 0.8 0.9 9.6 0.7 1.1 12.2 0.7 1.2 12.3 0.8 0.8 12.9 0.8 1.5 13.4 0.4 1.1
Burke Shaw 4.1 0.2 0.6 10.8 0.5 1.5 9.0 0.5 2.4 13.0 0.4 2.0 13.2 0.4 1.9 13.3 0.6 1.8 11.5 0.8 0.6 12.1 0.6 1.0 12.2 0.7 1.2 10.5 0.7 2.9 12.2 0.9 1.3 13.7 0.5 2.6
Ca Two Plus Quasip. 2.3 0.1 0.1 7.2 0.1 0.1 8.3 0.1 0.1 9.2 0.4 0.1 13.1 0.4 0.1 13.2 0.4 0.2 11.3 0.9 0.8 11.6 0.8 1.2 12.8 0.6 0.3 8.9 0.6 1.6 11.1 0.8 1.6 13.8 0.7 1.6
Chen Lee 3.9 0.4 0.3 4.9 0.3 0.9 4.0 0.5 1.6 4.8 0.3 0.4 4.6 0.3 1.0 8.9 0.7 1.3 7.2 0.9 3.6 7.9 0.7 2.1 8.0 0.7 1.7 4.8 0.5 1.4 11.0 0.9 1.2 6.4 0.5 4.0
Chua 5.9 0.1 0.1 7.4 0.2 0.2 5.9 0.2 0.2 7.6 0.1 0.2 11.7 0.2 0.1 13.1 0.3 0.4 11.8 0.6 1.0 12.2 0.5 1.3 11.7 0.6 0.6 12.0 0.6 1.4 11.3 0.5 1.5 13.4 0.4 0.8
Coullet 1.2 0.0 0.4 4.9 0.0 0.2 0.7 0.0 0.1 2.8 0.0 0.1 7.5 0.2 0.0 14.6 0.4 0.3 13.0 0.8 1.7 11.2 0.8 2.4 11.8 0.8 2.3 11.6 0.7 2.4 12.3 0.8 1.7 14.9 0.5 1.9
Dadras 4.1 0.2 1.5 7.7 0.6 0.9 4.7 0.5 1.2 13.1 0.5 0.5 12.1 0.4 0.3 10.2 0.7 1.0 12.7 0.8 2.6 12.9 0.7 2.4 13.0 0.8 2.7 8.8 0.8 4.2 11.9 0.8 2.5 12.8 0.4 2.8
Dequan Li 12.7 0.5 0.3 11.1 0.3 0.2 11.0 0.4 0.4 12.1 0.3 0.3 12.1 0.2 0.5 13.0 0.3 1.5 11.7 0.4 1.1 11.3 0.4 1.0 9.6 0.4 0.7 11.0 0.4 0.4 10.7 0.4 0.6 13.5 0.4 2.5
Finance 3.2 0.1 0.4 8.5 0.6 0.9 8.6 0.4 0.7 9.1 0.4 0.3 12.5 0.3 0.1 13.8 0.4 1.4 11.6 0.9 1.8 10.7 0.8 2.0 13.4 0.8 1.5 11.8 0.8 1.1 11.1 0.9 0.8 13.7 0.5 1.3
Genesio Tesi 8.2 0.2 0.3 7.9 0.4 0.7 8.5 0.2 0.7 7.7 0.2 0.7 9.4 0.2 0.8 13.0 0.3 1.6 10.4 0.7 0.7 10.6 0.7 0.8 10.3 0.8 1.0 9.6 0.8 1.5 11.0 0.9 0.9 12.6 0.6 4.1
Guckenheimer Holmes 9.6 0.1 0.2 11.8 0.1 0.4 12.0 0.1 0.4 12.6 0.3 0.2 12.5 0.4 0.6 12.1 0.4 0.2 9.6 0.7 0.2 9.6 0.7 0.4 9.4 0.6 0.2 11.8 0.6 0.4 10.2 0.7 0.5 12.7 0.5 0.6
Hadley 11.3 0.4 0.6 10.7 0.5 2.3 9.7 0.5 1.3 12.5 0.3 1.1 12.6 0.6 2.1 12.6 0.4 3.2 11.4 0.8 3.2 11.9 0.7 3.3 11.9 0.6 2.9 11.5 0.8 3.4 11.4 0.8 3.5 12.5 0.5 3.6
Halvorsen 4.3 0.3 1.0 9.7 0.6 0.3 8.2 0.2 0.8 7.5 0.2 0.8 13.1 0.2 0.6 13.3 0.4 1.3 10.0 0.8 4.7 10.3 0.7 4.4 9.7 0.8 4.4 10.9 0.7 4.3 11.3 0.9 3.4 13.4 0.5 3.5
Hindmarsh Rose 6.5 0.4 0.5 9.8 0.2 0.1 4.9 0.1 0.0 9.6 0.2 0.0 12.7 0.6 0.2 13.7 0.6 0.5 12.9 0.7 1.1 14.3 0.8 1.3 12.9 0.8 1.1 14.0 0.8 1.4 13.2 0.8 0.6 14.6 0.4 1.1
Isothermal Chemical 2.0 0.2 0.3 2.5 0.1 0.2 5.4 0.1 0.1 8.3 0.2 0.1 7.0 0.4 0.1 11.6 0.5 0.2 11.7 0.9 0.6 11.7 0.9 1.0 10.8 0.8 0.8 12.5 0.9 0.8 10.1 0.9 0.9 15.6 0.6 2.5
Itik Banks tumor 5.7 0.2 0.1 8.2 0.5 0.1 10.8 0.6 0.2 9.8 0.5 0.1 11.8 0.6 0.2 13.3 0.6 0.1 10.8 0.6 0.9 11.2 0.8 1.7 11.9 0.9 0.9 12.0 0.8 2.4 12.0 0.9 1.4 12.9 0.7 1.5
Jerk Circuit 4.3 0.3 0.4 8.8 0.6 1.1 9.1 0.4 0.8 10.6 0.3 1.0 11.7 0.3 0.2 9.8 0.6 1.1 12.5 0.9 1.9 13.5 0.9 1.6 12.9 0.9 0.7 13.1 1.0 1.5 10.3 0.8 0.5 14.1 0.5 3.1
Laser 4.5 0.2 0.5 7.5 0.5 5.0 8.2 0.3 1.9 11.5 0.2 0.6 12.8 0.2 0.7 13.9 0.6 1.3 12.6 0.7 2.4 12.8 0.5 1.7 13.0 0.7 1.3 12.8 0.7 2.7 11.8 0.8 2.4 13.3 0.5 2.2
Lorenz 3.4 0.2 0.6 9.8 0.6 0.5 8.5 0.3 0.6 13.0 0.4 0.7 12.6 0.7 0.8 13.3 0.6 0.3 12.8 0.8 1.0 13.0 0.7 1.9 13.2 0.6 1.4 13.0 0.6 1.3 11.8 0.8 1.3 13.4 0.5 1.1
Lorenz 84 7.9 0.2 1.6 10.7 0.5 3.3 11.0 0.4 4.4 12.6 0.4 4.8 11.0 0.4 0.8 13.2 0.6 1.8 11.8 0.9 2.0 12.2 0.9 2.6 11.5 0.9 2.4 12.3 0.9 2.5 11.2 0.9 2.2 13.4 0.6 2.9
Lu Chen 9.5 0.8 1.4 5.3 0.2 1.0 9.0 0.3 1.3 8.9 0.2 0.2 9.7 0.3 0.8 8.3 0.3 1.0 12.2 1.0 1.0 13.8 0.9 1.4 12.2 0.9 1.7 13.1 1.0 0.9 10.1 0.8 1.4 14.2 0.5 2.2
Moore Spiegel 9.2 0.6 0.9 9.1 0.6 0.9 10.9 0.7 0.9 12.5 0.5 0.8 11.5 0.6 0.8 11.3 0.4 0.9 11.4 0.9 2.6 11.7 0.8 3.6 12.7 0.9 3.6 12.0 0.8 3.3 10.5 0.9 4.6 13.1 0.7 4.0
Multi Chua 12.5 0.5 0.2 12.0 0.6 0.5 11.9 0.6 0.3 13.4 0.4 0.5 13.0 0.4 0.3 13.9 0.4 0.4 12.5 0.5 0.5 12.6 0.5 0.4 13.6 0.5 0.3 13.0 0.5 0.5 10.7 0.5 0.1 13.5 0.5 0.8
Nose Hoover 11.4 0.5 0.5 10.2 0.6 0.4 10.8 0.7 0.6 10.8 0.6 0.3 12.9 0.8 0.6 13.3 0.5 0.6 11.2 0.9 1.6 11.3 0.9 1.4 12.7 0.7 2.2 11.0 0.9 2.0 10.5 0.9 2.3 13.3 0.5 3.1
Pehlivan Wei 4.8 0.2 3.1 4.6 0.1 0.4 6.8 0.4 0.5 9.0 0.3 0.6 10.7 0.3 0.9 12.4 0.5 1.5 11.1 0.9 5.8 12.0 0.6 6.1 12.6 0.8 6.1 12.3 0.8 5.5 10.0 0.9 2.3 12.7 0.4 6.3
Sakarya 3.1 0.3 1.0 2.6 0.2 0.5 5.0 0.3 0.5 8.2 0.3 0.5 10.2 0.7 1.0 12.9 0.8 1.8 10.7 0.9 1.5 12.7 0.9 1.9 11.8 0.8 2.0 12.8 0.8 2.1 12.4 0.7 1.4 14.0 0.5 2.9
San Um Srisuchinwong 5.5 0.1 0.5 5.5 0.2 1.5 10.5 0.8 1.5 12.6 0.4 1.5 13.0 0.5 1.4 13.1 0.6 0.7 11.8 0.8 1.4 13.0 0.7 0.3 12.0 0.6 1.1 11.9 0.7 0.6 10.7 0.5 1.7 12.6 0.5 0.5
Sprott A 4.9 0.1 0.6 9.3 0.5 0.2 8.5 0.4 0.3 7.5 0.3 0.3 11.7 0.3 0.6 12.6 0.5 0.8 10.7 0.9 2.9 11.8 0.7 2.9 11.5 0.8 3.1 11.1 0.7 3.5 9.8 0.9 0.6 12.4 0.6 3.5
Sprott B 2.6 0.2 2.1 9.2 0.5 1.1 11.0 0.6 1.3 13.2 0.5 2.0 13.4 0.6 0.9 13.4 0.7 2.3 12.6 0.8 1.2 13.2 0.8 1.9 12.5 0.8 2.0 12.6 0.7 1.8 13.9 0.8 1.9 12.9 0.6 1.5
Sprott C 2.9 0.2 0.3 8.5 0.6 1.2 12.0 0.3 1.2 13.3 0.4 0.5 12.6 0.4 0.6 11.8 0.5 2.4 12.4 0.7 2.2 13.0 0.6 2.1 11.9 0.6 2.9 11.8 0.6 2.1 13.3 0.7 2.2 13.5 0.5 2.9
Sprott D 3.8 0.1 0.3 7.6 0.2 0.1 5.5 0.1 0.1 5.5 0.2 0.1 12.9 0.4 0.5 13.1 0.5 0.6 12.1 0.7 0.5 12.0 0.8 3.1 12.1 0.7 1.8 12.0 0.6 2.0 12.9 0.7 0.7 13.8 0.5 1.1
Sprott F 4.9 0.3 1.1 10.8 0.4 0.9 8.9 0.5 0.7 12.6 0.3 0.5 13.0 0.5 0.5 13.1 0.4 0.4 12.3 0.9 1.8 12.2 0.6 2.1 13.2 0.8 2.0 12.4 0.7 2.4 11.8 0.9 1.0 13.4 0.6 1.3
Sprott G 2.4 0.0 0.2 4.2 0.1 0.1 6.2 0.2 0.1 7.1 0.2 0.2 7.7 0.3 0.7 12.9 0.3 0.6 12.3 0.8 3.5 12.4 0.7 3.9 12.9 0.8 2.9 11.2 0.9 4.4 11.0 0.9 2.3 14.0 0.5 1.6
Sprott I 3.0 0.2 0.6 8.6 0.5 0.1 10.8 0.1 0.1 7.4 0.3 0.3 10.1 0.3 0.3 13.1 0.4 0.2 12.1 0.9 1.1 11.9 0.8 2.9 9.4 0.9 1.0 13.4 0.9 2.2 10.9 0.9 1.0 14.4 0.6 2.7
Sprott L 2.7 0.2 1.4 9.2 0.3 0.2 8.2 0.1 0.3 10.2 0.2 0.2 11.3 0.2 0.3 13.6 0.4 0.2 12.3 0.9 1.0 12.3 0.9 0.4 13.3 0.9 1.6 13.1 0.9 2.0 12.2 0.8 0.5 14.2 0.5 3.1
Sprott M 3.4 0.2 0.4 10.9 0.5 0.1 12.4 0.4 0.1 12.8 0.3 0.3 13.5 0.3 0.3 14.0 0.8 0.5 12.1 0.8 0.9 11.1 0.7 1.3 13.0 0.8 0.9 12.3 0.8 0.8 10.7 0.9 0.8 14.3 0.6 2.3
Sprott O 3.5 0.2 0.4 9.2 0.2 0.4 6.7 0.1 0.4 8.1 0.2 0.6 11.6 0.4 0.4 13.2 0.3 0.5 11.8 1.0 0.9 12.7 0.8 1.6 11.4 1.0 1.8 12.6 1.0 3.9 10.4 0.9 0.9 14.0 0.7 3.5
Sprott P 5.7 0.2 1.2 9.2 0.6 0.8 10.5 0.4 0.5 11.1 0.8 1.0 13.3 0.4 0.7 13.5 0.6 1.6 9.6 0.8 4.0 7.6 0.6 3.3 8.7 0.8 3.4 9.6 0.8 3.0 10.8 0.8 1.5 10.4 0.5 1.8
Sprott Q 5.8 0.2 1.1 9.5 0.5 0.5 9.4 0.5 0.4 10.0 0.7 0.5 12.9 0.5 1.3 12.6 0.3 1.3 11.1 0.9 0.8 13.0 0.7 1.8 12.5 0.7 0.6 11.5 0.8 1.7 9.7 0.9 0.4 13.2 0.4 4.0
Sprott R 5.1 0.3 0.5 6.9 0.4 0.1 6.7 0.3 0.1 6.5 0.2 0.0 12.5 0.3 0.3 14.2 0.5 0.4 11.0 0.9 1.4 8.8 0.7 1.0 10.6 0.9 1.0 9.4 0.9 0.9 11.4 0.9 1.7 13.8 0.6 2.3
Sprott S 2.9 0.0 0.3 9.6 0.4 0.3 10.7 0.2 0.2 11.7 0.3 0.3 13.0 0.4 0.3 13.3 0.4 0.7 10.2 0.9 0.6 10.2 0.8 1.6 11.2 0.9 1.7 11.0 0.8 2.0 10.7 0.9 1.9 12.8 0.5 4.0
Sprott Torus 3.6 0.2 0.2 6.8 0.6 0.2 12.0 0.5 0.2 13.0 0.5 0.3 14.3 0.7 0.3 14.7 0.6 0.7 7.0 0.9 0.6 12.1 0.4 0.8 7.9 0.9 0.6 12.5 0.7 0.1 12.2 0.9 0.3 14.9 0.5 1.0
Thomas 2.3 0.2 6.7 12.4 0.6 4.1 12.1 0.3 3.6 13.4 0.2 2.5 13.4 0.3 1.9 13.8 0.2 1.1 12.3 0.8 4.9 12.0 0.8 6.6 13.7 0.7 4.9 13.1 0.7 5.5 11.9 0.8 4.9 13.6 0.4 5.4
Thomas Labyrinth 6.5 0.1 5.2 12.5 0.2 4.8 9.8 0.4 5.6 11.8 0.3 3.4 12.2 0.3 4.6 12.5 0.4 4.0 11.5 0.6 4.4 12.3 0.6 4.2 11.5 0.6 4.2 11.5 0.6 3.8 12.5 0.6 3.3 12.5 0.3 2.6
Torus 10.9 0.9 2.3 12.4 0.4 0.1 11.9 0.2 0.1 12.0 0.2 0.2 12.9 0.5 1.0 12.9 0.7 1.5 12.0 1.0 2.1 12.6 0.8 2.3 12.6 1.0 2.0 12.4 1.0 3.1 8.7 0.9 1.7 13.1 0.6 4.6
Tsucs 2 7.2 0.6 1.7 10.9 0.7 1.0 10.3 0.8 1.2 10.0 0.2 1.6 12.0 0.6 1.7 13.3 0.7 2.2 8.5 0.8 2.2 12.3 0.8 2.2 7.5 0.7 2.4 9.4 0.7 2.3 11.5 0.8 2.6 9.8 0.5 2.1
Vallise El Nino 1.8 0.2 0.4 10.1 0.4 0.3 5.5 0.2 0.2 7.4 0.3 0.2 12.7 0.5 0.5 13.4 0.6 1.7 11.6 0.8 1.8 12.5 0.8 1.0 13.6 0.7 1.1 12.4 0.8 1.9 13.6 0.7 1.3 13.7 0.6 2.1
Wang Sun 5.3 0.2 0.7 11.8 0.5 1.2 10.3 0.4 0.4 13.2 0.5 2.0 10.4 0.5 0.4 10.5 0.7 0.7 9.3 0.7 1.9 11.9 0.7 1.8 11.2 0.7 1.8 9.9 0.6 1.7 13.3 0.7 1.5 12.6 0.5 1.0
Winsmi Reduced 5.0 0.3 1.4 8.9 0.4 1.8 11.2 0.4 1.6 13.6 0.4 1.6 13.6 0.5 1.5 14.0 0.7 1.0 11.7 0.9 2.1 11.7 0.8 2.4 11.9 0.9 1.4 10.4 0.8 2.7 11.4 0.9 0.9 14.3 0.6 2.1
Yu Wang 6.4 0.3 3.2 9.8 0.7 3.3 11.3 0.6 3.4 10.3 0.6 3.5 8.1 0.6 3.3 12.8 0.6 3.2 12.3 0.9 7.6 12.0 0.8 5.7 12.8 0.7 6.6 12.8 0.8 7.2 12.0 0.7 5.0 13.7 0.5 4.6
Yu Wang 2 3.4 0.1 0.3 10.0 0.2 0.1 10.6 0.4 0.2 12.7 0.2 0.1 12.2 0.2 0.5 13.8 0.4 0.5 10.8 0.9 1.4 9.4 0.8 2.5 8.6 0.7 1.6 10.0 0.7 1.4 11.6 0.9 1.8 10.7 0.3 1.2
Zhou Chen 10.0 0.4 0.2 10.4 0.4 1.3 13.7 0.5 3.0 12.7 0.4 0.9 12.8 0.3 0.4 12.8 0.4 0.9 13.2 0.5 1.6 11.6 0.5 2.0 13.5 0.5 1.8 9.7 0.5 2.2 12.8 0.6 2.7 11.6 0.4 2.9

### v2

@@
 Here we introduce DynaMix, a novel multivariate ALRNN-based mixture-of-experts architecture pre-trained for DSR, the first DSR model able to generalize zero-shot to out-of-domain DS.
-Just from a provided context signal, without any re-training, DynaMix faithfully forecasts the long-term evolution of novel DS where existing time series (TS) foundation models, like Chronos, fail – at a fraction of the number of parameters and orders of magnitude faster inference times.
+Just from a provided context signal, without any re-training, DynaMix faithfully forecasts the long-term evolution of novel DS where existing time series (TS) foundation models, like Chronos, fail – at a fraction of the number of parameters ($0.1\%$) and orders of magnitude faster inference times.
 DynaMix outperforms TS foundation models in terms of long-term statistics, and often also short-term forecasts, even on real-world time series, like traffic or weather data, typically used for training and evaluating TS models, but not at all part of DynaMix’ training corpus.
@@
 The core features of DynaMix are:
-  • Accurate zero-shot DSR: DynaMix achieves strong generalization across diverse DS, eliminating the need for extensive fine-tuning while maintaining accuracy in reproducing attractor geometry and long-term statistics.
+  • Accurate zero-shot DSR: DynaMix achieves strong generalization across diverse DS, eliminating the need for fine-tuning while maintaining accuracy in reproducing attractor geometry and long-term statistics.
   • Multivariate information transfer: Due to its multivariate architecture, the model efficiently captures dependencies among multiple system dimensions, enabling accurate reconstruction of their coupled dynamics.
-Further, it is neither bound to a specific dimensionality, but can flexibly adapt to other dimensions through specific embeddings, nor to a specific context length.
-  • Computational and parameter efficiency: The model achieves high performance with a very lightweight architecture, enabling orders of magnitude faster inference than other foundation models.
+Further, it is neither bound to a specific dimensionality nor to a specific context length, but can flexibly adapt to other dimensions through specific embeddings.
+  • Computational and parameter efficiency: The model achieves high performance with a very lightweight architecture ($\approx 10$k parameters) and small training corpus ($34$ systems), enabling orders of magnitude faster inference than other foundation models.
   • Interpretable dynamics composition: The model provides insight into the dynamical composition of reconstructed systems, elucidating similarities between different DS.
@@
 One idea is to simply use pretrained LLMs directly as TS forecasters [34, 71, 79, 93, 83], but the inherent differences between textual and many temporal data, such as continuity, present significant challenges [82].
-To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos [2] or Moirai [51], were specifically pretrained on a large corpus of time series data.
+To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos [2] or TimesFM [20], were specifically pretrained on a large corpus of time series data.
 Promising zero-shot forecasting capabilities have also been achieved with alternative architectural designs, such as Tiny Time Mixers [26] or state-space models like Mamba [5].
@@
 3.1 Model architecture
-To enable zero-shot reconstruction of novel DS, here we develop a specific mixture-of-experts (MoE) architecture that can be pretrained across many diverse DS (Fig.
-9), with different experts possibly specializing in different dynamical regimes.
-As a base model for the experts, we leverage a recent parameter-friendly RNN designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; [7]):
+To enable zero-shot reconstruction of novel DS, we develop a specific mixture-of-experts (MoE) architecture that can be pretrained across many diverse DS (Fig.
+10), with different experts possibly specializing in different dynamical regimes.
+As a SOTA DSR base model for the experts, we leverage a recent parameter-friendly RNN, which allows for highly efficient DSR training and is designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; [7]):
 $\bm{z}_{t}=\bm{A}\bm{z}_{t-1}+\bm{W}\Phi^{*}(\bm{z}_{t-1})+\bm{h}\;.$ (1)
@@
 The selection of AL-RNN experts is achieved through a gating network (Fig.
-2), which receives as inputs the (generally multivariate) context time series $\bm{C}=\{\bm{c}_{t}\}\in\mathbb{R}^{N\times T}$ as well as the current latent state $\bm{z}_{t}$.
+2), which receives as inputs the (generally multivariate) context time series $\bm{C}=\{\bm{c}_{t}\}\in\mathbb{R}^{N\times T_{C}}$ as well as the current latent state $\bm{z}_{t}$.
 Both are passed into a state attention mechanism defined as
-$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T}^{\top}\right|^{\top}\bm{1}_{N}}{T_{\text{att}}}\right)\in\mathbb{R}^{T}\;,$ (3)
-where $\bm{D}\in\mathbb{R}^{N\times M}$ is a learnable matrix which maps the latent state into observation space, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{\Sigma})$ exploration noise with a learnable covariance $\bm{\Sigma}$, $T_{\text{att}}$ a learnable temperature parameter, $\bm{1}_{\{T,N\}}$ column vectors of ones of length $T,N$, respectively, and $\sigma(\cdot)$ the softmax returning normalized weights.
+$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T_{C}}^{\top}\right|^{\top}\bm{1}_{N}}{\tau_{\text{att}}}\right)\in\mathbb{R}^{T}\;,$ (3)
+where $\bm{D}\in\mathbb{R}^{N\times M}$ is a learnable matrix which maps the latent state into observation space, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{\Sigma})$ exploration noise with a learnable covariance $\bm{\Sigma}$, $\tau_{\text{att}}$ a learnable temperature parameter, $\bm{1}_{\{T_{C},N\}}$ column vectors of ones of length $T_{C},N$, respectively, and $\sigma(\cdot)$ the softmax returning normalized weights.
 Hence, this mechanism computes attention weights based on some distance between projected latent states and actual context observations.
-At the same time, the context signal $\bm{C}$ is processed by a CNN, yielding temporal features $\tilde{\bm{C}}\in\mathbb{R}^{N\times T}$.
+At the same time, the context signal $\bm{C}$ is processed by a CNN, yielding temporal features $\tilde{\bm{C}}\in\mathbb{R}^{N\times T_{C}}$.
 These features are then weighted with the attention weights $\bm{w}^{att}_{t}$, and together with the current state $\bm{z}_{t}$ further processed by a multi-layer perceptron (MLP) followed by a softmax, yielding a set of expert weights at any time $t$:
-$\bm{w}^{exp}_{t}=\sigma\left(\frac{\text{MLP}(\tilde{\bm{C}}{\bm{w}^{att}_{t}},\bm{z}_{t})}{T_{\text{exp}}}\right)\in\mathbb{R}^{J}\;,$ (4)
-where $T_{\text{exp}}$ is another learnable temperature parameter.
+$\bm{w}^{exp}_{t}=\sigma\left(\frac{\text{MLP}(\tilde{\bm{C}}{\bm{w}^{att}_{t}},\bm{z}_{t})}{\tau_{\text{exp}}}\right)\in\mathbb{R}^{J}\;,$ (4)
+where $\tau_{\text{exp}}$ is another learnable temperature parameter.
 The forward-iterated latent states $\bm{z}_{t+1}^{j}$ of the individual experts $j\in\{1,...,J\}$ are then weighted by $\bm{w}^{exp}_{t}$ to yield the next time step prediction $\bm{z}_{t+1}=\sum_{j=1}^{J}w^{exp}_{j,t}\cdot\bm{z}_{t+1}^{j}$.
@@
 For training, a collection of just $34$ different chaotic and cyclic DS is used [29], see Appx.
-A.2 for full specification.
-From each, multivariate time series $\bm{X}\in\mathbb{R}^{N\times T}$ are simulated ($\approx 1.6$ million in total) and then standardized (dimension-wise), of which the first $T_{C}<T$ column entries are defined as the context signal $\bm{C}=\bm{X}_{1:T_{C}}$.
+A.2 for full specification and Fig.
+10 for illustration.
+From each, multivariate time series $\bm{X}\in\mathbb{R}^{N\times T}$ are simulated ($\approx 6\times 10^{5}$ in total) and then standardized dimension-wise (likewise for the test set), of which the first $T_{C}<T$ column entries are defined as the context signal $\bm{C}=\bm{X}_{1:T_{C}}$.
 The experts are initialized at $t_{0}=T_{C}-\Delta t+1$ and forward-iterated until time $T$, such that they overlap for a period $\Delta t$ with the context.
@@
 As shown in Fig.
-24, STF with optimal $\tau$ is essential for achieving good zero-shot DSR results.
+26, STF with optimal $\tau$ is essential for achieving good zero-shot DSR results.
 Note that STF is only used for training and turned off at test time.
@@
 For assessing agreement in long-term temporal properties, we used the Hellinger distance $D_{H}$ defined across power spectra of the true and model-generated trajectories.
-In addition, a short-term, $n$-step-ahead mean absolute prediction error was evaluated.
-As has been emphasized repeatedly in the statistical and DSR literature [89, 46, 6, 56, 64, 30], prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig.
-22).
+In addition, a short-term, $n$-step-ahead mean absolute prediction error (MAE) was evaluated.
+As has been repeatedly emphasized in the statistical and DSR literature [94, 48, 6, 58, 68, 30], prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig.
+24).
 Empirically observed time series are often just one-dimensional ($N^{*}=1$), but come from an inherently much higher-dimensional underlying DS.
@@
 4.1 Zero-shot DSR
-Figure 3: a) DynaMix zero-shot DSR (red) compared to ground truth (blue) using a $500$-step context (gray) for the Sprott M system.
-b) Zero-shot forecasts (red) for the Selkov DS (true vector field in blue) from different initial conditions outside the context range (gray).
-c) DSR quality as a function of context length.
+Figure 3: a) DynaMix zero-shot DSR (red) compared to ground truth (lightgray) using a $500$-step context (darkgray) for the Sprott M system.
+b) Zero-shot forecasts for the Selkov DS (true vector field in lightgray) from different initial conditions (red) outside the context range (darkgray).
+c) DSR quality as a function of context length for Lorenz-63.
 d) DSR quality as a function of the temporal resolution $\Delta t$ of the context signal.
+Error bands = STD
 Figs.
 1 (top) & 3a present example zero-shot DSRs of new chaotic DS, which were not part of DynaMix’ training domain, from just a provided context signal $\bm{C}$ (see Figs.
-10-14 for further examples).
+11-15 for further examples).
 Despite the inherent challenge of predicting chaotic systems, where small differences in initial conditions lead to rapidly diverging trajectories (Fig.
-22), DynaMix successfully captures the system’s temporal & geometrical behavior, i.e.
+24), DynaMix successfully captures the system’s temporal & geometrical behavior, i.e.
 accurately reconstructs the underlying state space and faithfully recovers the true system’s power spectrum.
@@
 Notably, the model even generalizes to new initial conditions outside the scope of the context data, i.e.
-correctly infers properties of the state space beyond the short context trajectory.
+correctly infers properties of the state space beyond the short context trajectory (see Fig.
+22 for further examples).
+Likewise, although trained only on $3d$ DS, DynaMix also successfully generalizes to higher-dimensional DS, as illustrated for the $6d$ Lorenz-96 system [54] in Fig.
+34 (see Appx.
+A.6 for details on the setup and comparisons to other models).
 As further shown in Fig.
@@
 sampling frequency.
-We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos [2], Mamba4Cast [5], and Tiny Time Mixers [26], using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see sect.
+We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos [2], Mamba4Cast [5], TimesFM [20], and Tiny Time Mixers [26], using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see Sect.
 A.1).
@@
 4a-b, DynaMix clearly and significantly outperforms all of these.
-In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the (strongest in our batch) Chronos-t5 models (t-tests, all $p>0.37$), as shown in Fig.
+In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the (strongest in our batch) Chronos-t5 models, as shown in Fig.
 4c.
 Table LABEL:tab:performance_DSR assembles further results on performance comparisons, while Figs.
-15-20 provide specific examples.
+16-21 provide specific examples.
 Moreover, a significant limitation of particularly the large Chronos models with millions of parameters is their extreme inefficiency in both inference time and computational costs, as compared in Fig.
@@
 In fact, their inference time can even exceed that of a custom-trained model.
-In contrast, DynaMix provides a highly efficient alternative, offering superior DSR performance while using much less parameters and computation time.
-Figure 4: Zero-shot DSR performance across all 54 test set DS for DynaMix and various TS foundation models for context length $T_{C}=2000$ (see Fig.
-21 for results with $T_{C}=512$).
-Median±$\pm$±MAD of a) $D_{stsp}$ (geometrical disagreement), b) $D_{H}$ (temporal disagreement), and c) short-term prediction error.
-Figure 5: Inference times for zero-shot generation of Lorenz-63 DS with $10000$ forecasting steps.
+In contrast, DynaMix provides a highly efficient alternative, offering superior DSR performance while using orders of magnitude less parameters ($\approx 10$k in total) and computation time.
+[…]
+Figure 5: Number of model parameters vs.
+inference times for zero-shot generation of Lorenz-63 DS with $10000$ forecasting steps.
 All models were provided the exact same context and run on the same hardware.
-Note the log-scale on the $x$-axis.
-For comparison, also the time needed for training and inference of a custom-trained AL-RNN is shown in red.
+Note the log-scale on the xx/$y$-axis.
+For comparison, also the time needed for training and inference of a custom-trained AL-RNN is shown (turquoise).
+To put DynaMix’ performance further in context, we also compared it to a couple of custom-trained DSR models, i.e.
+SOTA DSR models explicitly trained on the context data, incl.
+the plain AL-RNN trained by STF [7], Neural ODEs [37, 17], and reservoir computers which are commonly used for DSR [70] (see Appx.
+A.7 for details).
+Despite the custom trained models’ unduly advantage, performing – unlike DynaMix – only in-domain generalization [37], DynaMix keeps up with, and in some cases even outperforms, them (Appx.
+Tables 68).
 4.2 Reasons for the failure of TS foundation models on DSR
@@
 6a, results in poor long-term accuracy (see also Figs.
-15-20).
+16-21).
 This highlights a fundamental limitation: Univariate approaches are inherently insufficient for reconstructing multivariate, interacting DS.
@@
 6c).
-This is illustrated in more detail in Fig.
-6d for Chronos-t5-base, the best of the TS foundation models in our batch.
-It is particularly evident in the power spectrum (Fig.
-6d right), where Chronos’ sharp, narrow peaks wrongly suggest periodicity, while the true spectrum is usually rather broad and smeared out for chaotic systems, as the one exhibited by DynaMix.
-We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent [56], a measure for the exponential divergence rate of trajectories (see Appx.
-A.1).
-For Chronos, the value $\lambda_{Chronos}\approx 0.06$ close to zero confirms the nearly periodic behavior, while DynaMix produces exponentially diverging trajectories with $\lambda_{DynaMix}\approx 1.48$ close to the ground truth value of $\lambda_{Lorenz63}\approx 1.37$.
-Hence, unlike DynaMix, Chronos does not reconstruct the underlying dynamics but just repeats the context.
-On the other hand, if the underlying dynamics truly is cyclic, DynaMix settles into a cyclic pattern as well, exhibiting no intrinsic bias, as illustrated in Fig.
-6e.
-Figure 6: Common problems in TS foundation models (red: forecast, gray: context, blue: ground truth).
+Figure 6: Common problems of TS foundation models (red: forecast, darkgray: context, lightgray: ground truth).
 a) Dimensions are dynamically decoupled.
@@
 e) For comparison, if the underlying behavior is truly cyclic with prominent peaks in the power spectrum, both Chronos and DynaMix are able to capture it.
+This is illustrated in more detail in Fig.
+6d for Chronos-t5-base, the best of the TS foundation models in our batch.
+It is particularly evident in the power spectrum (Fig.
+6d right), where Chronos’ sharp, narrow peaks wrongly suggest periodicity, while the true spectrum is usually rather broad and smeared out for chaotic systems, as the one exhibited by DynaMix.
+We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent [58], a measure for the exponential divergence rate of trajectories (see Appx.
+A.1).
+For Chronos, the value $\lambda_{Chronos}\approx 0.02$ close to zero confirms the nearly periodic behavior, while DynaMix produces exponentially diverging trajectories with $\lambda_{DynaMix}\approx 0.96$ close to the ground truth value of $\lambda_{Lorenz63}\approx 0.87$.
+Hence, unlike DynaMix, Chronos does not reconstruct the underlying dynamics but just repeats the context.
+On the other hand, if the underlying dynamics truly is cyclic, DynaMix settles into a cyclic pattern as well, exhibiting no intrinsic bias, as illustrated in Fig.
+6e.
 4.3 Dynamical similarity
@@
 As shown in Fig.
-7a (bottom), different experts specialize in specific dynamical patterns or regimes.
-Hence, we can utilize the average expert usage $\bm{e}=\frac{1}{T}\sum_{t=1}^{T}\bm{w}^{exp}_{t}$ across the generated time series to construct a similarity metric
-$S=\frac{1}{S_{max}}\left(||\bm{e}_{1}-\bm{e}_{2}||_{1}\right)^{-1}\;,$ (7)
-defined here as the inverse $L_{1}$ norm of the $\bm{e}$ vectors between any two systems, rescaled to the range $[0,1]$.
+7a (bottom), different expert mixtures specialize in specific dynamical patterns or regimes.
+Hence, we can utilize the expert usage $\bm{w}^{exp}$ across the generated time series to construct a similarity metric $S=\frac{1}{1+W_{2}^{2}}\in[0,1]$, based on the squared 2-Wasserstein distance $W_{2}^{2}$ between the standardized expert weights $\bm{w}^{exp}$ of any two systems.
 Figure 7: a) Top: Vector fields of different experts which in combination determine the attractor.
-Bottom plot illustrates the dominant expert in each time window.
-b) Similarity matrix among all $84$ types of attractor dynamics in our dataset, determined by computing $S$ from the reconstructed time series.
-c) Examples of two pairs of attractors classified as similar (top), and one dissimilar pair.
+Bottom plot illustrates the expert usage across time.
+b) Similarity matrix based on $S$ among all $84$ DS in the dataset, sorted via hierarchical clustering.
+Structure in the matrix indicates there are clusters of DS with high dynamical similarity.
+c) Examples of two pairs of attractors classified as similar (top), and one dissimilar pair (bottom).
 4.4 Time series forecasting from a DS perspective
-Figure 8: Time series forecasts for different data domains (gray: context, blue: ground truth, other colors: various model predictions): Forecasts of a) partially ($1$d) observed Lorenz-63 DS, b) electricity transformer temperature data, c) hourly car traffic data with weekly cycle, d) Huawei cloud request data, e) soil temperature development, f) air pressure data, g) human functional magnetic resonance imaging (fMRI) data, h) human electroencephalogram (EEG) data.
+Figure 8: Comparison of DynaMix (red) to strongest competitor in terms of $D_{stsp}$ (other colors) on zero-shot forecasts of various empirical time series (see Fig.
+27-30 for comparisons to all other TS foundation models): Forecasts of a) partially (11d) observed Lorenz-63 DS, b) electricity transformer temperature, c) hourly car traffic data with weekly cycle, d) Huawei cloud requests, e) soil temperature development, f) air pressure, g) human functional magnetic resonance imaging (fMRI), h) human electroencephalogram (EEG).
 Table 1: Performance comparison on empirical time series in terms of geometrical divergence ($D_{\text{stsp}}$), long-term temporal distance ($D_{H}$), and forecast error (MAE).
 Best in red, second-best in blue.
-System DynaMix Chronos-t5-base Mamba4Cast TTM
-$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
-Partially obs. DS 0.11 0.25 0.32 0.11 0.32 0.39 8.14 0.35 0.76 6.54 0.46 0.89
-ETTh1 0.8 0.08 5.97 0.33 0.09 3.22 0.94 0.1 3.63 6.15 0.09 4.64
-Traffic 0.68 0.21 8.42 0.35 0.1 3.16 1.4 0.34 7.68 1.77 0.25 5.25
-Cloud Requests 0.17 0.14 147.28 1.57 0.3 557.35 7.91 0.16 282.95 1.39 0.26 147.54
-Weather (Temp.) 0.59 0.09 2.24 4.72 0.22 4.14 6.96 0.58 9.39 5.39 0.3 2.76
-Weather (Press.) 0.41 0.19 8.85 6.62 0.23 9.05 7.03 0.52 8.6 7.27 0.25 10.03
-Human fMRI 0.63 0.09 0.56 5.77 0.12 0.85 7.99 0.11 1.35 6.66 0.21 0.9
-Human EEG 0.78 0.23 1.09 10.17 0.24 1.1 6.0 0.62 1.12 2.48 0.37 0.92
+System DynaMix Chronos-t5-base Mamba4Cast TTM TimesFM
+$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
+Partially obs. DS 0.02 0.25 0.39 0.02 0.32 0.39 8.35 0.35 0.76 4.76 0.31 0.90 7.30 0.36 0.71
+ETTh1 0.22 0.08 4.92 0.23 0.09 3.22 0.80 0.10 3.63 6.36 0.09 4.64 0.24 0.10 3.04
+Traffic 0.81 0.21 6.93 0.40 0.10 3.16 1.59 0.34 7.68 1.36 0.25 5.25 0.43 0.11 3.66
+Cloud Requests 0.27 0.14 161.58 1.86 0.31 557.35 5.89 0.16 282.95 1.31 0.25 147.54 0.33 0.29 239.31
+Weather (Temp.) 0.66 0.09 2.24 4.78 0.21 4.14 6.86 0.57 9.39 1.28 0.12 1.92 6.86 0.13 3.31
+Weather (Press.) 0.39 0.19 9.03 6.68 0.23 9.05 7.07 0.52 8.60 7.31 0.26 10.03 9.59 0.60 8.41
+Human fMRI 0.17 0.09 0.45 5.78 0.12 0.85 7.97 0.11 1.35 4.87 0.39 1.54 6.19 0.21 1.11
+Human EEG 0.79 0.23 1.07 9.45 0.24 1.10 5.87 0.61 1.12 10.36 0.26 1.25 10.07 0.27 1.32
 We next tested and compared DynaMix on a variety of real-world datasets often used to probe time series models.
 As discussed in Sect.
-3.3, for DynaMix we used a multivariate embedding of the typically $1d$ empirical time series, to better reflect the underlying DS properties.
+3.3, for DynaMix we used an embedding of the typically $1d$ empirical time series, to better reflect the underlying DS properties (positional embedding for all empirical time series, akin to what is done in TS foundation models, and delay embedding for the Lorenz-63).
 Fig.
@@
 A.2 for further details on the datasets).
-Our model seems to particularly excel on human physiological signals, such as functional magnetic resonance imaging (fMRI) or electroencephalographic (EEG) recordings, where all other models essentially produce meaningless dynamics, at best capturing the mean trend (Fig.
+Our model seems to excel particularly on human physiological signals, such as functional magnetic resonance imaging (fMRI) or electroencephalographic (EEG) recordings, where all other models essentially produce meaningless dynamics, at best capturing the mean trend (Fig.
 8g,h).
 Table 1 quantitatively confirms that DynaMix mostly outperforms the other models tested on the DSR measures.
-Surprisingly, it often even outperforms them in terms of the forecasting error MAE (mean-absolute error), although its training corpus consisted purely of the simulated DS shown in Fig.
-9, i.e.
+Surprisingly, it often even outperforms them in terms of the forecasting error MAE, although its training corpus consisted purely of the simulated DS shown in Fig.
+10, i.e.
 did not include any empirical data at all, let alone traffic, cloud usage, electricity, or temperature data!
@@
 Many real-world data, like weather, human fMRI, or EEG data, are known to bear signatures of deterministic chaos [58, 50], potentially at least partly explaining DynaMix’ strongly competitive forecasts, although it has never seen such type of data in training.
+4.5 Which ingredients are most crucial for DynaMix’ performance? Ablation studies
+Which are the components of the DynaMix framework contributing most to its success in zero-shot out-of-domain generalization?
+Training by STF (Fig.
+26), a sufficiently diverse training corpus, a sufficient number of experts (Fig.
+25), the attention mechanism, and context preprocessing by the CNN, all appear vital to its strong performance (Fig.
+9).
+Replacing the final gating step (the MLP) by a simple linear operation and the exact expert model within the class of piecewise linear RNNs, however, had less of an impact.
+Similar results were obtained when the AL-RNN was swapped for the clipped shallow PLRNN [39], and a probabilistic version of the AL-RNN (adding a noise term $\bm{z}_{t}=F(\bm{z}_{t-1})+\bm{\eta},\bm{\eta}\sim\mathcal{N}(0,\Gamma)$) even slightly improved long-term performance, albeit at the cost of some short-term prediction accuracy.
+Using as experts LSTMs [41], vanilla RNNs or reservoir computers [65] instead led to decay in performance, in line with previous results establishing PLRNNs as SOTA DSR models which allow for efficient training by STF [6, 39, 7].
+Hence, it appears that the whole ‘DSR package’ consisting of 1) architectural choices, 2) a training algorithm specialized for DSR, and 3) properties of the DS training corpus, is crucial to DynaMix’ success.
+Figure 9: Ablation studies.
+a) Effect of removing or altering components of the gating network (no MLP, no CNN, no Attention, CNN to linear layer), including too few experts ($J<5$), replacing STF by standard BPTT training, restricting the training corpus to Lorenz-type systems only.
+b) Effect of replacing the AL-RNN by other RNN models.
+Error bars = STD
 5 Conclusions
-Here we introduce the first DSR foundation architecture that achieves zero-shot reconstructions of novel DS, accurately reproducing their long-term statistics, just from a provided context signal without any retraining or fine-tuning.
+Here we introduce the first DSR foundation model that achieves zero-shot reconstructions of novel DS, accurately reproducing their long-term statistics, just from a provided context signal without any retraining or fine-tuning.
 It hugely outperforms current TS foundation models like Chronos in terms of DSR quality, although much more lightweight with only a fraction (0.1%) of the number of parameters.
@@
 Besides DynaMix’ specific DSR architecture, control-theoretically motivated training techniques, and embedding principles, also the fact that it has seen many chaotic systems in training may play a role.
-Most complex DS in nature [15, 82, 57], engineering [74], and society [54, 87] are likely chaotic, featuring an infinite spectrum of potential temporal patterns (unstable periodic orbits; [36]).
+Most complex DS in nature [15, 86, 59], engineering [78], and society [56, 92] are likely chaotic.
+Moreover, many chaotic attractors, like the Lorenz-63, feature a skeleton of infinitely many (unstable) periodic orbits of all possible periods [36], thus inherently expressing a wide spectrum of temporal patterns.
 Limitations
@@
 8b).
-Changes in statistical properties over time, discontinuities, or widely differing time scales in the data observed, impose severe difficulties for zero-shot forecasting, as illustrated in Fig.
-23a,b, which also plague other TS foundation models.
-However, this may partly be amended by explicitly including non-stationary DS or DS evolving on many different time scales, featuring sharp transitions, etc., in the training corpus.
+Changes in statistical properties over time, tipping points, or widely differing time scales in the data, impose severe difficulties for zero-shot forecasting, as illustrated in Fig.
+31a,b, which also plague other TS foundation models.
+This may partly be amended by explicitly including such types of non-stationary and multiscale DS in the training corpus.
+Another potential solution may be adding explicit filtering and decomposition modules, similar as in Auto- [95] or FEDformers [101].
+A proof of concept of this idea is provided in Appx.
+A.5.
 Incorrect embeddings, which may not be immediately obvious, can also lead to forecast failures (Fig.
-23c).
+31c).
+Finally, continuous-time versions of the experts which can deal with sampling at irregular time points would be another fruitful direction.
 Acknowledgements
@@
 where $\tilde{\bm{z}}_{t}=(\bm{x}_{t},\bm{z}_{N+1:M,t})^{T}$, $F_{\theta}$ is the mixture of experts using the context $\bm{C}$, and $\tau=10$ here (see Fig.
-24).
+26).
 Importantly, STF is only used for training the model, applied after calculating the loss, and is turned off at test time.
@@
 $\mathcal{L}_{MSE}(\hat{\bm{X}},\bm{X})=\frac{1}{N\cdot(T-T_{C}+\Delta t)}\sum_{t=T_{C}-\Delta t+1}^{T}\left\lVert\hat{\bm{x}}_{t}-\bm{x}_{t}\right\rVert_{2}^{2}.$ (8)
-To this we add a regularization term meant to encourage the model to explore a wider region of state space based on the context, by enhancing the variance $\bm{\Sigma}$ of the exploration noise in the state attention:
-$\mathcal{L}_{reg}=\lambda\left(1+\sum_{i}\Sigma_{ii}\right)^{-1}\;,$ (10)
-where we chose $\lambda=0.01$.
+To this we add a regularization term meant to encourage the model to explore a wider region of state space based on the context, by enhancing the variance $\bm{\Sigma}$ of exploration noise in the state attention:
+$\mathcal{L}_{reg}=\lambda\frac{1}{N}\sum_{i}\exp{\left(-\left|\Sigma_{ii}\right|/c\right)}\;,$ (9)
+where we chose $\lambda=0.1$ and $c=0.01$.
 Rectified adaptive moment estimation (RADAM) [52] was employed as the optimizer, with $L=50$ batches of $S_{B}=16$ sequences per epoch, each of length $T=550$, and $2000$ epochs in total.
-We used a learning rate exponentially decaying from $\eta_{\text{start}}=2\cdot 10^{-3}$ to $\eta_{\text{end}}=10^{-5}$.
+We used a learning rate exponentially decaying from $\eta_{\text{start}}=5\cdot 10^{-3}$ to $\eta_{\text{end}}=10^{-5}$.
 The context length used in training was set to $T_{C}=500$, and the window of overlap with the model-generated time series to $\Delta t=50$, see Sect.
 3.2.
-Training was performed on a single CPU (18-Core Xeon Gold 6254), and a single epoch took $60$ seconds depending on sequence length and model size.
+Training was performed on a single CPU (18-Core Xeon Gold 6254), and a single epoch took $30$ seconds depending on sequence length and model size.
 At test time, no retraining or fine-tuning is performed, but utilizing the context $\bm{C}$ the model just forward-iterates from the last context time step.
-Hyperparameters for training were partly selected according to previous results in [7, 6], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.01,0.02\}$, $\Delta t=\{0,50,100\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
+Hyperparameters for training were partly selected according to previous results in [7, 6], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.1,0.2\}$, $\Delta t=\{0,50,80\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
 Model details
-We use $J=20$ AL-RNN experts for our model.
-Each expert has a latent dimension of $M=20$, of which $P=7$ are rectified-linear units (ReLUs).
-As suggested in [7], we only use the first $N$ of the $M-P$ linear units for the readout.
+We use $J=10$ AL-RNN experts for our model.
+Each expert has a latent dimension of $M=30$, of which $P=2$ are rectified-linear units (ReLUs).
+As suggested in [7], we only use the first $N$ of the MPM-P linear units for the readout.
 We followed the initialization protocol in [6, 81], drawing $\bm{W}$ from a Gaussian with mean $\bm{0}$ and $\bm{\sigma}=0.01\mathds{1}$, setting $\bm{h}=\mathbf{0}$, and $\bm{A}$ to the diagonal of a normalized positive-definite random matrix.
@@
 where $\bm{L}\in\mathbb{R}^{(M-N)\times N}$ is jointly learned with other model parameters.
-[…]
+The gating network is implemented using a single-layer CNN with three channels and a kernel size of 2, stride of 1, and zero padding, with the identity as activation function.
+The MLP, which takes as inputs the concatenated weighted CNN outputs and current latent state, consists of two layers with ReLU activation.
+We initialize the temperature weights $\tau_{\text{att}}$ and $\tau_{\text{exp}}$ to $0.1$, the covariance matrix $\bm{\Sigma}=0.05\cdot\mathds{1}$, matrix $\bm{D}=\mathcal{I}_{(N\times M)}$, and draw all other CNN and MLP matrices from a Gaussian with standard deviation $0.01$.
+Model hyperparameters were selected according to previous results in [7] and by probing a few relevant parameter settings, namely $J=\{1,3,5,10,20,30\}$ (Fig.
+25), $\text{CNN-layers}=\{1,2\}$, $\text{CNN-kernel-size}=\{2,3,5\}$, $\text{MLP-layers}=\{1,2,3\}$, and different activation functions $\{\mathds{1},\tanh{},\text{ReLU}\}$ for the MLP/CNN.
+We found that an extensive grid search was not necessary to obtain a well-performing model.
 Performance measures
@@
 $D_{\text{stsp}}\approx\sum_{i=1}^{K}\hat{p}_{true;i}\log\frac{\hat{p}_{true;i}}{\hat{p}_{gen;i}}.$ (12)
-We set $m=30$ bins per dimension, following [37]. To ensure a steady-state distribution is reached, long trajectories of $T=10,000$ time steps are sampled.
+We set $m=30$ bins per dimension, following [38].
+For the $6d$ Lorenz-96 we use $m=5$ and for the empirical time series $m=20$.
+To ensure a steady-state distribution is reached, long trajectories of $T=10,000$ time steps are sampled from the DS.
 To evaluate the agreement in long-term temporal dynamics, we compute the Hellinger distance $D_{H}$ between the power spectra of the true and model-generated time series [58, 39], defined as
-[…]
+$H(F(\omega),G(\omega))=\sqrt{1-\int_{-\infty}^{\infty}\sqrt{F(\omega)G(\omega)}d\omega}\;\in[0,1],$ (13)
 where $F(\omega)$ and $G(\omega)$ are the power spectra of the true and generated time series, respectively.
@@
 The aggregated Hellinger distance $D_{H}$ is then computed as the average across all dimension-wise spectral comparisons.
-The $n$-step ahead prediction error to assess short-term forecast quality (with $n=10$ in our evaluation) is given by
-$\text{PE}(n)=\left\|\bm{x}_{T_{C}+n}-\hat{\bm{x}}_{T_{C}+n}\right\|_{1}.$ (15)
-The MAE for evaluating short term predictions on the empirical time series [39] is given by
+For assessing short-term forecast quality aggregated across the whole test set of $54$ DS, we used a normalized $n$-step ahead prediction error (with $n=10$ in our evaluation) as recommended in [40], given by
+$\text{MASE}=\frac{1}{N}\sum_{i=1}^{N}\frac{\dfrac{1}{n}\sum_{t=T_{C}+1}^{T_{C}+n}\left|x_{i,t}-\hat{x}_{i,t}\right|}{\dfrac{1}{T}\sum_{t=T_{C}+1}^{T_{C}+T}\left|x_{i,t}-x_{i,t-1}\right|}$ (14)
+For individual comparisons on single empirical times, the MAE for evaluating short term predictions [40] was used:
 $\text{MAE}=\frac{1}{n}\sum_{t=T_{C}+1}^{T_{C}+n}\left|x_{t}-\hat{x}_{t}\right|\;,$ (15)
-where we chose $n$ according to the data’s temporal scale and resolution ({ETTh1,fMRI}: $n=40$,{traffic, cloud requests, partially obs.
+where we chose $n$ according to the data’s temporal scale and resolution ({ETTh1,fMRI}: n=40n=40,{traffic, cloud requests, partially obs.
 DS}: $n=80$, {human EEG, weather air pressure}: $n=120$, {weather temperature}: $n=200$).
@@
 First, the time series is embedded into a state space using time-delay embedding as in eq.
-5 (here using the PECUZAL algorithm [47]).
-For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity [68]) and $l_{s}=0.2$.
+5 (e.g.
+using the PECUZAL algorithm [49]).
+For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity [72]) and $l_{s}=0.25$.
 The local divergence between initially nearby trajectories is then tracked over a range of time steps $k=0\dots k_{max}$,
 $d_{i}(k)=\|\bm{x}_{i+k}-\bm{x}_{j(i)+k}\|_{2}\;.$ (16)
-The maximum Lyapunov exponent is given by
+An estimate of the maximum Lyapunov exponent is given by
 $d_{i}(k)\approx d_{i}(0)e^{\lambda_{\text{max}}k\Delta t}\;,$ (17)
@@
 Training data
-DynaMix is trained on about $1.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in [29].
-Gaussian noise of $5\%$ of the data standard deviation was added to all dimensions, and time series were standardized to ensure comparable scaling of all data.
+DynaMix is trained on about $0.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in [29].
+Time series were standardized to ensure comparable scaling of all data, and Gaussian noise of $5\%$ of the data standard deviation was added to all dimensions for the ground truth.
 Attractor dynamics were chosen to reflect different types of behavior, see Fig.
-9.
-Figure 9: Training data (blue) and their reconstruction in red.
+10.
+Figure 10: Training data (gray) and their reconstructions in red.
 Test data
 Our test set for DS consists of simulated time series of length $10^{5}$ sampled from $54$ different $3d$ DS collected in [29], which are not part of the training set.
-Furthermore, we evaluated the Selkov 2d DS [73] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
+Furthermore, we evaluated several $2d$ systems: The Selkov system [77] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
 $\frac{dx}{dt}$ $=-x+ay+x^{2}y,$ (19)
 $\frac{dy}{dt}$ $=b-ay-x^{2}y,$
-where we chose $a=0.1$ and $b=0.5$.
-For evaluation of time series forecasting, different real-world data were used.
-The traffic data are hourly recordings of the number of cars passing road junctions (https://www.kaggle.com/datasets/fedesoriano/traffic-prediction-dataset/data).
-The cloud data, also used in [80] to evaluate TS foundation models, are publicly available from Huawei Cloud https://github.com/sir-lab/data-release.
+where we chose $a=0.1$ and $b=0.5$. The Van-der-Pol system [87] describes self-sustaining oscillations in vacuum tubes by
+$\frac{dx}{dt}$ $=y,$ (20)
+$\frac{dy}{dt}$ $=\mu(1-x^{2})y-x,$
+where we chose $\mu=0.5$. Similar to the Van-der-Pol system, the Rayleigh oscillator [61] describes a self-sustained nonlinear oscillator through
+$\frac{dx}{dt}$ $=y,$ (21)
+$\frac{dy}{dt}$ $=\mu(1-\frac{y^{2}}{3})y-x,$
+where we chose $\mu=1.0$. As a higher dimensional DS test case we use the Lorenz-96 system [54] defined by
+$\frac{\text{d}x_{i}}{\text{d}t}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,$ (22)
+with system variables $x_{i}$, i=1,,Ni=1,...,N, and forcing term $F$ (here, $F=8$ and $N=6$, in the chaotic regime).
+Furthermore, cyclic boundary conditions are assumed with $x_{-1}=x_{N-1}$, $x_{0}=x_{N}$, $x_{N+1}=x_{1}$, and the system was solved with integration step $\Delta t=0.08$.
+We further probed forecasting and DSR on different types of real-world data: The traffic data are hourly recordings of the number of cars passing road junctions (https://www.kaggle.com/datasets/fedesoriano/traffic-prediction-dataset/data).
+The cloud data, also used in [84] to evaluate TS foundation models, are publicly available from Huawei Cloud (https://github.com/sir-lab/data-release).
 It consists of function requests from Huawei’s serverless cloud platform.
@@
 For our evaluation we used the standard pipeline as described in https://github.com/amazon-science/chronos-forecasting.
+TimesFM
+TimesFM is another transformer based TS foundation model, with a decoder-only style architecture using input patching [20].
+Its training corpus consists of synthetic as well as real-world time series data, and it exhibits generalization across different time series domains and context lengths.
+Evaluation is performed as in https://github.com/google-research/timesfm.
 Mamba4Cast
@@
 Tiny Time Mixers
-Tiny Time Mixers is another time series foundation model using a non-transformer architecture [26].
+Like Mamba4Cast, Tiny Time Mixers is a time series foundation model not founded on transformers [26].
 It builds upon the TSMixer backbone [25] to which it adds several key innovations such as adaptive patching, diverse resolution sampling, and multi-resolution prefix tuning to improve generalization across datasets with varying temporal resolutions.
@@
 A.4 Further results
-Figure 10: Zero-shot DSR (red) from a $2000$-step context of unseen DS not contained in the training corpus (ground truth in blue).
-Figure 11: Zero-shot DSR of chaotic Lorenz-63 system (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
-Figure 12: Zero-shot DSR of Lorenz-63 system in cyclic regime (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
-Figure 13: Zero-shot DSR of chaotic finance system (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
-Figure 14: Zero-shot DSR of chaotic Genesio Tesi system (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
+Figure 11: Zero-shot DSR (red) from a $2000$-step context of unseen DS not contained in the training corpus (ground truth in gray).
+Figure 12: Zero-shot DSR of chaotic Lorenz-63 system (darkgray: context, lightgray: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
+Figure 13: Zero-shot DSR of Lorenz-63 system in cyclic regime (darkgray: context, lightgray: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
+Figure 14: Zero-shot DSR of chaotic finance system (darkgray: context, lightgray: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
+Figure 15: Zero-shot DSR of chaotic Genesio Tesi system (darkgray: context, lightgray: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
 Figure 16: Comparison of zero-shot forecasting of finance system for DynaMix vs. different TS foundation models.
@@
 Figure 21: Comparison of zero-shot forecasting of Sprott A system for DynaMix vs. different TS foundation models.
+Figure 22: Zero-shot forecasts (red) for the a) Van-der-Pol system. b) Rayleigh oscillator (true vector field in lightgray) from different initial conditions outside the context range (darkgray).
 Figure 23: Zero-shot DSR performance across all 54 test set DS for DynaMix and various TS foundation models as in Fig.
 4, but for context length $T_{C}=512$.
-Detailed values in Table LABEL:tab:performance_DSR_CL512.
+Numerical details in Table LABEL:tab:performance_DSR_CL512.
 Figure 24: Exponential divergence of two initially nearby trajectories (in blue and red) illustrated for the chaotic Lorenz-63 system.
 The prediction error is still sensible on a short time scale ($\text{MAE}=0.09$), but then rapidly increases and breaks down as a suitable metric in the longer-term ($\text{MAE}=1.12$), although both trajectories were drawn from the exact same system with the very same parameters.
+Figure 25: DynaMix’ performance as a function of the number of experts $J$.
+While a minimum number of experts is necessary, performance already plateaus after a surprisingly small number of experts, suggesting there is a lot of room for scaling the model to larger datasets.
+Error bands = STD
+Figure 26: DSR performance for different teacher forcing intervals $\tau$, illustrating an optimal value is essential for successful training, see [58, 6]. Error bands = STD
+Figure 27: Comparison of DynaMix (red) to Chronos (blue) on zero-shot forecasts of various empirical time series: Forecasts of a) partially (11d) observed Lorenz-63 DS, b) electricity transformer temperature data, c) hourly car traffic data with weekly cycle, d) Huawei cloud request data, e) soil temperature development, f) air pressure data, g) human functional magnetic resonance imaging (fMRI) data, h) human electroencephalogram (EEG) data.
+Figure 28: Comparison of DynaMix (red) to Mamba4Cast (green) on zero-shot forecasts of various empirical time series: Forecasts of a) partially (11d) observed Lorenz-63 DS, b) electricity transformer temperature data, c) hourly car traffic data with weekly cycle, d) Huawei cloud request data, e) soil temperature development, f) air pressure data, g) human functional magnetic resonance imaging (fMRI) data, h) human electroencephalogram (EEG) data.
+Figure 29: Comparison of DynaMix (red) to Tiny-Time Mixers (purple) on zero-shot forecasts of various empirical time series: Forecasts of a) partially (11d) observed Lorenz-63 DS, b) electricity transformer temperature data, c) hourly car traffic data with weekly cycle, d) Huawei cloud request data, e) soil temperature development, f) air pressure data, g) human functional magnetic resonance imaging (fMRI) data, h) human electroencephalogram (EEG) data.
+Figure 30: Comparison of DynaMix (red) to TimesFM (orange) on zero-shot forecasts of various empirical time series: ground truth): Forecasts of a) partially (11d) observed Lorenz-63 DS, b) electricity transformer temperature data, c) hourly car traffic data with weekly cycle, d) Huawei cloud request data, e) soil temperature development, f) air pressure data, g) human functional magnetic resonance imaging (fMRI) data, h) human electroencephalogram (EEG) data.
 Figure 31: Failure of proper zero-shot DSR.
@@
 6.
-Figure 24: DSR performance for different teacher forcing intervals $\tau$, illustrating an optimal value is essential for successful training, see [56, 6].
+A.5 Using DynaMix for non-stationary time series
+Similar as in FEDformers [101] and related TS models which use specialized decomposition and filtering blocks for handling non-stationary data, one could add simple preprocessing operations to the DynaMix pipeline to separate out trend or other non-stationary components in the context signal.
+For a simple illustration, here we first apply a Box-Cox transformation to time series $x_{t}$,
+$x_{t}^{(\lambda)}=\begin{cases}\frac{x_{t}^{\lambda}-1}{\lambda},&\text{if }\lambda\neq 0\\ \log(x_{t}),&\text{if }\lambda=0\end{cases}$ (23)
+where $\lambda$ is estimated by maximizing the log-likelihood. Next, trend components of the form
+$f(t;\bm{\theta})=\theta_{1}t^{\theta_{2}}+\theta_{3}\;,$ (24)
+are inferred by least squares estimation and subtracted from the context signal.
+Standard DynaMix is then used to forecast the embedded residual context (see Sect.
+3.3), after which the estimated trend model is added back on.
+We tested this simple setup on the non-stationary Air Passengers dataset containing passenger counts of an airline (https://www.kaggle.com/datasets/chirag19/air-passengers).
+The results in Fig.
+32 illustrate that this is in principle a viable direction, although constituting just a proof-of-concept at this stage.
+Figure 32: Example forecast for the non-stationary Air Passenger data using DynaMix with preprocessing pipeline.
+A.6 Scaling up context dimension
+For reconstructing DS higher-dimensional than the context dimension defined by the architecture and used in training, one can use DynaMix without any modification by employing the delay-embedding theorems [80, 74]: Zero-shot infer the underlying DS from the observed TS and then delay-embed DynaMix’ output into a sufficiently high-dimensional space which assures a diffeomorphism between original and reconstructed attractor.
+This idea is illustrated for the $6d$ Lorenz-96 system in Fig.
+33.
+Here, only the first two dimensions of the simulated Lorenz-96 were provided as context, for which DynaMix then produced long-term forecasts.
+The good geometrical and temporal agreement between delay-embeddings of both the two ground truth and the forcasted TS confirms that DynaMix has correctly inferred the underlying $6d$ DS despite its own $3d$ structure.
+Figure 33: Zero-shot forecasting of partially observed ($2d$) Lorenz-96 system by DynaMix.
+Both context signal (ground truth) and DynaMix’ forecast were subsequently delay-embedded into $6d$, the system’s original dimension.
+The black line indicates the transition between context and forecast.
+Alternatively, if more dimensions are observed than the gating network can handle by design, $N^{*}>N$, and are all to be utilized for DSR and forecasting, then the gating network needs to be modified architecturally such that it can receive $N=N^{*}$ dimensional context signals.
+To allow for this without changing the $3d$ DS training corpus, observations $\bm{x}\in\mathbb{R}^{3}$ were embedded into a higher dimensional $N>3$ space using a nonlinear transformation of the form
+$\tilde{\bm{x}}=f_{emb}(\bm{x})=\left[\bm{x}^{T}\;\tanh(\bm{A}\bm{x})^{T}\right]^{T}\in\mathbb{R}^{N},$ (25)
+where the entries in $\bm{A}\in\mathbb{R}^{E\times 3}$ are chosen randomly from $a_{ij}\sim\mathcal{U}(-1,1)$, yielding an embedding dimension of $N=3+E$.
+DynaMix is then retrained on the embedded and standardized original $34$ DS training set.
+Note that this embedding does not change the nature of the dynamics of the trained-on systems in any way, because $f_{emb}$ is an instantaneous (time-independent) transform that is only applied posthoc (after the DS has been simulated).
+It essentially places trajectories onto a $3d$ manifold embedded within an $N$ dimensional ambient space.
+Note that this yields a general recipe for any desired $N$.
+As shown in Fig.
+34 and Table 2, using this embedding DynaMix is able to generalize to all $6d$ of the Lorenz-96 system, although only trained on $3d$ systems, while all time series foundation models fail.
+Figure 34: Comparison of zero-shot forecasting of $6d$ Lorenz-96 system by DynaMix vs. different TS foundation models. The black line indicates the transition between context and forecast.
+[…]
+Model $D_{\text{stsp}}$ $D_{H}$ MASE
+DynaMix $1.93\pm 0.08$ $0.07\pm 0.00$ $1.02\pm 0.06$
+Chronos-t5-base $12.74\pm 1.18$ $0.37\pm 0.06$ $2.44\pm 1.30$
+Mamba4Cast $13.34\pm 0.36$ $0.70\pm 0.00$ $2.93\pm 0.45$
+TTM $14.54\pm 0.02$ $0.26\pm 0.02$ $4.00\pm 0.49$
+TimesFM $11.00\pm 0.18$ $0.67\pm 0.01$ $4.82\pm 0.56$
+A.7 Comparison of DynaMix to custom-trained DSR models
+To put DynaMix’ zero-shot DSR performance further into context, we compared it to different SOTA DSR models explicitly trained on the respective context data.
+Note that for the custom-trained models – unlike DynaMix – this constitutes a form of in-domain generalization, as the context signal and the forecast come from the same ergodic distribution.
+For these reasons, it is also difficult to establish a fair comparison, since hyper-parameters of the custom-trained models could be arbitrarily fine-tuned on the in-domain data (with often better results as the model size is further increased, indicating a tendency toward over-fitting).
+For the comparisons performed in here, we therefore decided to mainly rely on previously reported hyper-parameter settings, which often were already fine-tuned for evaluation on similar DS, without additional fine-tuning (which we did not do for DynaMix either).
+Our first comparison model is the same AL-RNN used for DynaMix’ experts (with $\approx 500$ trainable parameters), i.e.
+given by eq.
+1.
+Training was performed as in Brenner et al.
+[7], and hyperparameter settings, as collected in Table 3, followed those in the original paper.
+Table 3: Hyperparameter settings of custom AL-RNN.
+Hyperparameter Setting
+$M$ 20
+$P$ 8 (Lorenz-96: 14)
+$\tau$ 16
+$T$ 200
+batch size 16
+$\eta_{\text{start}}$ $10^{-3}$
+$\eta_{\text{end}}$ $10^{-5}$
+epochs 2000
+As another SOTA DSR model we trained Neural-ODEs [17], using the same type of MLP architecture, training setup, and same hyper-parameter settings as in [37], see Table 4.
+With this, the total number of trainable parameters is already more than twice (!) than that used for DynaMix ($\approx 10$k), and exceeded the number of data points ($\leq 4$k) manifold.
+Indeed, results became worse with less parameters, potentially indicating that Neural-ODEs particularly struggle on the short training segments.
+Table 4: Hyperparameter settings of custom Neural-ODEs.
+Hyperparameter Setting
+hidden layer [100, 100, 100]
+activation ReLU
+$T$ 30
+batch size 32
+ODE solver Tsit5
+$\eta_{\text{start}}$ $10^{-3}$
+$\eta_{\text{end}}$ $10^{-5}$
+epochs 100000
+Finally, Reservoir Computers (RCs) are widely used for DSR.
+Here we employed the architecture provided in Patel and Ott [65], trained as in Göring et al.
+[37] and using the same hyper-parameter settings, summarized in Table 5.
+Note that this endows RCs with about $25\times$ more parameters in total, of which $\leq 3$k were trainable, than DynaMix had at its disposal.
+Table 5: Hyperparameter settings of custom Reservoir Computer.
+Hyperparameter Setting
+$M$ 500
+$\rho$ 1.0
+$\alpha$ 0.7
+$\sigma$ 0.2
+$\beta$ 0.5
+With these settings, performance results in Tables 6 and 7 indicate that DynaMix is within the same ballpark as the custom-trained DSR models, in contrast to all the TS foundation models.
+For the empirical time series, Table 8, DynaMix sometimes even seems to have an advantage over custom-trained models, potentially because it makes more efficient use of the relatively short context time series (for a fair comparison, custom-trained models were provided with the same positional embedding as DynaMix; using a standard delay embedding in $3d$ for the custom-trained models actually produced worse results).
+At the same time, DynaMix admits orders of magnitude faster inference times (see Fig.
+5) and performs true out-of-domain generalization without parameter fine-tuning.
+[…]
+Model $D_{\text{stsp}}$ $D_{H}$ MASE
+DynaMix $3.8\pm 1.4$ $0.16\pm 0.06$ $0.35\pm 0.17$
+AL-RNN $3.86\pm 2.48$ $0.08\pm 0.06$ $0.47\pm 0.23$
+Neural-ODE $3.43\pm 2.05$ $0.09\pm 0.04$ $0.18\pm 0.13$
+RC $1.33\pm 0.99$ $0.08\pm 0.07$ $0.86\pm 0.37$
+[…]
+Model $D_{\text{stsp}}$ $D_{H}$ MASE
+DynaMix $1.93\pm 0.08$ $0.07\pm 0.00$ $1.02\pm 0.06$
+AL-RNN $2.12\pm 0.18$ $0.15\pm 0.03$ $0.83\pm 0.29$
+Neural-ODE $76\pm 0.06 0.0$$06\pm 0.01$ $0.34\pm 0.16$
+[…]
+Table 8: Performance comparison on empirical time series in terms of geometrical divergence ($D_{\text{stsp}}$), long-term temporal distance ($D_{H}$), and forecast error (MAE).
+Best in red, second-best in blue.
+System DynaMix AL-RNN RC Neural-ODE
+$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
+ETTh1 0.22 0.08 4.92 0.26 0.08 6.82 0.60 0.75 44.41 0.70 0.09 4.55
+Traffic 0.81 0.21 6.93 0.95 0.13 7.29 0.19 0.70 7.08 0.85 0.17 5.03
+Cloud Requests 0.27 0.14 161.58 0.19 0.20 529.91 0.24 0.19 587.35 0.37 0.30 172.61
+Weather (Temp.) 0.66 0.09 2.24 0.53 0.11 2.08 1.20 0.11 2.50 0.65 0.11 2.23
+Weather (Press.) 0.39 0.19 9.03 0.28 0.41 8.84 0.55 0.45 61.31 1.13 0.34 11.08
+Human fMRI 0.17 0.09 0.45 0.10 0.11 1.33 0.16 0.19 0.95 0.23 0.06 0.78
+Human EEG 0.79 0.23 1.07 0.23 0.44 0.70 1.80 0.55 0.89 1.23 0.26 0.91
 Table 9: Performance comparison for zero-shot inference across all DS in the dataset from [29] using $T_{C}=2000$. Best value in red, second best in blue.
-System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM
-$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
-Atmospheric Regime 9.7 0.6 0.3 9.2 0.4 0.2 10.5 0.4 0.2 12.0 0.5 0.1 13.3 0.5 0.0 13.7 0.5 0.3 11.6 0.9 0.3 12.3 0.8 0.2 12.3 0.9 0.3 11.7 0.7 0.3 11.2 0.8 0.6 13.9 0.4 1.3
-BeerRNN 8.2 0.4 3.6 12.1 0.5 0.8 9.3 0.4 1.3 13.3 0.4 1.6 10.9 0.3 1.3 13.1 0.5 1.5 12.4 0.9 1.7 12.7 0.8 2.1 11.9 0.9 2.0 11.7 0.8 1.9 11.4 0.8 1.1 14.2 0.4 2.5
-Bouali 4.7 0.1 0.7 7.0 0.3 0.5 12.1 0.2 0.6 11.6 0.1 0.6 13.1 0.2 0.3 13.5 0.4 1.5 11.2 0.7 2.2 12.0 0.6 3.4 12.3 0.5 2.2 9.8 0.5 1.5 12.9 0.7 2.6 14.0 0.3 2.5
-Burke Shaw 3.4 0.1 0.2 8.0 0.1 0.3 5.6 0.1 0.2 5.1 0.2 0.1 11.1 0.2 0.4 13.2 0.6 1.0 11.0 0.6 0.9 11.0 0.7 0.5 11.7 0.8 0.2 10.0 0.7 1.1 10.2 0.7 0.6 12.9 0.4 0.8
-Ca Two Plus Quasip. 3.3 0.1 0.3 9.6 0.5 1.2 9.5 0.3 0.3 12.6 0.2 0.9 12.7 0.4 0.6 13.2 0.4 1.6 9.4 0.7 0.8 9.4 0.6 1.5 9.6 0.7 0.8 11.5 0.5 0.9 11.1 0.6 2.0 13.8 0.2 5.2
-Chen Lee 5.1 0.4 0.1 4.2 0.2 0.0 3.6 0.2 0.0 3.9 0.2 0.0 5.3 0.3 0.0 12.0 0.8 0.3 4.5 0.8 0.3 8.8 0.7 1.1 7.7 0.7 0.5 7.2 0.7 0.5 11.3 0.8 0.2 8.9 0.4 0.6
-Chua 3.0 0.1 0.4 12.9 0.2 0.4 12.4 0.4 0.4 13.3 0.2 0.1 13.5 0.4 0.4 13.5 0.3 0.9 11.9 0.6 1.0 11.7 0.4 0.6 11.7 0.4 0.6 12.0 0.4 0.7 12.1 0.4 0.2 13.4 0.3 1.0
-Coullet 0.8 0.0 0.3 0.6 0.0 0.1 6.3 0.0 0.1 3.9 0.1 0.1 13.7 0.3 0.1 14.6 0.3 0.6 11.4 0.7 0.6 12.8 0.7 0.4 11.8 0.6 0.7 10.7 0.6 1.1 13.3 0.6 0.7 15.0 0.2 1.8
-Dadras 3.0 0.2 13.9 10.0 0.5 9.2 9.2 0.5 4.2 11.8 0.4 7.5 11.4 0.3 4.7 12.9 0.8 9.0 10.5 0.8 8.9 9.4 0.7 9.0 11.0 0.6 8.8 8.8 0.5 8.9 13.1 0.8 7.9 12.2 0.4 9.2
-Dequan Li 10.0 0.5 0.1 12.6 0.5 0.1 13.2 0.4 0.1 13.1 0.3 0.1 12.9 0.4 0.5 13.1 0.3 0.4 12.4 0.4 0.4 12.7 0.4 0.3 12.2 0.4 0.3 12.6 0.4 0.6 12.3 0.4 1.1 13.0 0.4 0.4
-Finance 3.0 0.2 0.6 8.5 0.4 0.2 9.4 0.6 0.3 13.4 0.3 0.5 12.2 0.4 0.3 13.6 0.5 0.3 12.0 0.7 1.3 11.7 0.6 2.3 12.4 0.7 1.3 11.8 0.6 1.4 12.6 0.8 2.3 14.0 0.3 2.1
-Genesio Tesi 10.5 0.2 0.3 8.9 0.1 0.1 8.7 0.1 0.0 7.7 0.1 0.0 11.8 0.4 0.2 12.3 0.3 0.3 10.7 0.3 0.2 11.4 0.9 0.8 10.6 0.6 1.1 10.0 0.4 1.6 10.4 0.8 0.4 12.3 0.2 0.7
-Guckenheimer Holmes 7.9 0.0 0.2 10.6 0.4 0.1 11.7 0.2 0.0 11.3 0.4 0.1 13.2 0.7 0.1 12.5 0.2 0.1 10.7 0.6 0.1 10.6 0.6 0.2 9.2 0.6 0.1 10.3 0.6 0.1 10.4 0.5 0.5 11.6 0.3 0.6
-Hadley 8.4 0.3 0.2 10.8 0.5 0.2 9.3 0.6 0.2 11.9 0.4 0.2 12.5 0.4 0.2 12.5 0.3 0.4 10.9 0.7 0.4 10.7 0.7 1.0 9.5 0.6 0.3 10.2 0.4 0.4 11.7 0.7 0.8 12.5 0.3 1.6
-Halvorsen 2.8 0.1 0.2 6.0 0.1 0.2 8.4 0.1 0.2 5.9 0.1 0.3 12.6 0.2 0.3 13.3 0.3 1.3 11.8 0.8 1.1 9.5 0.7 0.8 8.9 0.8 0.4 10.4 0.5 0.4 12.2 0.6 1.3 13.2 0.3 1.5
-Hindmarsh Rose 6.8 0.5 0.6 9.7 0.3 0.1 6.7 0.5 0.3 8.4 0.1 0.1 11.5 0.4 0.7 13.8 0.6 1.0 12.7 0.7 1.3 12.5 0.7 1.5 13.9 0.6 1.1 13.4 0.7 1.1 13.9 0.6 0.7 14.0 0.2 1.6
-Isothermal Chemical 2.5 0.2 0.1 3.7 0.1 0.0 7.9 0.3 0.0 9.7 0.3 0.0 9.0 0.3 0.0 13.8 0.5 0.2 7.9 0.8 0.3 5.9 0.8 0.2 8.4 0.8 0.2 8.5 0.8 0.2 12.8 0.8 0.5 15.6 0.2 2.8
-Itik Banks tumor 9.1 0.2 1.2 10.0 0.6 0.5 8.4 0.5 0.3 10.5 0.4 0.3 12.8 0.3 0.2 12.1 0.7 0.4 9.2 0.7 1.2 9.5 0.7 1.2 9.7 0.8 1.2 9.6 0.7 1.0 13.3 0.9 0.5 13.2 0.3 1.7
-Jerk Circuit 2.7 0.3 0.4 10.1 0.4 1.0 9.7 0.4 0.4 7.5 0.3 0.4 11.1 0.4 0.4 12.3 0.3 0.2 10.8 0.8 1.6 10.9 0.5 2.0 10.5 0.9 1.9 10.1 0.9 1.8 11.1 0.9 1.8 13.9 0.4 2.1
-Laser 6.5 0.3 0.1 11.1 0.4 0.1 10.8 0.2 0.1 12.8 0.2 0.2 13.8 0.5 0.4 14.1 0.5 0.3 11.8 0.7 0.3 11.5 0.7 0.4 11.0 0.7 0.5 11.9 0.7 1.6 13.3 0.6 0.7 13.8 0.2 1.2
-Lorenz 1.1 0.1 0.7 6.6 0.2 0.3 6.0 0.2 0.3 10.0 0.2 0.2 12.0 0.4 0.3 13.0 0.5 0.3 10.9 0.7 0.3 11.3 0.5 0.6 12.7 0.6 0.9 12.8 0.3 1.8 11.4 0.7 0.2 13.0 0.3 1.9
-Lorenz 84 6.9 0.3 4.9 9.5 0.4 5.2 11.0 0.4 4.8 11.5 0.4 3.2 12.2 0.4 2.5 12.9 0.5 4.4 12.9 0.8 2.4 12.9 0.8 2.5 11.5 0.9 2.4 12.3 0.7 2.4 13.4 0.8 3.3 13.4 0.5 2.4
-Lu Chen 3.4 0.5 0.8 6.3 0.4 0.2 5.3 0.3 0.3 8.0 0.1 0.2 7.1 0.1 0.2 4.9 0.2 0.1 11.6 0.9 1.3 11.9 1.0 1.2 11.3 1.0 1.9 11.4 1.0 1.4 9.7 0.8 0.7 13.7 0.4 3.4
-Moore Spiegel 5.9 0.4 1.0 10.4 0.6 1.0 10.0 0.5 0.0 12.5 0.4 0.2 9.9 0.5 0.6 12.1 0.3 0.9 10.5 0.9 2.8 11.4 0.9 1.8 10.1 0.9 2.5 9.8 0.7 3.4 9.4 0.8 4.6 12.5 0.3 3.6
-Multi Chua 7.8 0.3 0.8 9.3 0.5 0.0 9.3 0.2 0.2 7.1 0.4 0.1 13.3 0.2 0.4 14.0 0.3 1.2 11.5 0.6 1.8 12.2 0.4 3.1 9.5 0.5 2.6 10.1 0.4 2.4 10.6 0.4 1.2 14.0 0.3 2.8
-Nose Hoover 10.0 0.3 0.3 9.9 0.5 0.1 9.6 0.6 0.1 11.5 0.3 0.5 12.5 0.7 0.5 13.0 0.5 0.5 10.6 0.9 0.8 10.6 0.9 0.7 10.1 0.7 0.7 10.2 0.5 0.4 10.2 0.7 0.5 12.9 0.4 0.4
-Pehlivan Wei 4.0 0.2 0.3 8.5 0.4 0.4 10.1 0.4 0.3 11.3 0.3 0.3 10.8 0.4 0.3 12.6 0.5 0.7 9.2 0.7 1.0 10.1 0.8 1.0 9.2 0.7 0.6 10.6 0.8 1.3 9.3 0.7 1.0 12.2 0.6 1.9
-Sakarya 1.5 0.2 0.3 5.6 0.5 0.2 5.0 0.3 0.2 7.5 0.2 0.3 11.1 0.4 0.2 12.5 0.4 0.4 10.2 0.9 0.4 10.0 0.6 1.4 7.7 0.9 0.9 9.1 0.4 0.8 13.1 0.8 0.8 12.6 0.4 0.7
-San Um Srisuchinwong 4.1 0.2 0.2 12.9 0.4 1.0 12.7 0.5 1.6 11.4 0.5 1.0 13.3 0.5 1.2 13.0 0.6 1.1 12.7 0.7 0.9 10.4 0.7 0.9 13.4 0.6 1.2 11.7 0.7 1.5 10.7 0.7 1.6 13.1 0.4 2.4
-Sprott A 2.7 0.1 0.3 8.0 0.2 0.3 8.9 0.3 0.3 6.6 0.2 0.2 11.8 0.3 0.3 12.3 0.3 1.1 10.5 0.8 1.0 8.1 0.5 1.6 10.0 0.7 0.7 10.4 0.7 1.0 9.9 0.7 0.8 12.0 0.3 3.0
-Sprott B 2.4 0.2 0.3 4.3 0.2 0.8 5.6 0.4 1.6 11.7 0.5 0.2 12.7 0.4 0.6 13.2 0.7 1.7 11.8 0.8 0.8 11.5 0.6 0.6 12.3 0.8 1.2 10.6 0.7 1.4 12.5 0.8 1.8 12.6 0.4 1.5
-Sprott C 2.3 0.1 0.6 11.6 0.4 1.2 12.4 0.4 0.7 11.9 0.4 1.2 13.3 0.3 0.3 13.1 0.4 2.4 10.0 0.5 4.7 9.2 0.5 3.9 10.7 0.5 5.3 9.3 0.5 4.3 12.8 0.6 2.9 13.3 0.2 2.4
-Sprott D 4.0 0.2 0.3 9.8 0.6 0.1 10.3 0.4 0.1 12.9 0.3 0.2 13.4 0.5 0.4 13.4 0.4 0.1 12.2 0.7 0.2 10.9 0.6 0.2 9.9 0.5 0.6 11.5 0.6 0.6 12.9 0.7 0.2 14.0 0.2 1.4
-Sprott F 3.0 0.2 0.3 9.6 0.5 0.3 9.2 0.5 0.6 12.3 0.3 0.4 12.6 0.4 0.7 12.7 0.4 0.6 9.9 0.7 1.5 10.4 0.7 1.3 10.4 0.6 1.0 12.1 0.5 1.1 11.4 0.8 1.1 13.3 0.3 0.9
-Sprott G 2.3 0.1 0.7 7.0 0.6 0.1 7.7 0.5 0.1 10.0 0.3 0.3 7.2 0.1 0.4 13.0 0.3 0.6 10.9 0.9 1.2 11.4 0.6 1.3 9.2 0.8 1.4 10.8 0.8 1.7 14.1 0.8 0.7 13.1 0.2 1.3
-Sprott I 2.7 0.1 0.5 9.7 0.5 0.4 11.3 0.2 0.4 9.0 0.1 0.4 11.9 0.3 0.5 13.7 0.2 1.1 10.2 0.9 2.1 9.8 0.9 3.4 8.4 0.6 2.9 9.1 0.6 2.8 12.2 0.7 1.9 14.4 0.2 6.2
-Sprott L 3.4 0.1 0.4 8.0 0.5 0.3 6.8 0.1 0.3 8.2 0.2 0.1 11.6 0.3 0.2 13.0 0.2 0.1 10.5 0.9 1.2 10.4 0.8 1.8 11.4 0.9 1.2 12.1 0.7 1.0 10.8 0.8 1.3 14.2 0.3 3.1
-Sprott M 3.8 0.1 0.4 9.8 0.4 0.0 10.5 0.4 0.1 12.4 0.1 0.1 12.9 0.2 0.2 14.0 0.6 0.6 9.8 0.8 1.4 10.3 0.6 1.6 9.9 0.8 1.5 11.8 0.6 1.6 11.6 0.7 1.5 14.3 0.2 3.2
-Sprott O 1.3 0.1 0.9 3.6 0.1 0.3 4.6 0.0 0.5 9.7 0.1 0.4 12.4 0.4 0.4 12.9 0.4 0.9 10.8 1.0 1.3 12.6 0.9 1.3 10.4 0.8 1.6 11.5 0.9 1.1 10.7 0.8 0.6 14.0 0.4 2.0
-Sprott P 5.5 0.2 0.6 10.8 0.3 2.2 11.9 0.3 1.3 12.7 0.3 1.1 12.7 0.3 1.2 13.3 0.3 2.0 7.3 0.5 3.1 10.6 0.6 4.5 8.3 0.6 4.2 10.0 0.6 5.0 11.1 0.7 2.7 12.2 0.3 5.7
-Sprott Q 4.5 0.1 0.5 10.1 0.2 0.2 7.6 0.1 0.2 7.3 0.1 0.3 11.6 0.2 0.2 13.2 0.3 0.8 11.8 0.8 2.5 11.6 0.7 1.1 11.9 0.6 0.5 11.2 0.6 2.8 10.9 0.7 0.7 13.1 0.4 4.3
-Sprott R 3.5 0.3 0.4 9.3 0.8 0.1 6.7 0.5 0.2 11.0 0.5 0.4 12.6 0.5 0.6 13.1 0.4 0.3 10.6 0.8 0.4 10.4 0.5 0.3 9.6 0.7 0.8 10.5 0.7 0.8 11.8 0.8 0.8 13.8 0.3 1.3
-Sprott S 5.3 0.1 0.6 10.1 0.9 1.7 9.9 0.6 0.7 12.3 0.2 0.9 12.6 0.4 0.4 13.1 0.3 1.1 10.6 0.7 0.6 9.7 0.8 3.5 11.0 0.6 1.8 10.4 0.4 3.7 11.8 0.8 1.8 11.6 0.3 4.6
-Sprott Torus 3.2 0.2 0.1 5.4 0.7 0.4 10.4 0.7 0.1 14.3 0.6 0.2 14.5 0.6 0.4 13.8 0.7 0.3 6.2 0.8 2.9 8.3 0.6 5.2 5.3 0.8 5.4 5.8 0.7 4.1 13.6 0.8 1.0 15.5 0.3 2.1
-Thomas 1.4 0.3 2.1 10.6 0.5 3.8 10.1 0.8 3.9 9.9 0.3 3.7 13.9 0.5 2.6 13.9 0.7 2.8 12.6 0.8 1.9 12.6 0.8 2.0 12.8 0.8 2.0 12.9 0.8 2.0 10.7 0.5 3.0 13.3 0.2 2.6
-Thomas Labyrinth 4.4 0.1 1.8 12.5 0.2 2.8 12.5 0.3 2.9 12.4 0.4 3.1 12.5 0.4 2.9 12.5 0.3 3.8 11.4 0.5 2.7 12.5 0.5 2.2 11.9 0.5 2.5 11.7 0.5 2.1 12.3 0.5 3.1 12.4 0.2 1.8
-Torus 11.5 0.9 2.7 9.8 0.3 0.1 9.4 0.2 0.2 10.9 0.3 0.2 12.3 0.3 0.3 12.4 0.5 0.3 12.1 0.9 1.9 10.5 1.0 2.0 11.6 0.9 1.8 11.9 0.5 1.9 9.1 0.8 1.9 12.6 0.5 3.2
-Tsucs 2 5.7 0.6 0.4 11.6 0.3 0.1 10.3 0.5 0.2 10.6 0.3 0.3 11.6 0.6 0.2 13.8 0.7 0.9 10.6 0.7 0.4 11.2 0.8 0.4 8.2 0.7 0.4 10.5 0.8 0.4 11.1 0.7 0.4 8.4 0.4 0.5
-Vallise El Nino 1.2 0.1 0.3 8.4 0.3 0.6 11.2 0.5 1.2 12.5 0.3 1.3 12.7 0.5 1.9 13.5 0.6 3.2 12.3 0.7 2.5 13.4 0.5 2.7 12.3 0.7 3.3 12.6 0.5 1.2 13.9 0.8 2.9 13.7 0.3 3.7
-Wang Sun 2.8 0.2 0.1 13.2 0.5 0.5 12.2 0.5 0.2 11.8 0.5 0.5 11.1 0.4 0.5 14.6 0.7 1.2 13.9 0.7 0.9 10.9 0.7 0.9 10.8 0.6 1.1 10.9 0.7 0.8 13.3 0.7 0.6 11.1 0.4 0.8
-Winsmi Reduced 6.5 0.4 0.2 11.4 0.4 0.1 10.7 0.5 0.1 13.5 0.6 1.1 13.7 0.6 0.1 14.0 0.4 1.3 11.9 0.9 3.7 12.2 0.8 4.0 10.5 0.6 3.6 9.4 0.3 2.3 12.0 0.8 3.2 14.2 0.2 3.3
-Yu Wang 5.5 0.2 0.6 9.8 0.7 0.8 8.4 0.3 0.3 12.4 0.3 0.2 12.4 0.5 0.7 13.0 0.6 0.9 12.4 0.9 3.6 12.5 0.7 3.2 12.8 0.8 3.3 11.9 0.9 3.9 12.5 0.8 2.3 13.5 0.4 2.9
-Yu Wang 2 3.8 0.0 0.1 9.6 0.5 0.1 10.0 0.6 0.1 12.4 0.3 0.2 12.2 0.3 0.1 14.1 0.7 0.2 9.9 0.2 0.2 9.5 0.7 0.4 9.9 0.6 0.6 10.4 0.4 0.3 10.6 0.7 0.6 12.6 0.2 0.4
-Zhou Chen 4.9 0.3 0.2 11.1 0.3 0.5 5.9 0.4 0.2 11.6 0.3 0.9 14.2 0.4 0.7 10.3 0.3 0.8 12.8 0.4 0.7 12.7 0.5 0.9 13.5 0.4 1.0 12.9 0.5 1.2 13.1 0.5 0.6 13.2 0.2 2.1
+System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM TimesFM
+$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
+Atmospheric Regime 10.4 nan 0.0 9.2 0.4 0.0 10.5 0.4 0.0 12.0 0.5 0.0 13.3 0.5 0.0 13.7 0.5 0.1 11.6 0.9 0.1 12.3 0.8 0.2 12.3 0.9 0.1 11.7 0.7 0.2 11.2 0.8 0.7 13.9 0.4 1.6 11.8 0.4 0.1
+BeerRNN 5.2 0.7 0.6 12.1 0.5 0.5 9.3 0.4 0.7 13.3 0.4 0.7 10.9 0.3 0.6 13.1 0.5 0.8 12.4 0.9 1.0 12.7 0.8 1.7 11.9 0.9 1.5 11.7 0.8 1.6 11.4 0.8 1.0 14.2 0.4 1.5 14.0 0.3 0.5
+Bouali 5.9 0.1 0.2 7.0 0.3 0.4 12.1 0.2 0.5 11.6 0.1 0.5 13.1 0.2 0.4 13.5 0.4 2.0 11.2 0.7 2.3 12.0 0.6 3.1 12.3 0.5 2.0 9.8 0.5 2.2 12.9 0.7 4.4 14.0 0.3 5.6 9.7 0.4 1.1
+Burke Shaw 3.8 0.2 0.3 8.0 0.1 0.5 5.6 0.1 0.4 5.1 0.2 0.2 11.1 0.2 0.5 13.2 0.6 1.1 11.0 0.6 1.3 11.0 0.7 0.9 11.7 0.8 0.9 10.0 0.7 1.2 10.2 0.7 1.2 12.9 0.4 2.1 12.4 0.3 0.8
+Ca Two Plus Quasip. 2.8 0.1 0.3 9.6 0.5 1.6 9.5 0.3 0.5 12.6 0.2 1.2 12.7 0.4 1.4 13.2 0.4 1.1 9.4 0.7 2.6 9.4 0.6 1.3 9.6 0.7 1.4 11.5 0.5 2.4 11.1 0.6 4.0 13.8 0.2 4.6 13.4 0.3 1.9
+Chen Lee 3.6 0.3 0.2 4.2 0.2 0.0 3.6 0.2 0.1 3.9 0.2 0.1 5.3 0.3 0.1 12.0 0.8 0.5 4.5 0.8 0.4 8.8 0.7 0.9 7.7 0.7 0.6 7.2 0.7 1.4 11.3 0.8 0.8 8.9 0.4 2.0 4.6 0.2 0.3
+Chua 2.9 0.1 0.8 12.9 0.2 0.6 12.4 0.4 0.8 13.3 0.2 0.2 13.5 0.4 1.0 13.5 0.3 2.1 11.9 0.6 2.4 11.7 0.4 1.3 11.7 0.4 1.8 12.0 0.4 2.2 12.1 0.4 0.8 13.4 0.3 4.0 12.7 0.2 1.0
+Coullet 0.9 0.0 0.2 0.6 0.0 0.3 6.3 0.0 0.2 3.9 0.1 0.2 13.7 0.3 0.2 14.6 0.3 0.8 11.4 0.7 1.1 12.8 0.7 1.5 11.8 0.6 1.1 10.7 0.6 3.0 13.3 0.6 1.5 15.0 0.2 3.5 9.3 0.2 0.5
+Dadras 2.2 0.2 1.3 10.0 0.5 4.2 9.2 0.5 1.9 11.8 0.4 2.3 11.4 0.3 3.5 12.9 0.8 4.1 10.5 0.8 5.4 9.4 0.7 5.0 11.0 0.6 6.8 8.8 0.5 5.6 13.1 0.8 5.0 12.2 0.4 6.0 13.1 0.6 4.5
+Dequan Li 8.5 0.1 0.8 12.6 0.5 0.2 13.2 0.4 0.3 13.1 0.3 0.3 12.9 0.4 1.4 13.1 0.3 1.0 12.4 0.4 0.8 12.7 0.4 0.8 12.2 0.4 0.6 12.6 0.4 1.4 12.3 0.4 3.3 13.0 0.4 1.7 13.0 0.3 1.3
+Finance 3.1 0.1 0.3 8.5 0.4 0.7 9.4 0.6 0.5 13.4 0.3 0.6 12.2 0.4 0.9 13.6 0.5 1.1 12.0 0.7 2.8 11.7 0.6 4.0 12.4 0.7 3.7 11.8 0.6 4.7 12.6 0.8 4.9 14.0 0.3 4.2 13.2 0.5 2.7
+Genesio Tesi 9.8 0.2 0.4 8.9 0.1 0.2 8.7 0.1 0.1 7.7 0.1 0.1 11.8 0.4 0.4 12.3 0.3 0.4 10.7 0.3 0.8 11.4 0.9 1.4 10.6 0.6 2.0 10.0 0.4 2.8 10.4 0.8 1.3 12.3 0.2 1.7 12.2 0.2 1.1
+Guckenheimer Holmes 7.4 0.0 0.2 10.6 0.4 0.3 11.7 0.2 0.3 11.3 0.4 0.4 13.2 0.7 0.6 12.5 0.2 0.6 10.7 0.6 0.7 10.6 0.6 1.0 9.2 0.6 0.4 10.3 0.6 0.8 10.4 0.5 2.0 11.6 0.3 3.2 11.0 0.2 0.7
+Hadley 8.6 0.3 0.3 10.8 0.5 0.2 9.3 0.6 0.4 11.9 0.4 0.4 12.5 0.4 0.6 12.5 0.3 0.9 10.9 0.7 0.6 10.7 0.7 2.0 9.5 0.6 0.7 10.2 0.4 1.8 11.7 0.7 1.7 12.5 0.3 5.4 11.7 0.3 0.4
+Halvorsen 1.8 0.1 0.2 6.0 0.1 0.3 8.4 0.1 0.5 5.9 0.1 0.3 12.6 0.2 0.5 13.3 0.3 1.4 11.8 0.8 1.6 9.5 0.7 2.1 8.9 0.8 1.8 10.4 0.5 2.3 12.2 0.6 3.4 13.2 0.3 2.8 12.8 0.2 0.8
+Hindmarsh Rose 6.1 0.5 1.1 9.7 0.3 0.3 6.7 0.5 0.7 8.4 0.1 0.3 11.5 0.4 0.9 13.8 0.6 1.5 12.7 0.7 1.5 12.5 0.7 2.2 13.9 0.6 1.2 13.4 0.7 2.2 13.9 0.6 2.5 14.0 0.2 2.7 14.2 0.6 1.5
+Isothermal Chemical 2.4 0.2 0.1 3.7 0.1 0.0 7.9 0.3 0.0 9.7 0.3 0.0 9.0 0.3 0.1 13.8 0.5 0.4 7.9 0.8 0.5 5.9 0.8 0.4 8.4 0.8 0.6 8.5 0.8 0.4 12.8 0.8 0.7 15.6 0.2 6.1 4.0 0.4 0.3
+Itik Banks tumor 8.7 0.2 7.0 10.0 0.6 1.7 8.4 0.5 2.4 10.5 0.4 1.5 12.8 0.3 0.8 12.1 0.7 3.2 9.2 0.7 3.6 9.5 0.7 4.6 9.7 0.8 5.1 9.6 0.7 5.1 13.3 0.9 6.4 13.2 0.3 8.2 12.7 0.3 3.8
+Jerk Circuit 3.2 0.4 3.3 10.1 0.4 1.3 9.7 0.4 1.0 7.5 0.3 0.5 11.1 0.4 0.5 12.3 0.3 0.5 10.8 0.8 1.5 10.9 0.5 2.1 10.5 0.9 1.7 10.1 0.9 1.9 11.1 0.9 1.6 13.9 0.4 2.8 12.6 0.3 1.4
+Laser 8.8 0.4 0.2 11.1 0.4 0.1 10.8 0.2 0.1 12.8 0.2 0.3 13.8 0.5 0.6 14.1 0.5 0.5 11.8 0.7 0.4 11.5 0.7 0.5 11.0 0.7 0.8 11.9 0.7 2.7 13.3 0.6 1.2 13.8 0.2 2.2 12.8 0.4 0.4
+Lorenz 1.2 0.1 0.2 6.6 0.2 0.4 6.0 0.2 0.3 10.0 0.2 0.3 12.0 0.4 0.3 13.0 0.5 0.3 10.9 0.7 0.5 11.3 0.5 0.5 12.7 0.6 1.1 12.8 0.3 3.0 11.4 0.7 1.2 13.0 0.3 2.2 13.1 0.5 0.8
+Lorenz 84 6.9 0.4 1.5 9.5 0.4 2.2 11.0 0.4 1.9 11.5 0.4 1.3 12.2 0.4 1.7 12.9 0.5 3.5 12.9 0.8 2.4 12.9 0.8 2.7 11.5 0.9 2.5 12.3 0.7 2.9 13.4 0.8 2.5 13.4 0.5 2.4 13.4 0.5 2.0
+Lu Chen 7.7 0.8 0.2 6.3 0.4 0.2 5.3 0.3 0.2 8.0 0.1 0.1 7.1 0.1 0.1 4.9 0.2 0.1 11.6 0.9 0.7 11.9 1.0 0.8 11.3 1.0 0.9 11.4 1.0 0.8 9.7 0.8 0.9 13.7 0.4 1.5 8.6 0.5 0.4
+Moore Spiegel 4.3 0.2 0.3 10.4 0.6 0.5 10.0 0.5 0.1 12.5 0.4 0.2 9.9 0.5 0.2 12.1 0.3 0.6 10.5 0.9 1.5 11.4 0.9 1.6 10.1 0.9 1.6 9.8 0.7 2.1 9.4 0.8 2.8 12.5 0.3 2.4 12.4 0.3 2.8
+Multi Chua 8.1 0.3 0.8 9.3 0.5 0.3 9.3 0.2 0.5 7.1 0.4 0.4 13.3 0.2 1.1 14.0 0.3 3.9 11.5 0.6 5.8 12.2 0.4 11.5 9.5 0.5 11.3 10.1 0.4 9.0 10.6 0.4 4.8 14.0 0.3 7.6 11.4 0.2 2.0
+Nose Hoover 9.5 0.2 0.1 9.9 0.5 0.1 9.6 0.6 0.1 11.5 0.3 0.5 12.5 0.7 0.8 13.0 0.5 0.5 10.6 0.9 0.6 10.6 0.9 0.6 10.1 0.7 0.8 10.2 0.5 1.0 10.2 0.7 1.1 12.9 0.4 1.6 13.3 0.3 0.9
+Pehlivan Wei 9.2 0.0 0.2 8.5 0.4 0.5 10.1 0.4 0.5 11.3 0.3 0.3 10.8 0.4 0.5 12.6 0.5 0.7 9.2 0.7 1.2 10.1 0.8 1.7 9.2 0.7 1.0 10.6 0.8 2.1 9.3 0.7 1.4 12.2 0.6 3.0 8.2 0.3 1.2
+Sakarya 2.0 0.2 0.7 5.6 0.5 0.3 5.0 0.3 0.3 7.5 0.2 0.3 11.1 0.4 0.4 12.5 0.4 0.4 10.2 0.9 1.0 10.0 0.6 1.6 7.7 0.9 2.2 9.1 0.4 1.3 13.1 0.8 1.5 12.6 0.4 1.2 11.9 0.3 2.3
+San Um Srisuchinwong 3.6 0.2 0.3 12.9 0.4 1.2 12.7 0.5 2.1 11.4 0.5 1.4 13.3 0.5 2.5 13.0 0.6 1.9 12.7 0.7 1.5 10.4 0.7 2.1 13.4 0.6 1.8 11.7 0.7 1.9 10.7 0.7 2.5 13.1 0.4 3.7 13.3 0.7 1.8
+Sprott A 2.7 0.1 0.3 8.0 0.2 0.3 8.9 0.3 0.3 6.6 0.2 0.3 11.8 0.3 0.3 12.3 0.3 0.9 10.5 0.8 0.9 8.1 0.5 1.6 10.0 0.7 1.5 10.4 0.7 1.6 9.9 0.7 2.5 12.0 0.3 5.7 9.8 0.2 1.0
+Sprott B 2.9 0.2 0.4 4.3 0.2 1.1 5.6 0.4 0.9 11.7 0.5 0.7 12.7 0.4 0.8 13.2 0.7 1.6 11.8 0.8 1.3 11.5 0.6 0.8 12.3 0.8 1.2 10.6 0.7 2.9 12.5 0.8 3.0 12.6 0.4 2.1 13.5 0.6 0.9
+Sprott C 2.4 0.1 0.4 11.6 0.4 2.3 12.4 0.4 1.3 11.9 0.4 2.3 13.3 0.3 1.1 13.1 0.4 4.6 10.0 0.5 7.7 9.2 0.5 6.8 10.7 0.5 8.0 9.3 0.5 7.6 12.8 0.6 5.2 13.3 0.2 6.2 13.4 0.2 4.5
+Sprott D 3.0 0.1 0.4 9.8 0.6 0.3 10.3 0.4 0.3 12.9 0.3 0.4 13.4 0.5 0.7 13.4 0.4 0.3 12.2 0.7 0.8 10.9 0.6 0.6 9.9 0.5 1.2 11.5 0.6 0.9 12.9 0.7 1.3 14.0 0.2 4.0 7.9 0.5 0.9
+Sprott F 5.0 0.3 0.7 9.6 0.5 0.2 9.2 0.5 0.6 12.3 0.3 0.4 12.6 0.4 1.0 12.7 0.4 0.7 9.9 0.7 1.3 10.4 0.7 1.5 10.4 0.6 0.9 12.1 0.5 1.6 11.4 0.8 1.6 13.3 0.3 2.8 13.0 0.6 1.7
+Sprott G 3.8 0.1 0.3 7.0 0.6 0.2 7.7 0.5 0.2 10.0 0.3 0.4 7.2 0.1 0.6 13.0 0.3 0.6 10.9 0.9 1.9 11.4 0.6 1.9 9.2 0.8 1.8 10.8 0.8 2.4 14.1 0.8 1.4 13.1 0.2 1.9 9.5 0.4 1.4
+Sprott I 2.7 0.1 0.5 9.7 0.5 0.7 11.3 0.2 0.6 9.0 0.1 0.5 11.9 0.3 0.6 13.7 0.2 1.0 10.2 0.9 3.5 9.8 0.9 4.9 8.4 0.6 4.6 9.1 0.6 4.1 12.2 0.7 2.0 14.4 0.2 8.8 8.3 0.4 1.0
+Sprott L 3.6 0.1 0.6 8.0 0.5 0.3 6.8 0.1 0.3 8.2 0.2 0.1 11.6 0.3 0.4 13.0 0.2 0.3 10.5 0.9 1.8 10.4 0.8 2.8 11.4 0.9 1.8 12.1 0.7 1.6 10.8 0.8 1.3 14.2 0.3 2.9 8.1 0.4 0.2
+Sprott M 3.6 0.1 0.2 9.8 0.4 0.2 10.5 0.4 0.5 12.4 0.1 0.3 12.9 0.2 0.3 14.0 0.6 0.9 9.8 0.8 3.7 10.3 0.6 3.7 9.9 0.8 3.3 11.8 0.6 3.1 11.6 0.7 2.9 14.3 0.2 6.5 13.4 0.4 1.6
+Sprott O 2.0 0.1 0.4 3.6 0.1 0.4 4.6 0.0 0.4 9.7 0.1 0.5 12.4 0.4 0.6 12.9 0.4 1.3 10.8 1.0 1.6 12.6 0.9 2.6 10.4 0.8 1.8 11.5 0.9 1.4 10.7 0.8 1.6 14.0 0.4 3.7 9.5 0.5 0.7
+Sprott P 3.8 0.2 1.4 10.8 0.3 3.9 11.9 0.3 2.6 12.7 0.3 2.6 12.7 0.3 2.1 13.3 0.3 3.8 7.3 0.5 5.7 10.6 0.6 9.6 8.3 0.6 8.7 10.0 0.6 10.7 11.1 0.7 6.3 12.2 0.3 10.2 11.8 0.3 4.4
+Sprott Q 4.5 0.1 0.2 10.1 0.2 0.3 7.6 0.1 0.4 7.3 0.1 0.5 11.6 0.2 0.4 13.2 0.3 1.3 11.8 0.8 3.6 11.6 0.7 1.9 11.9 0.6 1.0 11.2 0.6 4.1 10.9 0.7 1.1 13.1 0.4 7.5 8.8 0.4 0.4
+Sprott R 4.3 0.3 0.2 9.3 0.8 0.2 6.7 0.5 0.3 11.0 0.5 0.5 12.6 0.5 0.6 13.1 0.4 0.4 10.6 0.8 0.5 10.4 0.5 1.2 9.6 0.7 2.1 10.5 0.7 1.8 11.8 0.8 1.3 13.8 0.3 2.5 11.3 0.3 0.9
+Sprott S 4.2 0.1 0.3 10.1 0.9 2.2 9.9 0.6 1.1 12.3 0.2 1.2 12.6 0.4 0.6 13.1 0.3 1.5 10.6 0.7 1.2 9.7 0.8 6.1 11.0 0.6 3.1 10.4 0.4 7.2 11.8 0.8 4.1 11.6 0.3 10.2 12.8 0.2 3.7
+Sprott Torus 2.2 0.2 0.2 5.4 0.7 0.4 10.4 0.7 0.3 14.3 0.6 0.3 14.5 0.6 0.6 13.8 0.7 0.4 6.2 0.8 4.2 8.3 0.6 8.3 5.3 0.8 6.3 5.8 0.7 6.3 13.6 0.8 2.4 15.5 0.3 4.8 14.3 0.5 1.0
+Thomas 1.6 0.3 4.8 10.6 0.5 6.0 10.1 0.8 6.0 9.9 0.3 5.7 13.9 0.5 4.7 13.9 0.7 5.6 12.6 0.8 3.4 12.6 0.8 1.9 12.8 0.8 2.8 12.9 0.8 2.5 10.7 0.5 6.1 13.3 0.2 3.1 13.8 0.6 3.3
+Thomas Labyrinth 4.1 0.1 2.4 12.5 0.2 2.6 12.5 0.3 2.8 12.4 0.4 3.0 12.5 0.4 2.8 12.5 0.3 3.8 11.4 0.5 3.2 12.5 0.5 2.6 11.9 0.5 3.0 11.7 0.5 3.1 12.3 0.5 3.6 12.4 0.2 2.0 12.6 0.5 2.5
+Torus 11.9 0.9 1.1 9.8 0.3 0.1 9.4 0.2 0.1 10.9 0.3 0.1 12.3 0.3 0.2 12.4 0.5 0.3 12.1 0.9 0.8 10.5 1.0 0.8 11.6 0.9 0.7 11.9 0.5 0.7 9.1 0.8 0.5 12.6 0.5 1.1 10.9 0.5 0.3
+Tsucs 2 3.3 0.1 0.6 11.6 0.3 0.4 10.3 0.5 0.4 10.6 0.3 0.3 11.6 0.6 0.4 13.8 0.7 1.0 10.6 0.7 1.0 11.2 0.8 0.9 8.2 0.7 1.0 10.5 0.8 0.9 11.1 0.7 0.8 8.4 0.4 1.2 10.3 0.5 0.4
+Vallise El Nino 1.1 0.1 0.5 8.4 0.3 0.7 11.2 0.5 1.0 12.5 0.3 1.8 12.7 0.5 1.9 13.5 0.6 3.9 12.3 0.7 1.4 13.4 0.5 1.8 12.3 0.7 2.8 12.6 0.5 2.3 13.9 0.8 3.0 13.7 0.3 3.8 13.2 0.3 2.3
+Wang Sun 4.9 nan 0.1 13.2 0.5 0.5 12.2 0.5 0.5 11.8 0.5 0.6 11.1 0.4 0.8 14.6 0.7 1.2 13.9 0.7 1.5 10.9 0.7 1.7 10.8 0.6 1.8 10.9 0.7 1.6 13.3 0.7 1.6 11.1 0.4 1.0 12.3 0.5 0.9
+Winsmi Reduced 8.3 0.5 0.6 11.4 0.4 0.1 10.7 0.5 0.1 13.5 0.6 0.4 13.7 0.6 0.2 14.0 0.4 0.8 11.9 0.9 3.5 12.2 0.8 3.8 10.5 0.6 3.3 9.4 0.3 1.9 12.0 0.8 3.3 14.2 0.2 4.7 13.6 0.2 1.2
+Yu Wang 4.9 0.3 0.6 9.8 0.7 0.6 8.4 0.3 0.3 12.4 0.3 0.3 12.4 0.5 0.7 13.0 0.6 0.8 12.4 0.9 3.7 12.5 0.7 4.4 12.8 0.8 4.4 11.9 0.9 3.3 12.5 0.8 3.4 13.5 0.4 3.4 13.2 0.5 3.0
+Yu Wang 2 2.7 0.1 0.3 9.6 0.5 0.2 10.0 0.6 0.3 12.4 0.3 0.5 12.2 0.3 0.4 14.1 0.7 0.5 9.9 0.2 0.5 9.5 0.7 1.5 9.9 0.6 2.3 10.4 0.4 1.3 10.6 0.7 2.9 12.6 0.2 2.6 10.6 0.2 1.7
+Zhou Chen 4.6 nan 0.4 11.1 0.3 1.0 5.9 0.4 0.5 11.6 0.3 1.8 14.2 0.4 1.5 10.3 0.3 1.8 12.8 0.4 1.5 12.7 0.5 2.1 13.5 0.4 1.9 12.9 0.5 2.6 13.1 0.5 3.7 13.2 0.2 5.1 14.9 0.4 1.3
 Table 10: Performance comparison for zero-shot inference across all DS in the dataset from [29] using $T_{C}=512$. Best value in red, second best in blue.
-System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM
-$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
-Atmospheric Regime 8.4 0.6 2.8 9.7 0.5 0.4 11.1 0.4 0.6 10.5 0.6 0.6 10.1 0.5 1.3 12.0 0.4 0.5 11.9 0.9 3.6 12.0 0.9 3.8 13.5 0.9 2.7 12.7 0.8 4.9 11.3 0.9 1.2 14.2 0.6 3.6
-BeerRNN 11.7 0.8 4.7 8.0 0.5 0.5 3.3 0.2 0.4 9.5 0.2 0.2 7.3 0.4 0.2 11.9 0.4 0.3 12.3 0.9 1.4 12.5 0.9 2.2 12.1 0.9 1.8 12.6 0.8 1.4 13.0 0.9 1.9 14.1 0.6 4.1
-Bouali 4.1 0.1 0.2 8.3 0.5 0.0 7.3 0.3 0.1 8.6 0.1 0.0 11.1 0.3 0.1 13.2 0.5 0.1 10.6 0.8 0.9 9.6 0.7 1.1 12.2 0.7 1.2 12.3 0.8 0.8 12.9 0.8 1.5 13.4 0.4 1.1
-Burke Shaw 4.1 0.2 0.6 10.8 0.5 1.5 9.0 0.5 2.4 13.0 0.4 2.0 13.2 0.4 1.9 13.3 0.6 1.8 11.5 0.8 0.6 12.1 0.6 1.0 12.2 0.7 1.2 10.5 0.7 2.9 12.2 0.9 1.3 13.7 0.5 2.6
-Ca Two Plus Quasip. 2.3 0.1 0.1 7.2 0.1 0.1 8.3 0.1 0.1 9.2 0.4 0.1 13.1 0.4 0.1 13.2 0.4 0.2 11.3 0.9 0.8 11.6 0.8 1.2 12.8 0.6 0.3 8.9 0.6 1.6 11.1 0.8 1.6 13.8 0.7 1.6
-Chen Lee 3.9 0.4 0.3 4.9 0.3 0.9 4.0 0.5 1.6 4.8 0.3 0.4 4.6 0.3 1.0 8.9 0.7 1.3 7.2 0.9 3.6 7.9 0.7 2.1 8.0 0.7 1.7 4.8 0.5 1.4 11.0 0.9 1.2 6.4 0.5 4.0
-Chua 5.9 0.1 0.1 7.4 0.2 0.2 5.9 0.2 0.2 7.6 0.1 0.2 11.7 0.2 0.1 13.1 0.3 0.4 11.8 0.6 1.0 12.2 0.5 1.3 11.7 0.6 0.6 12.0 0.6 1.4 11.3 0.5 1.5 13.4 0.4 0.8
-Coullet 1.2 0.0 0.4 4.9 0.0 0.2 0.7 0.0 0.1 2.8 0.0 0.1 7.5 0.2 0.0 14.6 0.4 0.3 13.0 0.8 1.7 11.2 0.8 2.4 11.8 0.8 2.3 11.6 0.7 2.4 12.3 0.8 1.7 14.9 0.5 1.9
-Dadras 4.1 0.2 1.5 7.7 0.6 0.9 4.7 0.5 1.2 13.1 0.5 0.5 12.1 0.4 0.3 10.2 0.7 1.0 12.7 0.8 2.6 12.9 0.7 2.4 13.0 0.8 2.7 8.8 0.8 4.2 11.9 0.8 2.5 12.8 0.4 2.8
-Dequan Li 12.7 0.5 0.3 11.1 0.3 0.2 11.0 0.4 0.4 12.1 0.3 0.3 12.1 0.2 0.5 13.0 0.3 1.5 11.7 0.4 1.1 11.3 0.4 1.0 9.6 0.4 0.7 11.0 0.4 0.4 10.7 0.4 0.6 13.5 0.4 2.5
-Finance 3.2 0.1 0.4 8.5 0.6 0.9 8.6 0.4 0.7 9.1 0.4 0.3 12.5 0.3 0.1 13.8 0.4 1.4 11.6 0.9 1.8 10.7 0.8 2.0 13.4 0.8 1.5 11.8 0.8 1.1 11.1 0.9 0.8 13.7 0.5 1.3
-Genesio Tesi 8.2 0.2 0.3 7.9 0.4 0.7 8.5 0.2 0.7 7.7 0.2 0.7 9.4 0.2 0.8 13.0 0.3 1.6 10.4 0.7 0.7 10.6 0.7 0.8 10.3 0.8 1.0 9.6 0.8 1.5 11.0 0.9 0.9 12.6 0.6 4.1
-Guckenheimer Holmes 9.6 0.1 0.2 11.8 0.1 0.4 12.0 0.1 0.4 12.6 0.3 0.2 12.5 0.4 0.6 12.1 0.4 0.2 9.6 0.7 0.2 9.6 0.7 0.4 9.4 0.6 0.2 11.8 0.6 0.4 10.2 0.7 0.5 12.7 0.5 0.6
-Hadley 11.3 0.4 0.6 10.7 0.5 2.3 9.7 0.5 1.3 12.5 0.3 1.1 12.6 0.6 2.1 12.6 0.4 3.2 11.4 0.8 3.2 11.9 0.7 3.3 11.9 0.6 2.9 11.5 0.8 3.4 11.4 0.8 3.5 12.5 0.5 3.6
-Halvorsen 4.3 0.3 1.0 9.7 0.6 0.3 8.2 0.2 0.8 7.5 0.2 0.8 13.1 0.2 0.6 13.3 0.4 1.3 10.0 0.8 4.7 10.3 0.7 4.4 9.7 0.8 4.4 10.9 0.7 4.3 11.3 0.9 3.4 13.4 0.5 3.5
-Hindmarsh Rose 6.5 0.4 0.5 9.8 0.2 0.1 4.9 0.1 0.0 9.6 0.2 0.0 12.7 0.6 0.2 13.7 0.6 0.5 12.9 0.7 1.1 14.3 0.8 1.3 12.9 0.8 1.1 14.0 0.8 1.4 13.2 0.8 0.6 14.6 0.4 1.1
-Isothermal Chemical 2.0 0.2 0.3 2.5 0.1 0.2 5.4 0.1 0.1 8.3 0.2 0.1 7.0 0.4 0.1 11.6 0.5 0.2 11.7 0.9 0.6 11.7 0.9 1.0 10.8 0.8 0.8 12.5 0.9 0.8 10.1 0.9 0.9 15.6 0.6 2.5
-Itik Banks tumor 5.7 0.2 0.1 8.2 0.5 0.1 10.8 0.6 0.2 9.8 0.5 0.1 11.8 0.6 0.2 13.3 0.6 0.1 10.8 0.6 0.9 11.2 0.8 1.7 11.9 0.9 0.9 12.0 0.8 2.4 12.0 0.9 1.4 12.9 0.7 1.5
-Jerk Circuit 4.3 0.3 0.4 8.8 0.6 1.1 9.1 0.4 0.8 10.6 0.3 1.0 11.7 0.3 0.2 9.8 0.6 1.1 12.5 0.9 1.9 13.5 0.9 1.6 12.9 0.9 0.7 13.1 1.0 1.5 10.3 0.8 0.5 14.1 0.5 3.1
-Laser 4.5 0.2 0.5 7.5 0.5 5.0 8.2 0.3 1.9 11.5 0.2 0.6 12.8 0.2 0.7 13.9 0.6 1.3 12.6 0.7 2.4 12.8 0.5 1.7 13.0 0.7 1.3 12.8 0.7 2.7 11.8 0.8 2.4 13.3 0.5 2.2
-Lorenz 3.4 0.2 0.6 9.8 0.6 0.5 8.5 0.3 0.6 13.0 0.4 0.7 12.6 0.7 0.8 13.3 0.6 0.3 12.8 0.8 1.0 13.0 0.7 1.9 13.2 0.6 1.4 13.0 0.6 1.3 11.8 0.8 1.3 13.4 0.5 1.1
-Lorenz 84 7.9 0.2 1.6 10.7 0.5 3.3 11.0 0.4 4.4 12.6 0.4 4.8 11.0 0.4 0.8 13.2 0.6 1.8 11.8 0.9 2.0 12.2 0.9 2.6 11.5 0.9 2.4 12.3 0.9 2.5 11.2 0.9 2.2 13.4 0.6 2.9
-Lu Chen 9.5 0.8 1.4 5.3 0.2 1.0 9.0 0.3 1.3 8.9 0.2 0.2 9.7 0.3 0.8 8.3 0.3 1.0 12.2 1.0 1.0 13.8 0.9 1.4 12.2 0.9 1.7 13.1 1.0 0.9 10.1 0.8 1.4 14.2 0.5 2.2
-Moore Spiegel 9.2 0.6 0.9 9.1 0.6 0.9 10.9 0.7 0.9 12.5 0.5 0.8 11.5 0.6 0.8 11.3 0.4 0.9 11.4 0.9 2.6 11.7 0.8 3.6 12.7 0.9 3.6 12.0 0.8 3.3 10.5 0.9 4.6 13.1 0.7 4.0
-Multi Chua 12.5 0.5 0.2 12.0 0.6 0.5 11.9 0.6 0.3 13.4 0.4 0.5 13.0 0.4 0.3 13.9 0.4 0.4 12.5 0.5 0.5 12.6 0.5 0.4 13.6 0.5 0.3 13.0 0.5 0.5 10.7 0.5 0.1 13.5 0.5 0.8
-Nose Hoover 11.4 0.5 0.5 10.2 0.6 0.4 10.8 0.7 0.6 10.8 0.6 0.3 12.9 0.8 0.6 13.3 0.5 0.6 11.2 0.9 1.6 11.3 0.9 1.4 12.7 0.7 2.2 11.0 0.9 2.0 10.5 0.9 2.3 13.3 0.5 3.1
-Pehlivan Wei 4.8 0.2 3.1 4.6 0.1 0.4 6.8 0.4 0.5 9.0 0.3 0.6 10.7 0.3 0.9 12.4 0.5 1.5 11.1 0.9 5.8 12.0 0.6 6.1 12.6 0.8 6.1 12.3 0.8 5.5 10.0 0.9 2.3 12.7 0.4 6.3
-Sakarya 3.1 0.3 1.0 2.6 0.2 0.5 5.0 0.3 0.5 8.2 0.3 0.5 10.2 0.7 1.0 12.9 0.8 1.8 10.7 0.9 1.5 12.7 0.9 1.9 11.8 0.8 2.0 12.8 0.8 2.1 12.4 0.7 1.4 14.0 0.5 2.9
-San Um Srisuchinwong 5.5 0.1 0.5 5.5 0.2 1.5 10.5 0.8 1.5 12.6 0.4 1.5 13.0 0.5 1.4 13.1 0.6 0.7 11.8 0.8 1.4 13.0 0.7 0.3 12.0 0.6 1.1 11.9 0.7 0.6 10.7 0.5 1.7 12.6 0.5 0.5
-Sprott A 4.9 0.1 0.6 9.3 0.5 0.2 8.5 0.4 0.3 7.5 0.3 0.3 11.7 0.3 0.6 12.6 0.5 0.8 10.7 0.9 2.9 11.8 0.7 2.9 11.5 0.8 3.1 11.1 0.7 3.5 9.8 0.9 0.6 12.4 0.6 3.5
-Sprott B 2.6 0.2 2.1 9.2 0.5 1.1 11.0 0.6 1.3 13.2 0.5 2.0 13.4 0.6 0.9 13.4 0.7 2.3 12.6 0.8 1.2 13.2 0.8 1.9 12.5 0.8 2.0 12.6 0.7 1.8 13.9 0.8 1.9 12.9 0.6 1.5
-Sprott C 2.9 0.2 0.3 8.5 0.6 1.2 12.0 0.3 1.2 13.3 0.4 0.5 12.6 0.4 0.6 11.8 0.5 2.4 12.4 0.7 2.2 13.0 0.6 2.1 11.9 0.6 2.9 11.8 0.6 2.1 13.3 0.7 2.2 13.5 0.5 2.9
-Sprott D 3.8 0.1 0.3 7.6 0.2 0.1 5.5 0.1 0.1 5.5 0.2 0.1 12.9 0.4 0.5 13.1 0.5 0.6 12.1 0.7 0.5 12.0 0.8 3.1 12.1 0.7 1.8 12.0 0.6 2.0 12.9 0.7 0.7 13.8 0.5 1.1
-Sprott F 4.9 0.3 1.1 10.8 0.4 0.9 8.9 0.5 0.7 12.6 0.3 0.5 13.0 0.5 0.5 13.1 0.4 0.4 12.3 0.9 1.8 12.2 0.6 2.1 13.2 0.8 2.0 12.4 0.7 2.4 11.8 0.9 1.0 13.4 0.6 1.3
-Sprott G 2.4 0.0 0.2 4.2 0.1 0.1 6.2 0.2 0.1 7.1 0.2 0.2 7.7 0.3 0.7 12.9 0.3 0.6 12.3 0.8 3.5 12.4 0.7 3.9 12.9 0.8 2.9 11.2 0.9 4.4 11.0 0.9 2.3 14.0 0.5 1.6
-Sprott I 3.0 0.2 0.6 8.6 0.5 0.1 10.8 0.1 0.1 7.4 0.3 0.3 10.1 0.3 0.3 13.1 0.4 0.2 12.1 0.9 1.1 11.9 0.8 2.9 9.4 0.9 1.0 13.4 0.9 2.2 10.9 0.9 1.0 14.4 0.6 2.7
-Sprott L 2.7 0.2 1.4 9.2 0.3 0.2 8.2 0.1 0.3 10.2 0.2 0.2 11.3 0.2 0.3 13.6 0.4 0.2 12.3 0.9 1.0 12.3 0.9 0.4 13.3 0.9 1.6 13.1 0.9 2.0 12.2 0.8 0.5 14.2 0.5 3.1
-Sprott M 3.4 0.2 0.4 10.9 0.5 0.1 12.4 0.4 0.1 12.8 0.3 0.3 13.5 0.3 0.3 14.0 0.8 0.5 12.1 0.8 0.9 11.1 0.7 1.3 13.0 0.8 0.9 12.3 0.8 0.8 10.7 0.9 0.8 14.3 0.6 2.3
-Sprott O 3.5 0.2 0.4 9.2 0.2 0.4 6.7 0.1 0.4 8.1 0.2 0.6 11.6 0.4 0.4 13.2 0.3 0.5 11.8 1.0 0.9 12.7 0.8 1.6 11.4 1.0 1.8 12.6 1.0 3.9 10.4 0.9 0.9 14.0 0.7 3.5
-Sprott P 5.7 0.2 1.2 9.2 0.6 0.8 10.5 0.4 0.5 11.1 0.8 1.0 13.3 0.4 0.7 13.5 0.6 1.6 9.6 0.8 4.0 7.6 0.6 3.3 8.7 0.8 3.4 9.6 0.8 3.0 10.8 0.8 1.5 10.4 0.5 1.8
-Sprott Q 5.8 0.2 1.1 9.5 0.5 0.5 9.4 0.5 0.4 10.0 0.7 0.5 12.9 0.5 1.3 12.6 0.3 1.3 11.1 0.9 0.8 13.0 0.7 1.8 12.5 0.7 0.6 11.5 0.8 1.7 9.7 0.9 0.4 13.2 0.4 4.0
-Sprott R 5.1 0.3 0.5 6.9 0.4 0.1 6.7 0.3 0.1 6.5 0.2 0.0 12.5 0.3 0.3 14.2 0.5 0.4 11.0 0.9 1.4 8.8 0.7 1.0 10.6 0.9 1.0 9.4 0.9 0.9 11.4 0.9 1.7 13.8 0.6 2.3
-Sprott S 2.9 0.0 0.3 9.6 0.4 0.3 10.7 0.2 0.2 11.7 0.3 0.3 13.0 0.4 0.3 13.3 0.4 0.7 10.2 0.9 0.6 10.2 0.8 1.6 11.2 0.9 1.7 11.0 0.8 2.0 10.7 0.9 1.9 12.8 0.5 4.0
-Sprott Torus 3.6 0.2 0.2 6.8 0.6 0.2 12.0 0.5 0.2 13.0 0.5 0.3 14.3 0.7 0.3 14.7 0.6 0.7 7.0 0.9 0.6 12.1 0.4 0.8 7.9 0.9 0.6 12.5 0.7 0.1 12.2 0.9 0.3 14.9 0.5 1.0
-Thomas 2.3 0.2 6.7 12.4 0.6 4.1 12.1 0.3 3.6 13.4 0.2 2.5 13.4 0.3 1.9 13.8 0.2 1.1 12.3 0.8 4.9 12.0 0.8 6.6 13.7 0.7 4.9 13.1 0.7 5.5 11.9 0.8 4.9 13.6 0.4 5.4
-Thomas Labyrinth 6.5 0.1 5.2 12.5 0.2 4.8 9.8 0.4 5.6 11.8 0.3 3.4 12.2 0.3 4.6 12.5 0.4 4.0 11.5 0.6 4.4 12.3 0.6 4.2 11.5 0.6 4.2 11.5 0.6 3.8 12.5 0.6 3.3 12.5 0.3 2.6
-Torus 10.9 0.9 2.3 12.4 0.4 0.1 11.9 0.2 0.1 12.0 0.2 0.2 12.9 0.5 1.0 12.9 0.7 1.5 12.0 1.0 2.1 12.6 0.8 2.3 12.6 1.0 2.0 12.4 1.0 3.1 8.7 0.9 1.7 13.1 0.6 4.6
-Tsucs 2 7.2 0.6 1.7 10.9 0.7 1.0 10.3 0.8 1.2 10.0 0.2 1.6 12.0 0.6 1.7 13.3 0.7 2.2 8.5 0.8 2.2 12.3 0.8 2.2 7.5 0.7 2.4 9.4 0.7 2.3 11.5 0.8 2.6 9.8 0.5 2.1
-Vallise El Nino 1.8 0.2 0.4 10.1 0.4 0.3 5.5 0.2 0.2 7.4 0.3 0.2 12.7 0.5 0.5 13.4 0.6 1.7 11.6 0.8 1.8 12.5 0.8 1.0 13.6 0.7 1.1 12.4 0.8 1.9 13.6 0.7 1.3 13.7 0.6 2.1
-Wang Sun 5.3 0.2 0.7 11.8 0.5 1.2 10.3 0.4 0.4 13.2 0.5 2.0 10.4 0.5 0.4 10.5 0.7 0.7 9.3 0.7 1.9 11.9 0.7 1.8 11.2 0.7 1.8 9.9 0.6 1.7 13.3 0.7 1.5 12.6 0.5 1.0
-Winsmi Reduced 5.0 0.3 1.4 8.9 0.4 1.8 11.2 0.4 1.6 13.6 0.4 1.6 13.6 0.5 1.5 14.0 0.7 1.0 11.7 0.9 2.1 11.7 0.8 2.4 11.9 0.9 1.4 10.4 0.8 2.7 11.4 0.9 0.9 14.3 0.6 2.1
-Yu Wang 6.4 0.3 3.2 9.8 0.7 3.3 11.3 0.6 3.4 10.3 0.6 3.5 8.1 0.6 3.3 12.8 0.6 3.2 12.3 0.9 7.6 12.0 0.8 5.7 12.8 0.7 6.6 12.8 0.8 7.2 12.0 0.7 5.0 13.7 0.5 4.6
-Yu Wang 2 3.4 0.1 0.3 10.0 0.2 0.1 10.6 0.4 0.2 12.7 0.2 0.1 12.2 0.2 0.5 13.8 0.4 0.5 10.8 0.9 1.4 9.4 0.8 2.5 8.6 0.7 1.6 10.0 0.7 1.4 11.6 0.9 1.8 10.7 0.3 1.2
-Zhou Chen 10.0 0.4 0.2 10.4 0.4 1.3 13.7 0.5 3.0 12.7 0.4 0.9 12.8 0.3 0.4 12.8 0.4 0.9 13.2 0.5 1.6 11.6 0.5 2.0 13.5 0.5 1.8 9.7 0.5 2.2 12.8 0.6 2.7 11.6 0.4 2.9
+System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM TimesFM
+$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
+Atmospheric Regime 8.4 0.6 2.3 9.7 0.5 0.9 11.1 0.4 0.5 10.5 0.6 0.7 10.1 0.5 1.0 12.0 0.4 0.5 11.9 0.9 2.2 12.0 0.9 1.9 13.5 0.9 1.8 12.7 0.8 2.7 11.3 0.9 1.3 14.2 0.6 3.1 13.9 0.4 1.2
+BeerRNN 11.7 0.8 1.2 8.0 0.5 0.2 3.3 0.2 0.2 9.5 0.2 0.2 7.3 0.4 0.2 11.9 0.4 0.3 12.3 0.9 1.2 12.5 0.9 1.2 12.1 0.9 1.2 12.6 0.8 1.1 13.0 0.9 1.4 14.1 0.6 1.3 14.0 0.4 0.8
+Bouali 4.1 0.1 0.5 8.3 0.5 0.1 7.3 0.3 0.2 8.6 0.1 0.2 11.1 0.3 0.5 13.2 0.5 0.3 10.6 0.8 1.0 9.6 0.7 1.1 12.2 0.7 2.0 12.3 0.8 1.2 12.9 0.8 2.3 13.4 0.4 2.3 12.7 0.3 0.7
+Burke Shaw 4.1 0.2 0.5 10.8 0.5 1.2 9.0 0.5 1.5 13.0 0.4 1.1 13.2 0.4 1.5 13.3 0.6 0.9 11.5 0.8 1.3 12.1 0.6 2.4 12.2 0.7 2.2 10.5 0.7 4.8 12.2 0.9 3.7 13.7 0.5 3.6 13.9 0.4 1.5
+Ca Two Plus Quasip. 2.3 0.1 0.7 7.2 0.1 0.2 8.3 0.1 0.3 9.2 0.4 0.3 13.1 0.4 0.2 13.2 0.4 0.6 11.3 0.9 1.0 11.6 0.8 1.1 12.8 0.6 1.2 8.9 0.6 1.6 11.1 0.8 3.0 13.8 0.7 2.6 13.4 0.3 1.9
+Chen Lee 3.9 0.4 0.3 4.9 0.3 0.3 4.0 0.5 0.5 4.8 0.3 0.1 4.6 0.3 0.3 8.9 0.7 0.4 7.2 0.9 1.6 7.9 0.7 1.4 8.0 0.7 2.0 4.8 0.5 1.7 11.0 0.9 0.9 6.4 0.5 3.4 7.3 0.4 1.8
+Chua 5.9 0.1 0.3 7.4 0.2 0.6 5.9 0.2 0.7 7.6 0.1 0.7 11.7 0.2 0.4 13.1 0.3 1.4 11.8 0.6 1.9 12.2 0.5 2.2 11.7 0.6 1.7 12.0 0.6 2.7 11.3 0.5 3.2 13.4 0.4 2.7 10.4 0.2 0.7
+Coullet 1.2 0.0 0.2 4.9 0.0 0.3 0.7 0.0 0.1 2.8 0.0 0.1 7.5 0.2 0.1 14.6 0.4 0.6 13.0 0.8 2.8 11.2 0.8 3.8 11.8 0.8 3.4 11.6 0.7 4.9 12.3 0.8 2.4 14.9 0.5 3.3 10.7 0.2 0.5
+Dadras 4.1 0.2 1.1 7.7 0.6 0.5 4.7 0.5 0.5 13.1 0.5 0.3 12.1 0.4 0.2 10.2 0.7 0.5 12.7 0.8 1.7 12.9 0.7 2.0 13.0 0.8 2.2 8.8 0.8 3.8 11.9 0.8 1.7 12.8 0.4 2.9 11.8 0.4 1.2
+Dequan Li 12.7 0.5 2.1 11.1 0.3 1.1 11.0 0.4 0.7 12.1 0.3 0.9 12.1 0.2 1.9 13.0 0.3 3.1 11.7 0.4 3.1 11.3 0.4 2.8 9.6 0.4 2.4 11.0 0.4 1.8 10.7 0.4 3.1 13.5 0.4 4.5 12.8 0.3 2.0
+Finance 3.2 0.1 0.6 8.5 0.6 1.6 8.6 0.4 0.9 9.1 0.4 0.4 12.5 0.3 0.4 13.8 0.4 1.5 11.6 0.9 1.8 10.7 0.8 1.7 13.4 0.8 2.0 11.8 0.8 1.7 11.1 0.9 2.3 13.7 0.5 4.2 13.6 0.4 2.6
+Genesio Tesi 8.2 0.2 0.9 7.9 0.4 1.0 8.5 0.2 1.1 7.7 0.2 1.0 9.4 0.2 1.0 13.0 0.3 2.0 10.4 0.7 1.4 10.6 0.7 1.1 10.3 0.8 2.2 9.6 0.8 2.1 11.0 0.9 1.4 12.6 0.6 5.4 10.6 0.4 1.0
+Guckenheimer Holmes 9.6 0.1 0.5 11.8 0.1 0.8 12.0 0.1 0.4 12.6 0.3 0.5 12.5 0.4 1.1 12.1 0.4 0.6 9.6 0.7 0.6 9.6 0.7 2.5 9.4 0.6 1.0 11.8 0.6 1.4 10.2 0.7 1.0 12.7 0.5 3.5 11.5 0.3 0.9
+Hadley 11.3 0.4 1.3 10.7 0.5 2.3 9.7 0.5 1.8 12.5 0.3 1.5 12.6 0.6 1.8 12.6 0.4 2.8 11.4 0.8 5.1 11.9 0.7 4.7 11.9 0.6 3.9 11.5 0.8 5.7 11.4 0.8 6.3 12.5 0.5 7.1 12.4 0.5 3.8
+Halvorsen 4.3 0.3 0.6 9.7 0.6 0.4 8.2 0.2 0.7 7.5 0.2 0.7 13.1 0.2 0.6 13.3 0.4 1.9 10.0 0.8 5.0 10.3 0.7 4.3 9.7 0.8 4.6 10.9 0.7 4.1 11.3 0.9 3.4 13.4 0.5 4.2 13.3 0.3 1.6
+Hindmarsh Rose 6.5 0.4 0.7 9.8 0.2 0.2 4.9 0.1 0.1 9.6 0.2 0.1 12.7 0.6 0.3 13.7 0.6 0.9 12.9 0.7 1.6 14.3 0.8 2.0 12.9 0.8 1.7 14.0 0.8 1.9 13.2 0.8 2.5 14.6 0.4 1.4 14.3 0.5 1.1
+Isothermal Chemical 2.0 0.2 0.2 2.5 0.1 0.6 5.4 0.1 0.3 8.3 0.2 0.2 7.0 0.4 0.2 11.6 0.5 0.5 11.7 0.9 1.1 11.7 0.9 1.5 10.8 0.8 1.6 12.5 0.9 2.0 10.1 0.9 0.7 15.6 0.6 3.5 7.5 0.3 1.3
+Itik Banks tumor 5.7 0.2 0.5 8.2 0.5 0.5 10.8 0.6 0.4 9.8 0.5 0.3 11.8 0.6 0.4 13.3 0.6 0.3 10.8 0.6 3.4 11.2 0.8 3.8 11.9 0.9 2.9 12.0 0.8 5.0 12.0 0.9 2.9 12.9 0.7 4.1 13.4 0.3 1.1
+Jerk Circuit 4.3 0.3 0.8 8.8 0.6 0.4 9.1 0.4 0.3 10.6 0.3 0.2 11.7 0.3 0.2 9.8 0.6 0.3 12.5 0.9 1.0 13.5 0.9 0.8 12.9 0.9 0.8 13.1 1.0 0.9 10.3 0.8 0.6 14.1 0.5 1.9 13.7 0.4 1.1
+Laser 4.5 0.2 0.2 7.5 0.5 3.0 8.2 0.3 1.2 11.5 0.2 0.5 12.8 0.2 0.5 13.9 0.6 0.8 12.6 0.7 2.0 12.8 0.5 2.8 13.0 0.7 1.7 12.8 0.7 5.1 11.8 0.8 2.5 13.3 0.5 2.6 13.8 0.4 1.0
+Lorenz 3.4 0.2 0.5 9.8 0.6 1.5 8.5 0.3 1.6 13.0 0.4 0.7 12.6 0.7 2.0 13.3 0.6 0.8 12.8 0.8 2.6 13.0 0.7 3.4 13.2 0.6 3.0 13.0 0.6 2.3 11.8 0.8 2.7 13.4 0.5 3.1 13.1 0.3 2.1
+Lorenz 84 7.9 0.2 0.5 10.7 0.5 1.5 11.0 0.4 1.3 12.6 0.4 1.0 11.0 0.4 0.4 13.2 0.6 0.4 11.8 0.9 1.3 12.2 0.9 1.4 11.5 0.9 1.3 12.3 0.9 1.3 11.2 0.9 1.2 13.4 0.6 1.2 13.4 0.5 1.4
+Lu Chen 9.5 0.8 0.7 5.3 0.2 0.5 9.0 0.3 0.7 8.9 0.2 0.2 9.7 0.3 0.5 8.3 0.3 0.6 12.2 1.0 1.7 13.8 0.9 1.8 12.2 0.9 1.6 13.1 1.0 2.1 10.1 0.8 1.1 14.2 0.5 1.9 7.4 0.5 0.9
+Moore Spiegel 9.2 0.6 0.4 9.1 0.6 0.4 10.9 0.7 0.4 12.5 0.5 0.3 11.5 0.6 0.3 11.3 0.4 0.3 11.4 0.9 2.0 11.7 0.8 2.7 12.7 0.9 2.0 12.0 0.8 2.4 10.5 0.9 2.2 13.1 0.7 1.5 12.6 0.5 2.1
+Multi Chua 12.5 0.5 0.4 12.0 0.6 0.8 11.9 0.6 0.4 13.4 0.4 1.0 13.0 0.4 0.8 13.9 0.4 1.3 12.5 0.5 1.5 12.6 0.5 1.1 13.6 0.5 1.1 13.0 0.5 1.2 10.7 0.5 0.7 13.5 0.5 2.1 12.7 0.3 1.9
+Nose Hoover 11.4 0.5 0.9 10.2 0.6 0.5 10.8 0.7 0.5 10.8 0.6 0.5 12.9 0.8 0.6 13.3 0.5 1.0 11.2 0.9 2.0 11.3 0.9 1.4 12.7 0.7 2.0 11.0 0.9 2.2 10.5 0.9 2.9 13.3 0.5 3.3 12.8 0.4 0.5
+Pehlivan Wei 4.8 0.2 1.9 4.6 0.1 0.7 6.8 0.4 0.9 9.0 0.3 1.0 10.7 0.3 1.3 12.4 0.5 2.1 11.1 0.9 5.6 12.0 0.6 6.8 12.6 0.8 6.5 12.3 0.8 6.3 10.0 0.9 4.1 12.7 0.4 7.7 9.1 0.3 3.3
+Sakarya 3.1 0.3 0.4 2.6 0.2 0.2 5.0 0.3 0.3 8.2 0.3 0.3 10.2 0.7 0.6 12.9 0.8 0.9 10.7 0.9 0.5 12.7 0.9 2.5 11.8 0.8 0.9 12.8 0.8 1.7 12.4 0.7 0.8 14.0 0.5 2.5 13.8 0.5 1.5
+San Um Srisuchinwong 5.5 0.1 0.6 5.5 0.2 1.2 10.5 0.8 1.3 12.6 0.4 1.2 13.0 0.5 0.8 13.1 0.6 0.6 11.8 0.8 1.1 13.0 0.7 0.5 12.0 0.6 1.2 11.9 0.7 0.9 10.7 0.5 1.8 12.6 0.5 0.9 12.6 0.4 1.4
+Sprott A 4.9 0.1 1.0 9.3 0.5 0.8 8.5 0.4 0.8 7.5 0.3 0.4 11.7 0.3 0.9 12.6 0.5 2.2 10.7 0.9 4.1 11.8 0.7 4.5 11.5 0.8 4.4 11.1 0.7 5.0 9.8 0.9 2.1 12.4 0.6 6.0 11.2 0.2 1.1
+Sprott B 2.6 0.2 0.9 9.2 0.5 1.7 11.0 0.6 1.8 13.2 0.5 2.5 13.4 0.6 1.5 13.4 0.7 1.9 12.6 0.8 2.7 13.2 0.8 2.7 12.5 0.8 2.2 12.6 0.7 1.9 13.9 0.8 3.9 12.9 0.6 2.9 13.7 0.3 2.5
+Sprott C 2.9 0.2 0.4 8.5 0.6 1.1 12.0 0.3 0.5 13.3 0.4 1.1 12.6 0.4 0.8 11.8 0.5 2.0 12.4 0.7 2.0 13.0 0.6 2.2 11.9 0.6 2.5 11.8 0.6 1.7 13.3 0.7 2.6 13.5 0.5 2.9 13.5 0.4 1.9
+Sprott D 3.8 0.1 0.6 7.6 0.2 0.2 5.5 0.1 0.2 5.5 0.2 0.1 12.9 0.4 0.7 13.1 0.5 0.5 12.1 0.7 0.8 12.0 0.8 4.6 12.1 0.7 3.1 12.0 0.6 3.7 12.9 0.7 1.9 13.8 0.5 2.3 13.2 0.5 0.6
+Sprott F 4.9 0.3 0.5 10.8 0.4 0.3 8.9 0.5 0.5 12.6 0.3 0.3 13.0 0.5 0.2 13.1 0.4 0.3 12.3 0.9 1.2 12.2 0.6 1.3 13.2 0.8 1.2 12.4 0.7 1.3 11.8 0.9 1.3 13.4 0.6 1.8 13.2 0.5 1.2
+Sprott G 2.4 0.0 0.3 4.2 0.1 0.1 6.2 0.2 0.1 7.1 0.2 0.2 7.7 0.3 0.5 12.9 0.3 0.6 12.3 0.8 3.6 12.4 0.7 4.0 12.9 0.8 3.4 11.2 0.9 4.5 11.0 0.9 2.6 14.0 0.5 2.0 10.6 0.3 1.2
+Sprott I 3.0 0.2 0.3 8.6 0.5 0.1 10.8 0.1 0.2 7.4 0.3 0.3 10.1 0.3 0.3 13.1 0.4 0.2 12.1 0.9 1.2 11.9 0.8 2.5 9.4 0.9 1.0 13.4 0.9 2.4 10.9 0.9 0.9 14.4 0.6 3.2 9.1 0.4 0.3
+Sprott L 2.7 0.2 0.2 9.2 0.3 0.2 8.2 0.1 0.2 10.2 0.2 0.1 11.3 0.2 0.1 13.6 0.4 0.2 12.3 0.9 2.6 12.3 0.9 2.7 13.3 0.9 1.8 13.1 0.9 4.8 12.2 0.8 0.6 14.2 0.5 2.5 11.7 0.3 0.2
+Sprott M 3.4 0.2 0.4 10.9 0.5 0.1 12.4 0.4 0.2 12.8 0.3 0.3 13.5 0.3 0.6 14.0 0.8 0.7 12.1 0.8 1.9 11.1 0.7 2.3 13.0 0.8 1.2 12.3 0.8 1.5 10.7 0.9 1.3 14.3 0.6 3.3 13.5 0.4 0.6
+Sprott O 3.5 0.1 0.4 9.2 0.2 0.4 6.7 0.1 0.4 8.1 0.2 0.4 11.6 0.4 0.3 13.2 0.3 0.5 11.8 1.0 1.2 12.7 0.8 2.1 11.4 1.0 2.3 12.6 1.0 3.1 10.4 0.9 0.8 14.0 0.7 3.2 9.6 0.2 0.5
+Sprott P 5.7 0.2 0.4 9.2 0.6 0.9 10.5 0.4 0.6 11.1 0.8 0.9 13.3 0.4 0.5 13.5 0.6 1.7 9.6 0.8 6.8 7.6 0.6 5.7 8.7 0.8 4.7 9.6 0.8 5.8 10.8 0.8 1.4 10.4 0.5 3.2 9.1 0.5 1.9
+Sprott Q 5.8 0.2 0.8 9.5 0.5 0.5 9.4 0.5 0.5 10.0 0.7 0.4 12.9 0.5 1.3 12.6 0.3 1.5 11.1 0.9 1.7 13.0 0.7 2.2 12.5 0.7 0.7 11.5 0.8 3.8 9.7 0.9 1.1 13.2 0.4 6.0 12.3 0.4 0.8
+Sprott R 5.1 0.3 0.7 6.9 0.4 0.3 6.7 0.3 0.2 6.5 0.2 0.1 12.5 0.3 0.7 14.2 0.5 1.4 11.0 0.9 2.8 8.8 0.7 2.2 10.6 0.9 3.2 9.4 0.9 1.9 11.4 0.9 2.2 13.8 0.6 6.0 8.9 0.3 1.6
+Sprott S 2.9 0.0 0.3 9.6 0.4 0.4 10.7 0.2 0.3 11.7 0.3 0.5 13.0 0.4 0.2 13.3 0.4 0.6 10.2 0.9 0.8 10.2 0.8 3.4 11.2 0.9 3.1 11.0 0.8 3.5 10.7 0.9 2.2 12.8 0.5 6.0 11.3 0.4 1.9
+Sprott Torus 3.6 0.2 0.1 6.8 0.6 0.2 12.0 0.5 0.2 13.0 0.5 0.3 14.3 0.7 0.3 14.7 0.6 0.8 7.0 0.9 1.1 12.1 0.4 1.1 7.9 0.9 0.8 12.5 0.7 0.8 12.2 0.9 1.1 14.9 0.5 1.5 12.8 0.4 0.9
+Thomas 2.3 0.2 3.1 12.4 0.6 2.9 12.1 0.3 3.3 13.4 0.2 2.5 13.4 0.3 2.9 13.8 0.2 2.7 12.3 0.8 4.4 12.0 0.8 5.2 13.7 0.7 4.2 13.1 0.7 4.9 11.9 0.8 4.0 13.6 0.4 4.6 13.7 0.3 4.3
+Thomas Labyrinth 6.5 0.1 3.2 12.5 0.2 3.2 9.8 0.4 3.3 11.8 0.3 2.5 12.2 0.3 3.3 12.5 0.4 2.8 11.5 0.6 4.1 12.3 0.6 3.8 11.5 0.6 3.6 11.5 0.6 3.6 12.5 0.6 2.5 12.5 0.3 2.3 12.6 0.3 2.6
+Torus 10.9 0.9 0.9 12.4 0.4 0.0 11.9 0.2 0.0 12.0 0.2 0.0 12.9 0.5 0.1 12.9 0.7 0.5 12.0 1.0 0.5 12.6 0.8 0.7 12.6 1.0 0.6 12.4 1.0 0.9 8.7 0.9 0.3 13.1 0.6 1.3 12.5 0.3 0.4
+Tsucs 2 7.2 0.6 1.5 10.9 0.7 1.9 10.3 0.8 1.4 10.0 0.2 2.0 12.0 0.6 2.4 13.3 0.7 2.6 8.5 0.8 2.8 12.3 0.8 3.9 7.5 0.7 3.4 9.4 0.7 4.0 11.5 0.8 3.3 9.8 0.5 4.2 13.6 0.4 3.7
+Vallise El Nino 1.8 0.2 0.2 10.1 0.4 0.2 5.5 0.2 0.2 7.4 0.3 0.2 12.7 0.5 0.3 13.4 0.6 1.2 11.6 0.8 1.3 12.5 0.8 1.3 13.6 0.7 1.2 12.4 0.8 1.3 13.6 0.7 1.0 13.7 0.6 1.6 13.2 0.5 1.9
+Wang Sun 5.3 0.2 0.5 11.8 0.5 1.7 10.3 0.4 0.7 13.2 0.5 1.7 10.4 0.5 0.4 10.5 0.7 0.9 9.3 0.7 1.7 11.9 0.7 1.9 11.2 0.7 2.2 9.9 0.6 2.5 13.3 0.7 2.3 12.6 0.5 1.4 10.5 0.4 1.5
+Winsmi Reduced 5.0 0.3 1.8 8.9 0.4 3.4 11.2 0.4 3.9 13.6 0.4 3.4 13.6 0.5 3.7 14.0 0.7 4.0 11.7 0.9 2.7 11.7 0.8 4.4 11.9 0.9 3.0 10.4 0.8 5.3 11.4 0.9 2.1 14.3 0.6 4.5 13.8 0.3 2.4
+Yu Wang 6.4 0.3 0.3 9.8 0.7 0.5 11.3 0.6 0.6 10.3 0.6 0.7 8.1 0.6 0.6 12.8 0.6 0.6 12.3 0.9 2.9 12.0 0.8 1.8 12.8 0.7 2.1 12.8 0.8 2.2 12.0 0.7 2.7 13.7 0.5 1.8 13.4 0.4 1.7
+Yu Wang 2 3.4 0.1 0.2 10.0 0.2 0.2 10.6 0.4 0.4 12.7 0.2 0.1 12.2 0.2 0.7 13.8 0.4 0.6 10.8 0.9 2.6 9.4 0.8 5.0 8.6 0.7 4.0 10.0 0.7 4.8 11.6 0.9 1.9 10.7 0.3 1.7 13.5 0.3 2.0
+Zhou Chen 10.0 0.4 0.8 10.4 0.4 2.6 13.7 0.5 4.9 12.7 0.4 1.2 12.8 0.3 3.1 12.8 0.4 2.1 13.2 0.5 4.1 11.6 0.5 4.4 13.5 0.5 4.3 9.7 0.5 5.4 12.8 0.6 5.1 11.6 0.4 7.0 14.3 0.5 5.4

### v3

@@
 1 Introduction
-Figure 1: DynaMix achieves zero-shot DSR of attractor geometry and long-term temporal properties (power spectrum) from a short context signal while Chronos [2] fails.
-Most real-world processes, from atmospheric phenomena and stock markets to brain activity or ecological networks, can be described as dynamical systems (DS) [56, 15, 86, 59].
-Reconstructing these from observational data, called dynamical systems reconstruction (DSR), has been a long-standing challenge in scientific modeling [98, 23, 31].
+Figure 1: DynaMix achieves zero-shot DSR of attractor geometry and long-term temporal properties (power spectrum) from a short context signal while Chronos Ansari et al. (2024) fails.
+Most real-world processes, from atmospheric phenomena and stock markets to brain activity or ecological networks, can be described as dynamical systems (DS) Mandelbrot and Hudson (2007); Buzsaki (2006); Tziperman et al.
+(1997); Mumby et al.
+(2007).
+Reconstructing these from observational data, called dynamical systems reconstruction (DSR), has been a long-standing challenge in scientific modeling Yu and Wang (2024); Durstewitz et al.
+(2023); Gilpin (2024).
 DSR goes beyond conventional time series (TS) modeling, as we wish to have a generative model of the underlying process which exhibits long-term behavior with the same temporal and geometrical signatures (Fig.
 1), i.e.
-the same invariant or ‘climate’ statistics [66, 70], as important in scientific applications.
-Achieving this usually requires special control-theoretic training techniques [58, 39] or loss objectives [69, 70, 42, 76] which accentuate the system’s long-term dynamics.
-Numerous deep learning approaches for DSR, based on recurrent neural networks (RNNs; [85, 22, 88, 16, 6, 73, 39, 7, 70]), neural ODEs [17, 45, 1, 47], Koopman operators [14, 55, 63, 3, 60, 28, 91], or library-based methods [12, 53, 43, 19, 57], have been advanced over the years.
-However, all of these require purpose-training on the specific system observed and struggle to generalize beyond their training distribution [33].
-Inspired by the strong in-context and zero-shot generalization abilities of LLMs [10, 27, 21, 18], there has been a push recently to develop models with likewise properties for the time series domain.
-Time series foundation models like Chronos [2] or Mamba4Cast [5] are pre-trained on a huge database of real-world and artificial time series, and then tasked to forecast novel time series from which snippets are presented in-context, without any parameter fine-tuning.
+the same invariant or ‘climate’ statistics Patel and Ott (2023); Platt et al.
+(2023), as important in scientific applications.
+Achieving this usually requires special control-theoretic training techniques Mikhaeil et al.
+(2022); Hess et al.
+(2023) or loss objectives Platt et al.
+(2022, 2023); Jiang et al.
+(2023); Schiff et al.
+(2024) which accentuate the system’s long-term dynamics.
+Numerous deep learning approaches for DSR, based on recurrent neural networks (RNNs; Trischler and D’Eleuterio (2016); Durstewitz (2017); Vlachas et al.
+(2018); Cestnik and Abel (2019); Brenner et al.
+(2022); Rusch et al.
+(2022); Hess et al.
+(2023); Brenner et al.
+(2024a); Platt et al.
+(2023)), neural ODEs Chen et al.
+(2018); Karlsson and Svanström (2019); Alvarez et al.
+(2020); Ko et al.
+(2023), Koopman operators Brunton et al.
+(2021); Lusch et al.
+(2018); Otto and Rowley (2019); Azencot et al.
+(2020); Naiman and Azencot (2021); Geneva and Zabaras (2022); Wang et al.
+(2022), or library-based methods Brunton et al.
+(2016); Loiseau and Brunton (2018); Kaiser et al.
+(2018); Cortiella et al.
+(2021); Messenger and Bortz (2021), have been advanced over the years.
+However, all of these require purpose-training on the specific system observed and struggle to generalize beyond their training distribution Göring et al.
+(2024b).
+Inspired by the strong in-context and zero-shot generalization abilities of LLMs Brown et al.
+(2020); Garg et al.
+(2022); Dong et al.
+(2022); Coda-Forno et al.
+(2023), there has been a push recently to develop models with likewise properties for the time series domain.
+Time series foundation models like Chronos Ansari et al.
+(2024, 2025) or Mamba4Cast Bhethanabhotla et al.
+(2024) are pre-trained on a huge database of real-world and artificial time series, and then tasked to forecast novel time series from which snippets are presented in-context, without any parameter fine-tuning.
 They are not built for DSR, however, and – as we show here – typically fail to properly capture a system’s long-term behavior and the structure of its attractors (Fig.
-1; [99]).
+1; Zhang and Gilpin (2025)).
 To address this gap, we introduce DynaMix, a novel mixture-of-experts model designed for DSR and zero-shot forecasting.
@@
   • Accurate zero-shot DSR: DynaMix achieves strong generalization across diverse DS, eliminating the need for fine-tuning while maintaining accuracy in reproducing attractor geometry and long-term statistics.
+No other model tested achieved this.
   • Multivariate information transfer: Due to its multivariate architecture, the model efficiently captures dependencies among multiple system dimensions, enabling accurate reconstruction of their coupled dynamics.
 Further, it is neither bound to a specific dimensionality nor to a specific context length, but can flexibly adapt to other dimensions through specific embeddings.
-  • Computational and parameter efficiency: The model achieves high performance with a very lightweight architecture ($\approx 10$k parameters) and small training corpus ($34$ systems), enabling orders of magnitude faster inference than other foundation models.
+  • Computational and parameter efficiency: The model reaches high performance with a very lightweight architecture ($\approx 10$k parameters) and small training corpus ($34$ systems), enabling orders of magnitude faster inference than other foundation models.
   • Interpretable dynamics composition: The model provides insight into the dynamical composition of reconstructed systems, elucidating similarities between different DS.
@@
 Dynamical systems reconstruction (DSR)
-In DSR we seek to learn generative models from time series data that represent the underlying system dynamics with all its topological, geometrical, and temporal properties [23, 33, 70, 30, 13, 11], in that sense providing an approximation to the underlying system’s governing equations.
+In DSR we seek to learn generative models from time series data that represent the underlying system dynamics with all its topological, geometrical, and temporal properties Durstewitz et al.
+(2023); Göring et al.
+(2024b); Platt et al.
+(2023); Gilpin (2023); Brunton et al.
+(2017); Brunton and Kutz (2019), in that sense providing an approximation to the underlying system’s governing equations.
 By definition, such a model should not only render viable short-term forecasts, but also reproduce the long-term evolution of the DS it has been trained on, both in state space and in the time domain.
-Methods approaching this goal have been founded on predefined function libraries, such as Sparse Identification of Nonlinear Dynamics (SINDy; [12, 53, 19, 43, 57]), on reservoir computers [67, 69, 70], neural ODEs [17, 45, 1, 47], Koopman operators [14, 55, 63, 3, 60, 28, 91], or different types of RNNs [85, 22, 88, 16, 6, 73, 39, 7].
-More important than the architecture itself seems to be their proper training in order to ensure long-term (invariant) statistics of the underlying system are met: Methods like sparse [58, 6] or generalized [39] teacher forcing allow RNN-generated trajectories to ‘explore the future’ whilst training, yet keep loss gradients in check.
-Alternatively, long-term statistics based on the observed system’s Lyapunov spectrum, fractal geometry, or invariant measures may be added to regularize the loss [69, 70, 42, 76], but require to compute these properties from the data first.
-Despite these advances, out-of-domain generalization remains a key challenge in DSR [33].
-Meta- and hierarchical learning models, trained across many DS simultaneously, have recently been designed as steps toward foundation models for DSR [9, 62, 46, 97], but still require parameter fine-tuning and lack in-context inference capabilities.
+Methods approaching this goal have been founded on predefined function libraries, such as Sparse Identification of Nonlinear Dynamics (SINDy; Brunton et al.
+(2016); Loiseau and Brunton (2018); Cortiella et al.
+(2021); Kaiser et al.
+(2018); Messenger and Bortz (2021)), on reservoir computers Pathak et al.
+(2017); Platt et al.
+(2022, 2023), neural ODEs Chen et al.
+(2018); Karlsson and Svanström (2019); Alvarez et al.
+(2020); Ko et al.
+(2023), Koopman operators Brunton et al.
+(2021); Lusch et al.
+(2018); Otto and Rowley (2019); Azencot et al.
+(2020); Naiman and Azencot (2021); Geneva and Zabaras (2022); Wang et al.
+(2022), or different types of RNNs Trischler and D’Eleuterio (2016); Durstewitz (2017); Vlachas et al.
+(2018); Cestnik and Abel (2019); Brenner et al.
+(2022); Rusch et al.
+(2022); Hess et al.
+(2023); Brenner et al.
+(2024a).
+More important than the architecture itself seems to be their proper training in order to ensure long-term (invariant) statistics of the underlying system are met: Methods like sparse Mikhaeil et al.
+(2022); Brenner et al.
+(2022) or generalized Hess et al.
+(2023) teacher forcing allow RNN-generated trajectories to ‘explore the future’ whilst training, yet keep loss gradients in check.
+Alternatively, long-term statistics based on the observed system’s Lyapunov spectrum, fractal geometry, or invariant measures may be added to regularize the loss Platt et al.
+(2022, 2023); Jiang et al.
+(2023); Schiff et al.
+(2024), but require to compute these properties from the data first.
+Despite these advances, out-of-domain generalization remains a key challenge in DSR Göring et al.
+(2024b).
+Meta- and hierarchical learning models, trained across many DS simultaneously, have recently been designed as steps toward foundation models for DSR Brenner et al.
+(2025); Nzoyem et al.
+(2025); Kirchmeyer et al.
+(2022); Yin et al.
+(2021), but still require parameter fine-tuning and lack in-context inference capabilities.
 Time series foundation models
-Large language models (LLMs) exhibit an impressive ability to infer patterns from prompts (context) and generalize to novel situations without retraining [10, 27, 21, 18], although the underlying reasons for this are still a matter of debate [51, 90, 96].
-This inspired the development of general-purpose time series (TS) foundation models which could accurately forecast time series from a short segment provided ‘in-context’ [2, 26], without the need of task-specific fine-tuning [20].
-One idea is to simply use pretrained LLMs directly as TS forecasters [34, 71, 79, 93, 83], but the inherent differences between textual and many temporal data, such as continuity, present significant challenges [82].
-To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos [2] or TimesFM [20], were specifically pretrained on a large corpus of time series data.
-Promising zero-shot forecasting capabilities have also been achieved with alternative architectural designs, such as Tiny Time Mixers [26] or state-space models like Mamba [5].
-The success of TS foundation models raised hope these could also be utilized for zero-shot DSR, but the – to our knowledge – so far only previous study on this, based on Chronos, had mixed outcomes and fell short of a full DSR evaluation [99]: Successful DSR, if present at all, heavily depended on the initial conditions, and mechanisms such as context parroting visually gave the wrong illusion that features of the dynamics had been captured.
+Large language models (LLMs) exhibit an impressive ability to infer patterns from prompts (context) and generalize to novel situations without retraining Brown et al.
+(2020); Garg et al.
+(2022); Dong et al.
+(2022); Coda-Forno et al.
+(2023), although the underlying reasons for this are still a matter of debate Lin and Lee (2024); Von Oswald et al.
+(2023); Xie et al.
+(2021).
+This inspired the development of general-purpose time series (TS) foundation models which could accurately forecast time series from a short segment provided ‘in-context’ Ansari et al.
+(2024); Ekambaram et al.
+(2024), without the need of task-specific fine-tuning Das et al.
+(2024).
+One idea is to simply use pretrained LLMs directly as TS forecasters Gruver et al.
+(2023); Rasul et al.
+(2023); Sun et al.
+(2023); Williams et al.
+(2024); Tang et al.
+(2025), but the inherent differences between textual and many temporal data, such as continuity, present significant challenges Tan et al.
+(2024).
+To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos Ansari et al.
+(2024, 2025) or TimesFM Das et al.
+(2024), were specifically pretrained on a large corpus of time series data.
+Promising zero-shot forecasting capabilities have also been achieved with alternative architectural designs, such as Tiny Time Mixers Ekambaram et al.
+(2024) or state-space models like Mamba Bhethanabhotla et al.
+(2024).
+The success of TS foundation models raised hope these could also be utilized for zero-shot DSR, but the – to our knowledge – so far only previous study on this, based on Chronos, had mixed outcomes and fell short of a full DSR evaluation Zhang and Gilpin (2025): Successful DSR, if present at all, heavily depended on the initial conditions, and mechanisms such as context parroting visually gave the wrong illusion that features of the dynamics had been captured.
 In fact, here we show that existing TS foundation models are generally not capable of producing valid DSRs (Fig.
 1), and highlight some of their failure modes.
-This is in contrast to the zero-shot DSR foundation architecture we develop here, based on models [7] and training algorithms [58] successful in DSR.
+This is in contrast to the zero-shot DSR foundation architecture we develop here, based on models Brenner et al.
+(2024a) and training algorithms Mikhaeil et al.
+(2022) successful in DSR.
 3 Methods
@@
 10), with different experts possibly specializing in different dynamical regimes.
-As a SOTA DSR base model for the experts, we leverage a recent parameter-friendly RNN, which allows for highly efficient DSR training and is designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; [7]):
+As a SOTA DSR base model for the experts, we leverage a recent parameter-friendly RNN, which allows for highly efficient DSR training and is designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; Brenner et al.
+(2024a)):
 $\bm{z}_{t}=\bm{A}\bm{z}_{t-1}+\bm{W}\Phi^{*}(\bm{z}_{t-1})+\bm{h}\;.$ (1)
@@
 Both are passed into a state attention mechanism defined as
-$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T_{C}}^{\top}\right|^{\top}\bm{1}_{N}}{\tau_{\text{att}}}\right)\in\mathbb{R}^{T}\;,$ (3)
+$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T_{C}}^{\top}\right|^{\top}\bm{1}_{N}}{\tau_{\text{att}}}\right)\in\mathbb{R}^{T_{C}}\;,$ (3)
 where $\bm{D}\in\mathbb{R}^{N\times M}$ is a learnable matrix which maps the latent state into observation space, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{\Sigma})$ exploration noise with a learnable covariance $\bm{\Sigma}$, $\tau_{\text{att}}$ a learnable temperature parameter, $\bm{1}_{\{T_{C},N\}}$ column vectors of ones of length $T_{C},N$, respectively, and $\sigma(\cdot)$ the softmax returning normalized weights.
@@
 3.2 Model training
-For training, a collection of just $34$ different chaotic and cyclic DS is used [29], see Appx.
+For training, a collection of just $34$ different chaotic and cyclic DS is used Gilpin (2022), see Appx.
 A.2 for full specification and Fig.
@@
 A MSE loss is then computed across ground truth, $\bm{X}_{T_{C}-\Delta t+1:T}$, and respective model-generated time series $\hat{\bm{X}}$.
-The model is trained by sparse teacher forcing (STF; [6, 58]), a control-theoretic technique specifically designed for DSR.
-STF replaces part of the forward-iterated latent states $\bm{z}_{t}$ by data-inferred states $\hat{\bm{z}}_{t}$ at optimally chosen intervals $\tau$, to avoid exploding gradients even for chaotic systems while enabling the DSR model “to explore the future” (see [58] for detailed theoretical motivation).
+The model is trained by sparse teacher forcing (STF; Brenner et al.
+(2022); Mikhaeil et al.
+(2022)), a control-theoretic technique specifically designed for DSR.
+STF replaces part of the forward-iterated latent states $\bm{z}_{t}$ by data-inferred states $\hat{\bm{z}}_{t}$ at optimally chosen intervals $\tau$, to avoid exploding gradients even for chaotic systems while enabling the DSR model “to explore the future” (see Mikhaeil et al.
+(2022) for detailed theoretical motivation).
 As shown in Fig.
@@
 3.3 DSR evaluation & time series forecasting
-We used two established measures to assess DSR quality [48, 6, 39, 94, 99, 70, 58, 64, 30]: First, to quantify agreement in state space (attractor) geometry, we employed a Kullback-Leibler divergence defined across space, $D_{stsp}$ (see Appx.
+We used two established measures to assess DSR quality Koppe et al.
+(2019); Brenner et al.
+(2022); Hess et al.
+(2023); Wood (2010); Zhang and Gilpin (2025); Platt et al.
+(2023); Mikhaeil et al.
+(2022); Pals et al.
+(2024); Gilpin (2023): First, to quantify agreement in state space (attractor) geometry, we employed a Kullback-Leibler divergence defined across space, $D_{stsp}$ (see Appx.
 A.1 for details).
 For assessing agreement in long-term temporal properties, we used the Hellinger distance $D_{H}$ defined across power spectra of the true and model-generated trajectories.
+Crucially, to properly assess the long-term dynamics, both these measures were evaluated in the limit of large $T=10,000$.
 In addition, a short-term, $n$-step-ahead mean absolute prediction error (MAE) was evaluated.
-As has been repeatedly emphasized in the statistical and DSR literature [94, 48, 6, 58, 68, 30], prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig.
+As has been repeatedly emphasized in the statistical and DSR literature Wood (2010); Koppe et al.
+(2019); Brenner et al.
+(2022); Mikhaeil et al.
+(2022); Platt et al.
+(2021); Gilpin (2023), prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig.
 24).
 Empirically observed time series are often just one-dimensional ($N^{*}=1$), but come from an inherently much higher-dimensional underlying DS.
-Our approach can in principle handle context signals of dimension $N^{*}<N$ smaller than those used in training by, e.g., just filling up $N-N^{*}$ dimensions in $\bm{C}$ with zeros.
-However, to work efficiently in a DSR setting, the context signal should represent the underlying DS as good as possible, for which it commonly needs to be lifted into a higher-dimensional space where sets explored by the dynamics ideally become diffeomorphic to those in the system’s true state space.
-Temporal delay embedding is the most popular technique to achieve this [80, 74].
-For a $1d$ time series $\{x_{t}\}$, a $d$-dimensional embedding can be defined as [49]
+To work efficiently in a DSR setting, the context signal should represent the underlying DS as good as possible, for which it commonly needs to be lifted into a higher-dimensional space where sets explored by the dynamics ideally become diffeomorphic to those in the system’s true state space.
+Temporal delay embedding is the most popular technique to achieve this Takens (1981); Sauer et al.
+(1991).
+For a $1d$ time series $\{x_{t}\}$, a $d$-dimensional embedding can be defined as Kraemer et al.
+(2021)
 $\bm{x}^{emb}_{t}=(x_{t},x_{t-\tau_{1}},...,x_{t-\tau_{d-1}})\;,$ (5)
@@
 Fig.
-3b highlights further aspects of DynaMix’ reconstructions: Even though not trained on any $2d$ systems, the model successfully recreates the limit cycle from the $2d$ Selkov system [77].
+3b highlights further aspects of DynaMix’ reconstructions: Even though not trained on any $2d$ systems, the model successfully recreates the limit cycle from the $2d$ Selkov system SEL’KOV (1968).
 Notably, the model even generalizes to new initial conditions outside the scope of the context data, i.e.
@@
 22 for further examples).
-Likewise, although trained only on $3d$ DS, DynaMix also successfully generalizes to higher-dimensional DS, as illustrated for the $6d$ Lorenz-96 system [54] in Fig.
+Likewise, although trained only on $3d$ DS, DynaMix also successfully generalizes to higher-dimensional DS, as illustrated for the $6d$ Lorenz-96 system Lorenz (1996) in Fig.
 34 (see Appx.
@@
 sampling frequency.
-We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos [2], Mamba4Cast [5], TimesFM [20], and Tiny Time Mixers [26], using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see Sect.
+We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos Ansari et al.
+(2024), Chronos-2 Ansari et al.
+(2025), Panda Lai et al.
+(2025), Mamba4Cast Bhethanabhotla et al.
+(2024), TimesFM Das et al.
+(2024), and Tiny Time Mixers Ekambaram et al.
+(2024), using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see Sect.
 A.1).
@@
 4a-b, DynaMix clearly and significantly outperforms all of these.
-In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the (strongest in our batch) Chronos-t5 models, as shown in Fig.
+In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the strongest competitors in our batch, as shown in Fig.
 4c.
+This is especially surprising when compared to Panda, which included variations of our test set DS in its training repertoire, biasing the evaluation in its favor.
 Table LABEL:tab:performance_DSR assembles further results on performance comparisons, while Figs.
@@
 SOTA DSR models explicitly trained on the context data, incl.
-the plain AL-RNN trained by STF [7], Neural ODEs [37, 17], and reservoir computers which are commonly used for DSR [70] (see Appx.
+the plain AL-RNN trained by STF Brenner et al.
+(2024a), Neural ODEs Göring et al.
+(2024a); Chen et al.
+(2018), and reservoir computers which are commonly used for DSR Platt et al.
+(2023) (see Appx.
 A.7 for details).
-Despite the custom trained models’ unduly advantage, performing – unlike DynaMix – only in-domain generalization [37], DynaMix keeps up with, and in some cases even outperforms, them (Appx.
+Despite the custom trained models’ unduly advantage, performing – unlike DynaMix – only in-domain generalization Göring et al.
+(2024a), DynaMix keeps up with, and in some cases even outperforms, them (Appx.
 Tables 68).
@@
 1 & 6b, in the longer run they often converge to fixed points or simple cycles where the true dynamics is chaotic, not reflecting the context anymore.
-Third, context parroting, a phenomenon originally identified in LLMs [4], has also been observed in TS foundation models [99], where they tend to repetitively reproduce the exact input pattern rather than dynamically adapting to the system’s true evolution.
+Third, context parroting, a phenomenon originally identified in LLMs Bender et al.
+(2021), has also been observed in TS foundation models Zhang and Gilpin (2025), where they tend to repetitively reproduce the exact input pattern rather than dynamically adapting to the system’s true evolution.
 Thus, TS foundation models simply produce oscillations, where the true dynamics is chaotic and therefore by definition irregular and non-repeating (Fig.
@@
 This is illustrated in more detail in Fig.
-6d for Chronos-t5-base, the best of the TS foundation models in our batch.
+6d for Chronos-t5-base, the best of the TS foundation models according to $D_{stsp}$.
 It is particularly evident in the power spectrum (Fig.
 6d right), where Chronos’ sharp, narrow peaks wrongly suggest periodicity, while the true spectrum is usually rather broad and smeared out for chaotic systems, as the one exhibited by DynaMix.
-We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent [58], a measure for the exponential divergence rate of trajectories (see Appx.
+We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent Mikhaeil et al.
+(2022), a measure for the exponential divergence rate of trajectories (see Appx.
 A.1).
@@
 8a confirms for a $1d$ time series from the $3d$ chaotic Lorenz-63 system that this works well and DynaMix, in contrast to all other models, is still able to reconstruct the chaotic behavior.
-When applied to real-world time series, such as traffic, weather, or cloud request data, where previous TS foundation models have shown inconsistent results [84], our model produces accurate forecasts which reflect the dynamics well (Fig.
+When applied to real-world time series, such as traffic, weather, or cloud request data, where previous TS foundation models have shown inconsistent results Toner et al.
+(2025), our model produces accurate forecasts which reflect the dynamics well (Fig.
 8b-f, see Appx.
@@
 This is in stark contrast to the other TS foundation models tested here, which mostly had examples of such empirical data in their training corpus and hence might be expected to have an edge over DynaMix.
-Many real-world data, like weather, human fMRI, or EEG data, are known to bear signatures of deterministic chaos [58, 50], potentially at least partly explaining DynaMix’ strongly competitive forecasts, although it has never seen such type of data in training.
+Many real-world data, like weather, human fMRI, or EEG data, are known to bear signatures of deterministic chaos Mikhaeil et al.
+(2022); Kramer et al.
+(2022), potentially at least partly explaining DynaMix’ strongly competitive forecasts, although it has never seen such type of data in training.
 4.5 Which ingredients are most crucial for DynaMix’ performance? Ablation studies
@@
 Replacing the final gating step (the MLP) by a simple linear operation and the exact expert model within the class of piecewise linear RNNs, however, had less of an impact.
-Similar results were obtained when the AL-RNN was swapped for the clipped shallow PLRNN [39], and a probabilistic version of the AL-RNN (adding a noise term $\bm{z}_{t}=F(\bm{z}_{t-1})+\bm{\eta},\bm{\eta}\sim\mathcal{N}(0,\Gamma)$) even slightly improved long-term performance, albeit at the cost of some short-term prediction accuracy.
-Using as experts LSTMs [41], vanilla RNNs or reservoir computers [65] instead led to decay in performance, in line with previous results establishing PLRNNs as SOTA DSR models which allow for efficient training by STF [6, 39, 7].
+Similar results were obtained when the AL-RNN was swapped for the clipped shallow PLRNN Hess et al.
+(2023), and a probabilistic version of the AL-RNN (adding a noise term $\bm{z}_{t}=F(\bm{z}_{t-1})+\bm{\eta},\bm{\eta}\sim\mathcal{N}(0,\Gamma)$) even slightly improved long-term performance, albeit at the cost of some short-term prediction accuracy.
+Using as experts LSTMs Hochreiter and Schmidhuber (1997), vanilla RNNs or reservoir computers Patel and Ott (2022) instead led to decay in performance, in line with previous results establishing PLRNNs as SOTA DSR models which allow for efficient training by STF Brenner et al.
+(2022); Hess et al.
+(2023); Brenner et al.
+(2024a).
 Hence, it appears that the whole ‘DSR package’ consisting of 1) architectural choices, 2) a training algorithm specialized for DSR, and 3) properties of the DS training corpus, is crucial to DynaMix’ success.
@@
 Here we introduce the first DSR foundation model that achieves zero-shot reconstructions of novel DS, accurately reproducing their long-term statistics, just from a provided context signal without any retraining or fine-tuning.
-It hugely outperforms current TS foundation models like Chronos in terms of DSR quality, although much more lightweight with only a fraction (0.1%) of the number of parameters.
+It hugely outperforms current TS foundation models like Chronos or Panda in terms of DSR quality, although much more lightweight with only a fraction (0.1%) of the number of parameters.
 Consequently, DynaMix features inference times about an order of magnitude faster than the closest competitor.
-Most surprisingly, it even outperforms current TS foundation models on empirical data of types it has – in contrast to its competitors – never seen in training, often producing also better short-term forecasts.
+Most surprisingly, it even outperforms current TS foundation models on empirical data of types it has – in contrast to competitors – never seen in training, often producing also better short-term forecasts.
 All of this is achieved with a fairly narrow training corpus consisting mainly of $3d$ synthetic data from chaotic systems, which suggests there may be considerable room for further improvement by extending DynaMix’ training to include, e.g., various empirical time series.
@@
 Besides DynaMix’ specific DSR architecture, control-theoretically motivated training techniques, and embedding principles, also the fact that it has seen many chaotic systems in training may play a role.
-Most complex DS in nature [15, 86, 59], engineering [78], and society [56, 92] are likely chaotic.
-Moreover, many chaotic attractors, like the Lorenz-63, feature a skeleton of infinitely many (unstable) periodic orbits of all possible periods [36], thus inherently expressing a wide spectrum of temporal patterns.
+Most complex DS in nature Buzsaki (2006); Tziperman et al.
+(1997); Mumby et al.
+(2007), engineering Strogatz (2024), and society Mandelbrot and Hudson (2007); Watts and Strogatz (1998) are likely chaotic.
+Moreover, many chaotic attractors, like the Lorenz-63, feature a skeleton of infinitely many (unstable) periodic orbits of all possible periods Guckenheimer and Holmes (1983), thus inherently expressing a wide spectrum of temporal patterns.
 Limitations
-For now, we have mainly focused on rather stationary synthetic and empirical time series (but see Fig.
+For now, we have mainly focused on rather stationary time series (but see Fig.
 8b).
@@
 This may partly be amended by explicitly including such types of non-stationary and multiscale DS in the training corpus.
-Another potential solution may be adding explicit filtering and decomposition modules, similar as in Auto- [95] or FEDformers [101].
+Another potential solution may be adding explicit filtering and decomposition modules, similar as in Auto- Wu et al.
+(2021) or FEDformers Zhou et al.
+(2022).
 A proof of concept of this idea is provided in Appx.
@@
 Training method
-For training our model, we used a variant of sparse teacher forcing (STF), a control-theoretic method for DSR, with the theoretical background and rationale developed in [58].
-STF and related training methods [6, 8, 39, 58, 24] circumvent the exploding gradient problem when training on chaotic systems, while still enabling the DSR model to sufficiently forward-iterate trajectories into the future to capture long-term properties (see [58] for details), and produces state-of-the-art results for DSR [7, 9].
-In STF, current latent states states $\bm{z}_{t}$ are replaced with states inferred from the data at fixed intervals $\tau$ by (pseudo-)inversion of the decoder model, $\tilde{\bm{z}}_{t}=g^{-1}(\bm{x}_{t})$, thus recalibrating the trajectory at times ideally chosen based on the system’s Lyapunov spectrum [58] (or simply treating $\tau$ as a hyper-parameter).
+For training our model, we used a variant of sparse teacher forcing (STF), a control-theoretic method for DSR, with the theoretical background and rationale developed in Mikhaeil et al.
+[2022].
+STF and related training methods Brenner et al.
+[2022, 2024b], Hess et al.
+[2023], Mikhaeil et al.
+[2022], Eisenmann et al.
+[2024] circumvent the exploding gradient problem when training on chaotic systems, while still enabling the DSR model to sufficiently forward-iterate trajectories into the future to capture long-term properties (see Mikhaeil et al.
+[2022] for details), and produces state-of-the-art results for DSR Brenner et al.
+[2024a, 2025].
+In STF, current latent states states $\bm{z}_{t}$ are replaced with states inferred from the data at fixed intervals $\tau$ by (pseudo-)inversion of the decoder model, $\tilde{\bm{z}}_{t}=g^{-1}(\bm{x}_{t})$, thus recalibrating the trajectory at times ideally chosen based on the system’s Lyapunov spectrum Mikhaeil et al.
+[2022] (or simply treating $\tau$ as a hyper-parameter).
 In our case of an identity mapping from a latent subspace to the observations, $\hat{\bm{x}}_{t}=\mathcal{I}_{(N\times M)}\bm{z}_{t}$, this inversion becomes trivial, where $N$ is the number of observations (and thus readout neurons) and $\mathcal{I}_{(N\times M)}$ an $N\times M$ matrix with $\mathcal{I}_{(N\times M),rr}=1$ for $r\leq N$ and zeros else.
@@
 where we chose $\lambda=0.1$ and $c=0.01$.
-Rectified adaptive moment estimation (RADAM) [52] was employed as the optimizer, with $L=50$ batches of $S_{B}=16$ sequences per epoch, each of length $T=550$, and $2000$ epochs in total.
+Rectified adaptive moment estimation (RADAM) Liu et al.
+[2020] was employed as the optimizer, with $L=50$ batches of $S_{B}=16$ sequences per epoch, each of length $T=550$, and $2000$ epochs in total.
 We used a learning rate exponentially decaying from $\eta_{\text{start}}=5\cdot 10^{-3}$ to $\eta_{\text{end}}=10^{-5}$.
@@
 At test time, no retraining or fine-tuning is performed, but utilizing the context $\bm{C}$ the model just forward-iterates from the last context time step.
-Hyperparameters for training were partly selected according to previous results in [7, 6], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.1,0.2\}$, $\Delta t=\{0,50,80\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
+Hyperparameters for training were partly selected according to previous results in Brenner et al.
+[2024a, 2022], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.1,0.2\}$, $\Delta t=\{0,50,80\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
 Model details
@@
 Each expert has a latent dimension of $M=30$, of which $P=2$ are rectified-linear units (ReLUs).
-As suggested in [7], we only use the first $N$ of the MPM-P linear units for the readout.
-We followed the initialization protocol in [6, 81], drawing $\bm{W}$ from a Gaussian with mean $\bm{0}$ and $\bm{\sigma}=0.01\mathds{1}$, setting $\bm{h}=\mathbf{0}$, and $\bm{A}$ to the diagonal of a normalized positive-definite random matrix.
+As suggested in Brenner et al.
+[2024a], we only use the first $N$ of the MPM-P linear units for the readout.
+We followed the initialization protocol in Brenner et al.
+[2022], Talathi and Vartak [2016], drawing $\bm{W}$ from a Gaussian with mean $\bm{0}$ and $\bm{\sigma}=0.01\mathds{1}$, setting $\bm{h}=\mathbf{0}$, and $\bm{A}$ to the diagonal of a normalized positive-definite random matrix.
 The initial latent state was estimated as
@@
 We initialize the temperature weights $\tau_{\text{att}}$ and $\tau_{\text{exp}}$ to $0.1$, the covariance matrix $\bm{\Sigma}=0.05\cdot\mathds{1}$, matrix $\bm{D}=\mathcal{I}_{(N\times M)}$, and draw all other CNN and MLP matrices from a Gaussian with standard deviation $0.01$.
-Model hyperparameters were selected according to previous results in [7] and by probing a few relevant parameter settings, namely $J=\{1,3,5,10,20,30\}$ (Fig.
+Model hyperparameters were selected according to previous results in Brenner et al.
+[2024a] and by probing a few relevant parameter settings, namely $J=\{1,3,5,10,20,30\}$ (Fig.
 25), $\text{CNN-layers}=\{1,2\}$, $\text{CNN-kernel-size}=\{2,3,5\}$, $\text{MLP-layers}=\{1,2,3\}$, and different activation functions $\{\mathds{1},\tanh{},\text{ReLU}\}$ for the MLP/CNN.
@@
 Performance measures
-As in [48, 58, 39, 8, 64, 30, 99], we assess the geometric (dis)similarity between true and model-generated attractors using a Kullback-Leibler divergence ($D_{\text{stsp}}$) defined across state space.
+As in Koppe et al.
+[2019], Mikhaeil et al.
+[2022], Hess et al.
+[2023], Brenner et al.
+[2024b], Pals et al.
+[2024], Gilpin [2023], Zhang and Gilpin [2025], we assess the geometric (dis)similarity between true and model-generated attractors using a Kullback-Leibler divergence ($D_{\text{stsp}}$) defined across state space.
 Specifically, $D_{\text{stsp}}$ determines the overlap between the distributions of true trajectory points, $p_{true}(\bm{x})$, and of model-generated trajectories, $p_{gen}(\bm{x}|\bm{z})$, given by
 $D_{\text{stsp}}(p_{true}(\bm{x})\lVert p_{gen}(\bm{x}|\bm{z}))=\int p_{true}(\bm{x})\log\frac{p_{true}(\bm{x})}{p_{gen}(\bm{x}|\bm{z})}d\bm{x}.$ (11)
-While especially for higher-dimensional state spaces these distributions may be approximated by Gaussian mixture models [48], here a discretized version was sufficient, with the state space parcellated into $K=m^{N}$ bins, where $m$ is the number of bins per dimension and $N$ the dimensionality of the system.
+While especially for higher-dimensional state spaces these distributions may be approximated by Gaussian mixture models Koppe et al.
+[2019], here a discretized version was sufficient, with the state space parcellated into $K=m^{N}$ bins, where $m$ is the number of bins per dimension and $N$ the dimensionality of the system.
 We estimate the occupation probabilities $p_{i}$ of each bin via the relative frequencies $\hat{p}_{i}$ of trajectory visits and approximate $D_{\text{stsp}}$ as
 $D_{\text{stsp}}\approx\sum_{i=1}^{K}\hat{p}_{true;i}\log\frac{\hat{p}_{true;i}}{\hat{p}_{gen;i}}.$ (12)
-We set $m=30$ bins per dimension, following [38].
+We set $m=30$ bins per dimension, following Hemmer et al.
+[2024].
 For the $6d$ Lorenz-96 we use $m=5$ and for the empirical time series $m=20$.
 To ensure a steady-state distribution is reached, long trajectories of $T=10,000$ time steps are sampled from the DS.
-To evaluate the agreement in long-term temporal dynamics, we compute the Hellinger distance $D_{H}$ between the power spectra of the true and model-generated time series [58, 39], defined as
+To evaluate the agreement in long-term temporal dynamics, we compute the Hellinger distance $D_{H}$ between the power spectra of the true and model-generated time series Mikhaeil et al.
+[2022], Hess et al.
+[2023], defined as
 $H(F(\omega),G(\omega))=\sqrt{1-\int_{-\infty}^{\infty}\sqrt{F(\omega)G(\omega)}d\omega}\;\in[0,1],$ (13)
@@
 Power spectra are obtained via the dimension-wise Fast Fourier Transform (FFT), smoothed using a Gaussian kernel (with $\sigma=20$ for DS and $\sigma=2$ for the short empirical time series), and normalized to ensure comparability.
-High-frequency tails dominated by noise are truncated as described in [39].
+High-frequency tails dominated by noise are truncated as described in Hess et al.
+[2023].
 The aggregated Hellinger distance $D_{H}$ is then computed as the average across all dimension-wise spectral comparisons.
-For assessing short-term forecast quality aggregated across the whole test set of $54$ DS, we used a normalized $n$-step ahead prediction error (with $n=10$ in our evaluation) as recommended in [40], given by
+Note that both $D_{\text{stsp}}$ and $D_{H}$ are assumed to assess (dis-)agreement in long-term properties in the limit $T\rightarrow\infty$.
+Hence, they are only sensible on rather long-term horizons, with $T=10,000$ used here.
+For assessing short-term forecast quality aggregated across the whole test set of $54$ DS, we used a normalized $n$-step ahead prediction error (with $n=10$ in our evaluation) as recommended in Hewamalage et al.
+[2023], given by
 $\text{MASE}=\frac{1}{N}\sum_{i=1}^{N}\frac{\dfrac{1}{n}\sum_{t=T_{C}+1}^{T_{C}+n}\left|x_{i,t}-\hat{x}_{i,t}\right|}{\dfrac{1}{T}\sum_{t=T_{C}+1}^{T_{C}+T}\left|x_{i,t}-x_{i,t-1}\right|}$ (14)
-For individual comparisons on single empirical times, the MAE for evaluating short term predictions [40] was used:
+For individual comparisons on single empirical times, the MAE for evaluating short term predictions Hewamalage et al. [2023] was used:
 $\text{MAE}=\frac{1}{n}\sum_{t=T_{C}+1}^{T_{C}+n}\left|x_{t}-\hat{x}_{t}\right|\;,$ (15)
@@
 Lyapunov exponent
-For numerically estimating the maximum Lyapunov exponent $\lambda_{\text{max}}$ from a one-dimensional empirical time series $\{x_{t}\},t=1\dots T$, one common approach is the Rosenstein algorithm [72].
+For numerically estimating the maximum Lyapunov exponent $\lambda_{\text{max}}$ from a one-dimensional empirical time series $\{x_{t}\},t=1\dots T$, one common approach is the Rosenstein algorithm Rosenstein et al.
+[1993].
 First, the time series is embedded into a state space using time-delay embedding as in eq.
 5 (e.g.
-using the PECUZAL algorithm [49]).
-For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity [72]) and $l_{s}=0.25$.
+using the PECUZAL algorithm Kraemer et al.
+[2021]).
+For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity Rosenstein et al.
+[1993]) and $l_{s}=0.25$.
 The local divergence between initially nearby trajectories is then tracked over a range of time steps $k=0\dots k_{max}$,
@@
 $\langle\ln d(k)\rangle\approx\lambda_{\max}k\Delta t+C.$ (18)
-In chaotic systems, for a suitable choice of $k_{max}$ (here: $50$), this quantity initially grows approximately linearly with time [44], with the slope given by $\lambda_{\max}>0$.
+In chaotic systems, for a suitable choice of $k_{max}$ (here: $50$), this quantity initially grows approximately linearly with time Kantz and Schreiber [2004], with the slope given by $\lambda_{\max}>0$.
 In contrast, periodic (cyclic) systems have $\lambda_{\max}=0$.
+As for $D_{\text{stsp}}$ and $D_{H}$, Lyapunov exponents can only be reasonably assessed for sufficiently long time series, with practical guidelines of $T\approx 10^{d}-30^{d}$, where $d$ is the attractor’s fractal dimension Wolf et al.
+[1985], Rosenstein et al.
+[1993].
 A.2 Datasets
 Training data
-DynaMix is trained on about $0.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in [29].
+DynaMix is trained on about $0.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in Gilpin [2022].
 Time series were standardized to ensure comparable scaling of all data, and Gaussian noise of $5\%$ of the data standard deviation was added to all dimensions for the ground truth.
@@
 Test data
-Our test set for DS consists of simulated time series of length $10^{5}$ sampled from $54$ different $3d$ DS collected in [29], which are not part of the training set.
-Furthermore, we evaluated several $2d$ systems: The Selkov system [77] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
+Our test set for DS consists of simulated time series of length $10^{5}$ sampled from $54$ different $3d$ DS collected in Gilpin [2022], which are not part of the training set.
+Furthermore, we evaluated several $2d$ systems: The Selkov system SEL’KOV [1968] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
 $\frac{dx}{dt}$ $=-x+ay+x^{2}y,$ (19)
 $\frac{dy}{dt}$ $=b-ay-x^{2}y,$
-where we chose $a=0.1$ and $b=0.5$. The Van-der-Pol system [87] describes self-sustaining oscillations in vacuum tubes by
+where we chose $a=0.1$ and $b=0.5$. The Van-der-Pol system Van der Pol [1926] describes self-sustaining oscillations in vacuum tubes by
 $\frac{dx}{dt}$ $=y,$ (20)
 $\frac{dy}{dt}$ $=\mu(1-x^{2})y-x,$
-where we chose $\mu=0.5$. Similar to the Van-der-Pol system, the Rayleigh oscillator [61] describes a self-sustained nonlinear oscillator through
+where we chose $\mu=0.5$. Similar to the Van-der-Pol system, the Rayleigh oscillator Nayfeh and Mook [2024] describes a self-sustained nonlinear oscillator through
 $\frac{dx}{dt}$ $=y,$ (21)
 $\frac{dy}{dt}$ $=\mu(1-\frac{y^{2}}{3})y-x,$
-where we chose $\mu=1.0$. As a higher dimensional DS test case we use the Lorenz-96 system [54] defined by
+where we chose $\mu=1.0$. As a higher dimensional DS test case we use the Lorenz-96 system Lorenz [1996] defined by
 $\frac{\text{d}x_{i}}{\text{d}t}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,$ (22)
@@
 We further probed forecasting and DSR on different types of real-world data: The traffic data are hourly recordings of the number of cars passing road junctions (https://www.kaggle.com/datasets/fedesoriano/traffic-prediction-dataset/data).
-The cloud data, also used in [84] to evaluate TS foundation models, are publicly available from Huawei Cloud (https://github.com/sir-lab/data-release).
+The cloud data, also used in Toner et al.
+[2025] to evaluate TS foundation models, are publicly available from Huawei Cloud (https://github.com/sir-lab/data-release).
 It consists of function requests from Huawei’s serverless cloud platform.
 The weather data consists of daily sampled soil temperature and air pressure measured in the city of Mannheim, Germany (source: Deutscher Wetterdienst [German weather service]), and can be accessed via https://www.dwd.de/EN/ourservices/cdc/cdc_ueberblick-klimadaten_en.html.
-The functional magnetic resonance imaging (fMRI) data comes from human subjects performing cognitive tasks and is publicly available on GitHub [50].
+The functional magnetic resonance imaging (fMRI) data comes from human subjects performing cognitive tasks and is publicly available on GitHub Kramer et al.
+[2022].
 We followed Kramer et al.
@@
 Kramer et al.
-[50] report a positive maximum Lyapunov exponent for models reconstructed from these time series, indicating their chaotic nature (see also [89]).
+[2022] report a positive maximum Lyapunov exponent for models reconstructed from these time series, indicating their chaotic nature (see also Volkmann et al.
+[2024]).
 The ETTh1 dataset is part of the Electricity Transformer Temperature (ETT) benchmark, which is widely used for evaluating TS forecasting models.
-It contains hourly data collected from a power transformer station [100] and can be accessed at https://github.com/zhouhaoyi/ETDataset.
+It contains hourly data collected from a power transformer station Zhou et al.
+[2021] and can be accessed at https://github.com/zhouhaoyi/ETDataset.
 Electroencephalogram (EEG) data were taken from a study by Schalk et al.
@@
 Chronos
-Chronos is a recent TS foundation framework that adapts a transformer-based LLM architecture for probabilistic time series forecasting [2].
+Chronos is a recent TS foundation framework that adapts a transformer-based LLM architecture for probabilistic time series forecasting Ansari et al.
+[2024].
 Key to this approach is the tokenization of real-valued time series observations through scaling and quantization, transforming them into token sequences accessible to language models.
-The model is pretrained on extensive datasets, including both synthetic data generated via Gaussian processes, as well as an extensive batch of empirical time series from [2, 32], including traffic, weather/ climate, electricity and web data, enabling it to achieve strong zero-shot performance across diverse datasets without task-specific fine-tuning.
+The model is pretrained on extensive datasets, including both synthetic data generated via Gaussian processes, as well as an extensive batch of empirical time series from Ansari et al.
+[2024], Godahewa et al.
+[2021], including traffic, weather/ climate, electricity and web data, enabling it to achieve strong zero-shot performance across diverse datasets without task-specific fine-tuning.
 For our evaluation we used the standard pipeline as described in https://github.com/amazon-science/chronos-forecasting.
 TimesFM
-TimesFM is another transformer based TS foundation model, with a decoder-only style architecture using input patching [20].
+TimesFM is another transformer based TS foundation model, with a decoder-only style architecture using input patching Das et al.
+[2024].
 Its training corpus consists of synthetic as well as real-world time series data, and it exhibits generalization across different time series domains and context lengths.
@@
 Mamba4Cast
-Mamba4Cast is a zero-shot time series forecasting model based on the Mamba architecture [35], a type of linear (‘state-space’) RNN with nonlinear input & output gating, and inspired by Prior-data Fitted Networks (PFNs) [5].
+Mamba4Cast is a zero-shot time series forecasting model based on the Mamba architecture Gu and Dao [2023], a type of linear (‘state-space’) RNN with nonlinear input & output gating, and inspired by Prior-data Fitted Networks (PFNs) Bhethanabhotla et al.
+[2024].
 Trained exclusively on synthetic data, Mamba4Cast can generate zero-shot forecasts when provided with time series context information.
@@
 Tiny Time Mixers
-Like Mamba4Cast, Tiny Time Mixers is a time series foundation model not founded on transformers [26].
-It builds upon the TSMixer backbone [25] to which it adds several key innovations such as adaptive patching, diverse resolution sampling, and multi-resolution prefix tuning to improve generalization across datasets with varying temporal resolutions.
+Like Mamba4Cast, Tiny Time Mixers is a time series foundation model not founded on transformers Ekambaram et al.
+[2024].
+It builds upon the TSMixer backbone Ekambaram et al.
+[2023] to which it adds several key innovations such as adaptive patching, diverse resolution sampling, and multi-resolution prefix tuning to improve generalization across datasets with varying temporal resolutions.
 Evaluation here is performed as in https://github.com/glehet/TTM1.
+Panda
+Panda is a recently proposed foundation model for short-term forecasting of DS Lai et al.
+[2025].
+It is based on a transformer architecture, which relies on patching the DS-generated time series.
+The model is trained on $2\cdot 10^{4}$ DS produced by combining base DS from the same database used to train and evaluate DynaMix in skew-product form Gilpin [2022].
+Evaluation is performed as in https://github.com/abao1999/panda.
 For comparability, all models were evaluated on the exact same CPU (18-Core Xeon Gold 6254) and GPU (Nvidia RTX 2080 Ti) using 512GB of RAM.
@@
 Numerical details in Table LABEL:tab:performance_DSR_CL512.
+Statistical testing based on Wilcoxon signed-rank tests.
 Figure 24: Exponential divergence of two initially nearby trajectories (in blue and red) illustrated for the chaotic Lorenz-63 system.
@@
 Error bands = STD
-Figure 26: DSR performance for different teacher forcing intervals $\tau$, illustrating an optimal value is essential for successful training, see [58, 6]. Error bands = STD
+Figure 26: DSR performance for different teacher forcing intervals $\tau$, illustrating an optimal value is essential for successful training, see Mikhaeil et al.
+[2022], Brenner et al.
+[2022].
+Error bands = STD
 Figure 27: Comparison of DynaMix (red) to Chronos (blue) on zero-shot forecasts of various empirical time series: Forecasts of a) partially (11d) observed Lorenz-63 DS, b) electricity transformer temperature data, c) hourly car traffic data with weekly cycle, d) Huawei cloud request data, e) soil temperature development, f) air pressure data, g) human functional magnetic resonance imaging (fMRI) data, h) human electroencephalogram (EEG) data.
@@
 A.5 Using DynaMix for non-stationary time series
-Similar as in FEDformers [101] and related TS models which use specialized decomposition and filtering blocks for handling non-stationary data, one could add simple preprocessing operations to the DynaMix pipeline to separate out trend or other non-stationary components in the context signal.
+Similar as in FEDformers Zhou et al.
+[2022] and related TS models which use specialized decomposition and filtering blocks for handling non-stationary data, one could add simple preprocessing operations to the DynaMix pipeline to separate out trend or other non-stationary components in the context signal.
 For a simple illustration, here we first apply a Box-Cox transformation to time series $x_{t}$,
@@
 A.6 Scaling up context dimension
-For reconstructing DS higher-dimensional than the context dimension defined by the architecture and used in training, one can use DynaMix without any modification by employing the delay-embedding theorems [80, 74]: Zero-shot infer the underlying DS from the observed TS and then delay-embed DynaMix’ output into a sufficiently high-dimensional space which assures a diffeomorphism between original and reconstructed attractor.
+For reconstructing DS higher-dimensional than the context dimension defined by the architecture and used in training, one can use DynaMix without any modification by employing the delay-embedding theorems Takens [1981], Sauer et al.
+[1991]: Zero-shot infer the underlying DS from the observed TS and then delay-embed DynaMix’ output into a sufficiently high-dimensional space which assures a diffeomorphism between original and reconstructed attractor.
 This idea is illustrated for the $6d$ Lorenz-96 system in Fig.
@@
 Training was performed as in Brenner et al.
-[7], and hyperparameter settings, as collected in Table 3, followed those in the original paper.
+[2024a], and hyperparameter settings, as collected in Table 3, followed those in the original paper.
 Table 3: Hyperparameter settings of custom AL-RNN.
@@
 epochs 2000
-As another SOTA DSR model we trained Neural-ODEs [17], using the same type of MLP architecture, training setup, and same hyper-parameter settings as in [37], see Table 4.
+As another SOTA DSR model we trained Neural-ODEs Chen et al.
+[2018], using the same type of MLP architecture, training setup, and same hyper-parameter settings as in Göring et al.
+[2024a], see Table 4.
 With this, the total number of trainable parameters is already more than twice (!) than that used for DynaMix ($\approx 10$k), and exceeded the number of data points ($\leq 4$k) manifold.
@@
 Here we employed the architecture provided in Patel and Ott [2022], trained as in Göring et al.
-[37] and using the same hyper-parameter settings, summarized in Table 5.
+[2024a] and using the same hyper-parameter settings, summarized in Table 5.
 Note that this endows RCs with about $25\times$ more parameters in total, of which $\leq 3$k were trainable, than DynaMix had at its disposal.
@@
 Human EEG 0.79 0.23 1.07 0.23 0.44 0.70 1.80 0.55 0.89 1.23 0.26 0.91
-Table 9: Performance comparison for zero-shot inference across all DS in the dataset from [29] using $T_{C}=2000$. Best value in red, second best in blue.
-System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM TimesFM
+Table 9: Performance comparison for zero-shot inference across all DS in the dataset from Gilpin [2022] using $T_{C}=2000$. Best value in red, second best in blue.
+System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-2 Panda Mamba4Cast TTM TimesFM
 $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
-Atmospheric Regime 10.4 nan 0.0 9.2 0.4 0.0 10.5 0.4 0.0 12.0 0.5 0.0 13.3 0.5 0.0 13.7 0.5 0.1 11.6 0.9 0.1 12.3 0.8 0.2 12.3 0.9 0.1 11.7 0.7 0.2 11.2 0.8 0.7 13.9 0.4 1.6 11.8 0.4 0.1
-BeerRNN 5.2 0.7 0.6 12.1 0.5 0.5 9.3 0.4 0.7 13.3 0.4 0.7 10.9 0.3 0.6 13.1 0.5 0.8 12.4 0.9 1.0 12.7 0.8 1.7 11.9 0.9 1.5 11.7 0.8 1.6 11.4 0.8 1.0 14.2 0.4 1.5 14.0 0.3 0.5
-Bouali 5.9 0.1 0.2 7.0 0.3 0.4 12.1 0.2 0.5 11.6 0.1 0.5 13.1 0.2 0.4 13.5 0.4 2.0 11.2 0.7 2.3 12.0 0.6 3.1 12.3 0.5 2.0 9.8 0.5 2.2 12.9 0.7 4.4 14.0 0.3 5.6 9.7 0.4 1.1
-Burke Shaw 3.8 0.2 0.3 8.0 0.1 0.5 5.6 0.1 0.4 5.1 0.2 0.2 11.1 0.2 0.5 13.2 0.6 1.1 11.0 0.6 1.3 11.0 0.7 0.9 11.7 0.8 0.9 10.0 0.7 1.2 10.2 0.7 1.2 12.9 0.4 2.1 12.4 0.3 0.8
-Ca Two Plus Quasip. 2.8 0.1 0.3 9.6 0.5 1.6 9.5 0.3 0.5 12.6 0.2 1.2 12.7 0.4 1.4 13.2 0.4 1.1 9.4 0.7 2.6 9.4 0.6 1.3 9.6 0.7 1.4 11.5 0.5 2.4 11.1 0.6 4.0 13.8 0.2 4.6 13.4 0.3 1.9
-Chen Lee 3.6 0.3 0.2 4.2 0.2 0.0 3.6 0.2 0.1 3.9 0.2 0.1 5.3 0.3 0.1 12.0 0.8 0.5 4.5 0.8 0.4 8.8 0.7 0.9 7.7 0.7 0.6 7.2 0.7 1.4 11.3 0.8 0.8 8.9 0.4 2.0 4.6 0.2 0.3
-Chua 2.9 0.1 0.8 12.9 0.2 0.6 12.4 0.4 0.8 13.3 0.2 0.2 13.5 0.4 1.0 13.5 0.3 2.1 11.9 0.6 2.4 11.7 0.4 1.3 11.7 0.4 1.8 12.0 0.4 2.2 12.1 0.4 0.8 13.4 0.3 4.0 12.7 0.2 1.0
-Coullet 0.9 0.0 0.2 0.6 0.0 0.3 6.3 0.0 0.2 3.9 0.1 0.2 13.7 0.3 0.2 14.6 0.3 0.8 11.4 0.7 1.1 12.8 0.7 1.5 11.8 0.6 1.1 10.7 0.6 3.0 13.3 0.6 1.5 15.0 0.2 3.5 9.3 0.2 0.5
-Dadras 2.2 0.2 1.3 10.0 0.5 4.2 9.2 0.5 1.9 11.8 0.4 2.3 11.4 0.3 3.5 12.9 0.8 4.1 10.5 0.8 5.4 9.4 0.7 5.0 11.0 0.6 6.8 8.8 0.5 5.6 13.1 0.8 5.0 12.2 0.4 6.0 13.1 0.6 4.5
-Dequan Li 8.5 0.1 0.8 12.6 0.5 0.2 13.2 0.4 0.3 13.1 0.3 0.3 12.9 0.4 1.4 13.1 0.3 1.0 12.4 0.4 0.8 12.7 0.4 0.8 12.2 0.4 0.6 12.6 0.4 1.4 12.3 0.4 3.3 13.0 0.4 1.7 13.0 0.3 1.3
-Finance 3.1 0.1 0.3 8.5 0.4 0.7 9.4 0.6 0.5 13.4 0.3 0.6 12.2 0.4 0.9 13.6 0.5 1.1 12.0 0.7 2.8 11.7 0.6 4.0 12.4 0.7 3.7 11.8 0.6 4.7 12.6 0.8 4.9 14.0 0.3 4.2 13.2 0.5 2.7
-Genesio Tesi 9.8 0.2 0.4 8.9 0.1 0.2 8.7 0.1 0.1 7.7 0.1 0.1 11.8 0.4 0.4 12.3 0.3 0.4 10.7 0.3 0.8 11.4 0.9 1.4 10.6 0.6 2.0 10.0 0.4 2.8 10.4 0.8 1.3 12.3 0.2 1.7 12.2 0.2 1.1
-Guckenheimer Holmes 7.4 0.0 0.2 10.6 0.4 0.3 11.7 0.2 0.3 11.3 0.4 0.4 13.2 0.7 0.6 12.5 0.2 0.6 10.7 0.6 0.7 10.6 0.6 1.0 9.2 0.6 0.4 10.3 0.6 0.8 10.4 0.5 2.0 11.6 0.3 3.2 11.0 0.2 0.7
-Hadley 8.6 0.3 0.3 10.8 0.5 0.2 9.3 0.6 0.4 11.9 0.4 0.4 12.5 0.4 0.6 12.5 0.3 0.9 10.9 0.7 0.6 10.7 0.7 2.0 9.5 0.6 0.7 10.2 0.4 1.8 11.7 0.7 1.7 12.5 0.3 5.4 11.7 0.3 0.4
-Halvorsen 1.8 0.1 0.2 6.0 0.1 0.3 8.4 0.1 0.5 5.9 0.1 0.3 12.6 0.2 0.5 13.3 0.3 1.4 11.8 0.8 1.6 9.5 0.7 2.1 8.9 0.8 1.8 10.4 0.5 2.3 12.2 0.6 3.4 13.2 0.3 2.8 12.8 0.2 0.8
-Hindmarsh Rose 6.1 0.5 1.1 9.7 0.3 0.3 6.7 0.5 0.7 8.4 0.1 0.3 11.5 0.4 0.9 13.8 0.6 1.5 12.7 0.7 1.5 12.5 0.7 2.2 13.9 0.6 1.2 13.4 0.7 2.2 13.9 0.6 2.5 14.0 0.2 2.7 14.2 0.6 1.5
-Isothermal Chemical 2.4 0.2 0.1 3.7 0.1 0.0 7.9 0.3 0.0 9.7 0.3 0.0 9.0 0.3 0.1 13.8 0.5 0.4 7.9 0.8 0.5 5.9 0.8 0.4 8.4 0.8 0.6 8.5 0.8 0.4 12.8 0.8 0.7 15.6 0.2 6.1 4.0 0.4 0.3
-Itik Banks tumor 8.7 0.2 7.0 10.0 0.6 1.7 8.4 0.5 2.4 10.5 0.4 1.5 12.8 0.3 0.8 12.1 0.7 3.2 9.2 0.7 3.6 9.5 0.7 4.6 9.7 0.8 5.1 9.6 0.7 5.1 13.3 0.9 6.4 13.2 0.3 8.2 12.7 0.3 3.8
-Jerk Circuit 3.2 0.4 3.3 10.1 0.4 1.3 9.7 0.4 1.0 7.5 0.3 0.5 11.1 0.4 0.5 12.3 0.3 0.5 10.8 0.8 1.5 10.9 0.5 2.1 10.5 0.9 1.7 10.1 0.9 1.9 11.1 0.9 1.6 13.9 0.4 2.8 12.6 0.3 1.4
-Laser 8.8 0.4 0.2 11.1 0.4 0.1 10.8 0.2 0.1 12.8 0.2 0.3 13.8 0.5 0.6 14.1 0.5 0.5 11.8 0.7 0.4 11.5 0.7 0.5 11.0 0.7 0.8 11.9 0.7 2.7 13.3 0.6 1.2 13.8 0.2 2.2 12.8 0.4 0.4
-Lorenz 1.2 0.1 0.2 6.6 0.2 0.4 6.0 0.2 0.3 10.0 0.2 0.3 12.0 0.4 0.3 13.0 0.5 0.3 10.9 0.7 0.5 11.3 0.5 0.5 12.7 0.6 1.1 12.8 0.3 3.0 11.4 0.7 1.2 13.0 0.3 2.2 13.1 0.5 0.8
-Lorenz 84 6.9 0.4 1.5 9.5 0.4 2.2 11.0 0.4 1.9 11.5 0.4 1.3 12.2 0.4 1.7 12.9 0.5 3.5 12.9 0.8 2.4 12.9 0.8 2.7 11.5 0.9 2.5 12.3 0.7 2.9 13.4 0.8 2.5 13.4 0.5 2.4 13.4 0.5 2.0
-Lu Chen 7.7 0.8 0.2 6.3 0.4 0.2 5.3 0.3 0.2 8.0 0.1 0.1 7.1 0.1 0.1 4.9 0.2 0.1 11.6 0.9 0.7 11.9 1.0 0.8 11.3 1.0 0.9 11.4 1.0 0.8 9.7 0.8 0.9 13.7 0.4 1.5 8.6 0.5 0.4
-Moore Spiegel 4.3 0.2 0.3 10.4 0.6 0.5 10.0 0.5 0.1 12.5 0.4 0.2 9.9 0.5 0.2 12.1 0.3 0.6 10.5 0.9 1.5 11.4 0.9 1.6 10.1 0.9 1.6 9.8 0.7 2.1 9.4 0.8 2.8 12.5 0.3 2.4 12.4 0.3 2.8
-Multi Chua 8.1 0.3 0.8 9.3 0.5 0.3 9.3 0.2 0.5 7.1 0.4 0.4 13.3 0.2 1.1 14.0 0.3 3.9 11.5 0.6 5.8 12.2 0.4 11.5 9.5 0.5 11.3 10.1 0.4 9.0 10.6 0.4 4.8 14.0 0.3 7.6 11.4 0.2 2.0
-Nose Hoover 9.5 0.2 0.1 9.9 0.5 0.1 9.6 0.6 0.1 11.5 0.3 0.5 12.5 0.7 0.8 13.0 0.5 0.5 10.6 0.9 0.6 10.6 0.9 0.6 10.1 0.7 0.8 10.2 0.5 1.0 10.2 0.7 1.1 12.9 0.4 1.6 13.3 0.3 0.9
-Pehlivan Wei 9.2 0.0 0.2 8.5 0.4 0.5 10.1 0.4 0.5 11.3 0.3 0.3 10.8 0.4 0.5 12.6 0.5 0.7 9.2 0.7 1.2 10.1 0.8 1.7 9.2 0.7 1.0 10.6 0.8 2.1 9.3 0.7 1.4 12.2 0.6 3.0 8.2 0.3 1.2
-Sakarya 2.0 0.2 0.7 5.6 0.5 0.3 5.0 0.3 0.3 7.5 0.2 0.3 11.1 0.4 0.4 12.5 0.4 0.4 10.2 0.9 1.0 10.0 0.6 1.6 7.7 0.9 2.2 9.1 0.4 1.3 13.1 0.8 1.5 12.6 0.4 1.2 11.9 0.3 2.3
-San Um Srisuchinwong 3.6 0.2 0.3 12.9 0.4 1.2 12.7 0.5 2.1 11.4 0.5 1.4 13.3 0.5 2.5 13.0 0.6 1.9 12.7 0.7 1.5 10.4 0.7 2.1 13.4 0.6 1.8 11.7 0.7 1.9 10.7 0.7 2.5 13.1 0.4 3.7 13.3 0.7 1.8
-Sprott A 2.7 0.1 0.3 8.0 0.2 0.3 8.9 0.3 0.3 6.6 0.2 0.3 11.8 0.3 0.3 12.3 0.3 0.9 10.5 0.8 0.9 8.1 0.5 1.6 10.0 0.7 1.5 10.4 0.7 1.6 9.9 0.7 2.5 12.0 0.3 5.7 9.8 0.2 1.0
-Sprott B 2.9 0.2 0.4 4.3 0.2 1.1 5.6 0.4 0.9 11.7 0.5 0.7 12.7 0.4 0.8 13.2 0.7 1.6 11.8 0.8 1.3 11.5 0.6 0.8 12.3 0.8 1.2 10.6 0.7 2.9 12.5 0.8 3.0 12.6 0.4 2.1 13.5 0.6 0.9
-Sprott C 2.4 0.1 0.4 11.6 0.4 2.3 12.4 0.4 1.3 11.9 0.4 2.3 13.3 0.3 1.1 13.1 0.4 4.6 10.0 0.5 7.7 9.2 0.5 6.8 10.7 0.5 8.0 9.3 0.5 7.6 12.8 0.6 5.2 13.3 0.2 6.2 13.4 0.2 4.5
-Sprott D 3.0 0.1 0.4 9.8 0.6 0.3 10.3 0.4 0.3 12.9 0.3 0.4 13.4 0.5 0.7 13.4 0.4 0.3 12.2 0.7 0.8 10.9 0.6 0.6 9.9 0.5 1.2 11.5 0.6 0.9 12.9 0.7 1.3 14.0 0.2 4.0 7.9 0.5 0.9
-Sprott F 5.0 0.3 0.7 9.6 0.5 0.2 9.2 0.5 0.6 12.3 0.3 0.4 12.6 0.4 1.0 12.7 0.4 0.7 9.9 0.7 1.3 10.4 0.7 1.5 10.4 0.6 0.9 12.1 0.5 1.6 11.4 0.8 1.6 13.3 0.3 2.8 13.0 0.6 1.7
-Sprott G 3.8 0.1 0.3 7.0 0.6 0.2 7.7 0.5 0.2 10.0 0.3 0.4 7.2 0.1 0.6 13.0 0.3 0.6 10.9 0.9 1.9 11.4 0.6 1.9 9.2 0.8 1.8 10.8 0.8 2.4 14.1 0.8 1.4 13.1 0.2 1.9 9.5 0.4 1.4
-Sprott I 2.7 0.1 0.5 9.7 0.5 0.7 11.3 0.2 0.6 9.0 0.1 0.5 11.9 0.3 0.6 13.7 0.2 1.0 10.2 0.9 3.5 9.8 0.9 4.9 8.4 0.6 4.6 9.1 0.6 4.1 12.2 0.7 2.0 14.4 0.2 8.8 8.3 0.4 1.0
-Sprott L 3.6 0.1 0.6 8.0 0.5 0.3 6.8 0.1 0.3 8.2 0.2 0.1 11.6 0.3 0.4 13.0 0.2 0.3 10.5 0.9 1.8 10.4 0.8 2.8 11.4 0.9 1.8 12.1 0.7 1.6 10.8 0.8 1.3 14.2 0.3 2.9 8.1 0.4 0.2
-Sprott M 3.6 0.1 0.2 9.8 0.4 0.2 10.5 0.4 0.5 12.4 0.1 0.3 12.9 0.2 0.3 14.0 0.6 0.9 9.8 0.8 3.7 10.3 0.6 3.7 9.9 0.8 3.3 11.8 0.6 3.1 11.6 0.7 2.9 14.3 0.2 6.5 13.4 0.4 1.6
-Sprott O 2.0 0.1 0.4 3.6 0.1 0.4 4.6 0.0 0.4 9.7 0.1 0.5 12.4 0.4 0.6 12.9 0.4 1.3 10.8 1.0 1.6 12.6 0.9 2.6 10.4 0.8 1.8 11.5 0.9 1.4 10.7 0.8 1.6 14.0 0.4 3.7 9.5 0.5 0.7
-Sprott P 3.8 0.2 1.4 10.8 0.3 3.9 11.9 0.3 2.6 12.7 0.3 2.6 12.7 0.3 2.1 13.3 0.3 3.8 7.3 0.5 5.7 10.6 0.6 9.6 8.3 0.6 8.7 10.0 0.6 10.7 11.1 0.7 6.3 12.2 0.3 10.2 11.8 0.3 4.4
-Sprott Q 4.5 0.1 0.2 10.1 0.2 0.3 7.6 0.1 0.4 7.3 0.1 0.5 11.6 0.2 0.4 13.2 0.3 1.3 11.8 0.8 3.6 11.6 0.7 1.9 11.9 0.6 1.0 11.2 0.6 4.1 10.9 0.7 1.1 13.1 0.4 7.5 8.8 0.4 0.4
-Sprott R 4.3 0.3 0.2 9.3 0.8 0.2 6.7 0.5 0.3 11.0 0.5 0.5 12.6 0.5 0.6 13.1 0.4 0.4 10.6 0.8 0.5 10.4 0.5 1.2 9.6 0.7 2.1 10.5 0.7 1.8 11.8 0.8 1.3 13.8 0.3 2.5 11.3 0.3 0.9
-Sprott S 4.2 0.1 0.3 10.1 0.9 2.2 9.9 0.6 1.1 12.3 0.2 1.2 12.6 0.4 0.6 13.1 0.3 1.5 10.6 0.7 1.2 9.7 0.8 6.1 11.0 0.6 3.1 10.4 0.4 7.2 11.8 0.8 4.1 11.6 0.3 10.2 12.8 0.2 3.7
-Sprott Torus 2.2 0.2 0.2 5.4 0.7 0.4 10.4 0.7 0.3 14.3 0.6 0.3 14.5 0.6 0.6 13.8 0.7 0.4 6.2 0.8 4.2 8.3 0.6 8.3 5.3 0.8 6.3 5.8 0.7 6.3 13.6 0.8 2.4 15.5 0.3 4.8 14.3 0.5 1.0
-Thomas 1.6 0.3 4.8 10.6 0.5 6.0 10.1 0.8 6.0 9.9 0.3 5.7 13.9 0.5 4.7 13.9 0.7 5.6 12.6 0.8 3.4 12.6 0.8 1.9 12.8 0.8 2.8 12.9 0.8 2.5 10.7 0.5 6.1 13.3 0.2 3.1 13.8 0.6 3.3
-Thomas Labyrinth 4.1 0.1 2.4 12.5 0.2 2.6 12.5 0.3 2.8 12.4 0.4 3.0 12.5 0.4 2.8 12.5 0.3 3.8 11.4 0.5 3.2 12.5 0.5 2.6 11.9 0.5 3.0 11.7 0.5 3.1 12.3 0.5 3.6 12.4 0.2 2.0 12.6 0.5 2.5
-Torus 11.9 0.9 1.1 9.8 0.3 0.1 9.4 0.2 0.1 10.9 0.3 0.1 12.3 0.3 0.2 12.4 0.5 0.3 12.1 0.9 0.8 10.5 1.0 0.8 11.6 0.9 0.7 11.9 0.5 0.7 9.1 0.8 0.5 12.6 0.5 1.1 10.9 0.5 0.3
-Tsucs 2 3.3 0.1 0.6 11.6 0.3 0.4 10.3 0.5 0.4 10.6 0.3 0.3 11.6 0.6 0.4 13.8 0.7 1.0 10.6 0.7 1.0 11.2 0.8 0.9 8.2 0.7 1.0 10.5 0.8 0.9 11.1 0.7 0.8 8.4 0.4 1.2 10.3 0.5 0.4
-Vallise El Nino 1.1 0.1 0.5 8.4 0.3 0.7 11.2 0.5 1.0 12.5 0.3 1.8 12.7 0.5 1.9 13.5 0.6 3.9 12.3 0.7 1.4 13.4 0.5 1.8 12.3 0.7 2.8 12.6 0.5 2.3 13.9 0.8 3.0 13.7 0.3 3.8 13.2 0.3 2.3
-Wang Sun 4.9 nan 0.1 13.2 0.5 0.5 12.2 0.5 0.5 11.8 0.5 0.6 11.1 0.4 0.8 14.6 0.7 1.2 13.9 0.7 1.5 10.9 0.7 1.7 10.8 0.6 1.8 10.9 0.7 1.6 13.3 0.7 1.6 11.1 0.4 1.0 12.3 0.5 0.9
-Winsmi Reduced 8.3 0.5 0.6 11.4 0.4 0.1 10.7 0.5 0.1 13.5 0.6 0.4 13.7 0.6 0.2 14.0 0.4 0.8 11.9 0.9 3.5 12.2 0.8 3.8 10.5 0.6 3.3 9.4 0.3 1.9 12.0 0.8 3.3 14.2 0.2 4.7 13.6 0.2 1.2
-Yu Wang 4.9 0.3 0.6 9.8 0.7 0.6 8.4 0.3 0.3 12.4 0.3 0.3 12.4 0.5 0.7 13.0 0.6 0.8 12.4 0.9 3.7 12.5 0.7 4.4 12.8 0.8 4.4 11.9 0.9 3.3 12.5 0.8 3.4 13.5 0.4 3.4 13.2 0.5 3.0
-Yu Wang 2 2.7 0.1 0.3 9.6 0.5 0.2 10.0 0.6 0.3 12.4 0.3 0.5 12.2 0.3 0.4 14.1 0.7 0.5 9.9 0.2 0.5 9.5 0.7 1.5 9.9 0.6 2.3 10.4 0.4 1.3 10.6 0.7 2.9 12.6 0.2 2.6 10.6 0.2 1.7
-Zhou Chen 4.6 nan 0.4 11.1 0.3 1.0 5.9 0.4 0.5 11.6 0.3 1.8 14.2 0.4 1.5 10.3 0.3 1.8 12.8 0.4 1.5 12.7 0.5 2.1 13.5 0.4 1.9 12.9 0.5 2.6 13.1 0.5 3.7 13.2 0.2 5.1 14.9 0.4 1.3
-Table 10: Performance comparison for zero-shot inference across all DS in the dataset from [29] using $T_{C}=512$. Best value in red, second best in blue.
-System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM TimesFM
+Atmospheric Regime 10.4 nan 0.0 9.2 0.4 0.0 10.5 0.4 0.0 12.0 0.5 0.0 13.3 0.5 0.0 11.6 0.9 0.1 12.3 0.8 0.2 12.3 0.9 0.1 10.9 0.5 0.1 12.0 0.4 0.1 11.2 0.8 0.7 13.9 0.4 1.6 11.8 0.4 0.1
+BeerRNN 5.2 0.7 0.6 12.1 0.5 0.5 9.3 0.4 0.7 13.3 0.4 0.7 10.9 0.3 0.6 12.4 0.9 1.0 12.7 0.8 1.7 11.9 0.9 1.5 13.5 0.3 0.3 14.0 0.3 0.4 11.4 0.8 1.0 14.2 0.4 1.5 14.0 0.3 0.5
+Bouali 5.9 0.1 0.2 7.0 0.3 0.4 12.1 0.2 0.5 11.6 0.1 0.5 13.1 0.2 0.4 11.2 0.7 2.3 12.0 0.6 3.1 12.3 0.5 2.0 6.7 0.1 0.6 13.0 0.2 0.6 12.9 0.7 4.4 14.0 0.3 5.6 9.7 0.4 1.1
+Burke Shaw 3.8 0.2 0.3 8.0 0.1 0.5 5.6 0.1 0.4 5.1 0.2 0.2 11.1 0.2 0.5 11.0 0.6 1.3 11.0 0.7 0.9 11.7 0.8 0.9 12.4 0.3 0.9 12.3 0.2 0.4 10.2 0.7 1.2 12.9 0.4 2.1 12.4 0.3 0.8
+Ca Two Plus Quasip. 2.8 0.1 0.3 9.6 0.5 1.6 9.5 0.3 0.5 12.6 0.2 1.2 12.7 0.4 1.4 9.4 0.7 2.6 9.4 0.6 1.3 9.6 0.7 1.4 8.4 0.2 1.1 12.8 0.2 0.7 11.1 0.6 4.0 13.8 0.2 4.6 13.4 0.3 1.9
+Chen Lee 3.6 0.3 0.2 4.2 0.2 0.0 3.6 0.2 0.1 3.9 0.2 0.1 5.3 0.3 0.1 4.5 0.8 0.4 8.8 0.7 0.9 7.7 0.7 0.6 3.9 0.2 0.2 6.2 0.2 0.3 11.3 0.8 0.8 8.9 0.4 2.0 4.6 0.2 0.3
+Chua 2.9 0.1 0.8 12.9 0.2 0.6 12.4 0.4 0.8 13.3 0.2 0.2 13.5 0.4 1.0 11.9 0.6 2.4 11.7 0.4 1.3 11.7 0.4 1.8 12.9 0.2 1.9 11.6 0.3 0.6 12.1 0.4 0.8 13.4 0.3 4.0 12.7 0.2 1.0
+Coullet 0.9 0.0 0.2 0.6 0.0 0.3 6.3 0.0 0.2 3.9 0.1 0.2 13.7 0.3 0.2 11.4 0.7 1.1 12.8 0.7 1.5 11.8 0.6 1.1 9.8 0.2 0.3 7.3 0.2 0.4 13.3 0.6 1.5 15.0 0.2 3.5 9.3 0.2 0.5
+Dadras 2.2 0.2 1.3 10.0 0.5 4.2 9.2 0.5 1.9 11.8 0.4 2.3 11.4 0.3 3.5 10.5 0.8 5.4 9.4 0.7 5.0 11.0 0.6 6.8 11.1 0.3 3.3 12.4 0.3 2.8 13.1 0.8 5.0 12.2 0.4 6.0 13.1 0.6 4.5
+Dequan Li 8.5 0.1 0.8 12.6 0.5 0.2 13.2 0.4 0.3 13.1 0.3 0.3 12.9 0.4 1.4 12.4 0.4 0.8 12.7 0.4 0.8 12.2 0.4 0.6 10.0 0.2 0.6 12.8 0.4 1.0 12.3 0.4 3.3 13.0 0.4 1.7 13.0 0.3 1.3
+Finance 3.1 0.1 0.3 8.5 0.4 0.7 9.4 0.6 0.5 13.4 0.3 0.6 12.2 0.4 0.9 12.0 0.7 2.8 11.7 0.6 4.0 12.4 0.7 3.7 12.8 0.3 0.8 13.2 0.2 0.3 12.6 0.8 4.9 14.0 0.3 4.2 13.2 0.5 2.7
+Genesio Tesi 9.8 0.2 0.4 8.9 0.1 0.2 8.7 0.1 0.1 7.7 0.1 0.1 11.8 0.4 0.4 10.7 0.3 0.8 11.4 0.9 1.4 10.6 0.6 2.0 11.5 0.4 0.6 10.1 0.2 0.2 10.4 0.8 1.3 12.3 0.2 1.7 12.2 0.2 1.1
+Guckenheimer Holmes 7.4 0.0 0.2 10.6 0.4 0.3 11.7 0.2 0.3 11.3 0.4 0.4 13.2 0.7 0.6 10.7 0.6 0.7 10.6 0.6 1.0 9.2 0.6 0.4 6.8 0.1 0.3 9.1 0.1 0.7 10.4 0.5 2.0 11.6 0.3 3.2 11.0 0.2 0.7
+Hadley 8.6 0.3 0.3 10.8 0.5 0.2 9.3 0.6 0.4 11.9 0.4 0.4 12.5 0.4 0.6 10.9 0.7 0.6 10.7 0.7 2.0 9.5 0.6 0.7 12.1 0.2 0.4 11.9 0.2 0.5 11.7 0.7 1.7 12.5 0.3 5.4 11.7 0.3 0.4
+Halvorsen 1.8 0.1 0.2 6.0 0.1 0.3 8.4 0.1 0.5 5.9 0.1 0.3 12.6 0.2 0.5 11.8 0.8 1.6 9.5 0.7 2.1 8.9 0.8 1.8 13.0 0.2 0.3 12.9 0.2 0.3 12.2 0.6 3.4 13.2 0.3 2.8 12.8 0.2 0.8
+Hindmarsh Rose 6.1 0.5 1.1 9.7 0.3 0.3 6.7 0.5 0.7 8.4 0.1 0.3 11.5 0.4 0.9 12.7 0.7 1.5 12.5 0.7 2.2 13.9 0.6 1.2 9.4 0.3 0.4 8.2 0.1 1.8 13.9 0.6 2.5 14.0 0.2 2.7 14.2 0.6 1.5
+Isothermal Chemical 2.4 0.2 0.1 3.7 0.1 0.0 7.9 0.3 0.0 9.7 0.3 0.0 9.0 0.3 0.1 7.9 0.8 0.5 5.9 0.8 0.4 8.4 0.8 0.6 3.7 0.3 0.2 7.0 0.2 0.3 12.8 0.8 0.7 15.6 0.2 6.1 4.0 0.4 0.3
+Itik Banks tumor 8.7 0.2 7.0 10.0 0.6 1.7 8.4 0.5 2.4 10.5 0.4 1.5 12.8 0.3 0.8 9.2 0.7 3.6 9.5 0.7 4.6 9.7 0.8 5.1 9.1 0.3 1.4 12.2 0.3 1.4 13.3 0.9 6.4 13.2 0.3 8.2 12.7 0.3 3.8
+Jerk Circuit 3.2 0.4 3.3 10.1 0.4 1.3 9.7 0.4 1.0 7.5 0.3 0.5 11.1 0.4 0.5 10.8 0.8 1.5 10.9 0.5 2.1 10.5 0.9 1.7 11.7 0.5 1.2 12.4 0.3 0.9 11.1 0.9 1.6 13.9 0.4 2.8 12.6 0.3 1.4
+Laser 8.8 0.4 0.2 11.1 0.4 0.1 10.8 0.2 0.1 12.8 0.2 0.3 13.8 0.5 0.6 11.8 0.7 0.4 11.5 0.7 0.5 11.0 0.7 0.8 10.9 0.2 0.4 12.7 0.3 0.3 13.3 0.6 1.2 13.8 0.2 2.2 12.8 0.4 0.4
+Lorenz 1.2 0.1 0.2 6.6 0.2 0.4 6.0 0.2 0.3 10.0 0.2 0.3 12.0 0.4 0.3 10.9 0.7 0.5 11.3 0.5 0.5 12.7 0.6 1.1 12.8 0.3 0.2 11.9 0.2 0.5 11.4 0.7 1.2 13.0 0.3 2.2 13.1 0.5 0.8
+Lorenz 84 6.9 0.4 1.5 9.5 0.4 2.2 11.0 0.4 1.9 11.5 0.4 1.3 12.2 0.4 1.7 12.9 0.8 2.4 12.9 0.8 2.7 11.5 0.9 2.5 13.4 0.7 1.6 13.4 0.4 0.8 13.4 0.8 2.5 13.4 0.5 2.4 13.4 0.5 2.0
+Lu Chen 7.7 0.8 0.2 6.3 0.4 0.2 5.3 0.3 0.2 8.0 0.1 0.1 7.1 0.1 0.1 11.6 0.9 0.7 11.9 1.0 0.8 11.3 1.0 0.9 6.4 0.3 0.3 12.2 0.3 0.3 9.7 0.8 0.9 13.7 0.4 1.5 8.6 0.5 0.4
+Moore Spiegel 4.3 0.2 0.3 10.4 0.6 0.5 10.0 0.5 0.1 12.5 0.4 0.2 9.9 0.5 0.2 10.5 0.9 1.5 11.4 0.9 1.6 10.1 0.9 1.6 12.6 0.5 1.9 13.2 0.4 0.5 9.4 0.8 2.8 12.5 0.3 2.4 12.4 0.3 2.8
+Multi Chua 8.1 0.3 0.8 9.3 0.5 0.3 9.3 0.2 0.5 7.1 0.4 0.4 13.3 0.2 1.1 11.5 0.6 5.8 12.2 0.4 11.5 9.5 0.5 11.3 9.7 0.3 1.3 11.0 0.3 0.5 10.6 0.4 4.8 14.0 0.3 7.6 11.4 0.2 2.0
+Nose Hoover 9.5 0.2 0.1 9.9 0.5 0.1 9.6 0.6 0.1 11.5 0.3 0.5 12.5 0.7 0.8 10.6 0.9 0.6 10.6 0.9 0.6 10.1 0.7 0.8 11.2 0.3 0.3 12.6 0.3 0.7 10.2 0.7 1.1 12.9 0.4 1.6 13.3 0.3 0.9
+Pehlivan Wei 9.2 0.0 0.2 8.5 0.4 0.5 10.1 0.4 0.5 11.3 0.3 0.3 10.8 0.4 0.5 9.2 0.7 1.2 10.1 0.8 1.7 9.2 0.7 1.0 8.6 0.2 0.5 8.6 0.2 0.4 9.3 0.7 1.4 12.2 0.6 3.0 8.2 0.3 1.2
+Sakarya 2.0 0.2 0.7 5.6 0.5 0.3 5.0 0.3 0.3 7.5 0.2 0.3 11.1 0.4 0.4 10.2 0.9 1.0 10.0 0.6 1.6 7.7 0.9 2.2 13.4 0.5 1.2 12.7 0.3 0.6 13.1 0.8 1.5 12.6 0.4 1.2 11.9 0.3 2.3
+San Um Srisuchinwong 3.6 0.2 0.3 12.9 0.4 1.2 12.7 0.5 2.1 11.4 0.5 1.4 13.3 0.5 2.5 12.7 0.7 1.5 10.4 0.7 2.1 13.4 0.6 1.8 12.5 0.3 1.4 12.2 0.3 1.6 10.7 0.7 2.5 13.1 0.4 3.7 13.3 0.7 1.8
+Sprott A 2.7 0.1 0.3 8.0 0.2 0.3 8.9 0.3 0.3 6.6 0.2 0.3 11.8 0.3 0.3 10.5 0.8 0.9 8.1 0.5 1.6 10.0 0.7 1.5 9.3 0.3 0.3 7.0 0.1 0.4 9.9 0.7 2.5 12.0 0.3 5.7 9.8 0.2 1.0
+Sprott B 2.9 0.2 0.4 4.3 0.2 1.1 5.6 0.4 0.9 11.7 0.5 0.7 12.7 0.4 0.8 11.8 0.8 1.3 11.5 0.6 0.8 12.3 0.8 1.2 13.1 0.5 0.7 12.8 0.2 0.5 12.5 0.8 3.0 12.6 0.4 2.1 13.5 0.6 0.9
+Sprott C 2.4 0.1 0.4 11.6 0.4 2.3 12.4 0.4 1.3 11.9 0.4 2.3 13.3 0.3 1.1 10.0 0.5 7.7 9.2 0.5 6.8 10.7 0.5 8.0 13.7 0.3 2.3 13.1 0.3 1.1 12.8 0.6 5.2 13.3 0.2 6.2 13.4 0.2 4.5
+Sprott D 3.0 0.1 0.4 9.8 0.6 0.3 10.3 0.4 0.3 12.9 0.3 0.4 13.4 0.5 0.7 12.2 0.7 0.8 10.9 0.6 0.6 9.9 0.5 1.2 9.1 0.2 0.4 9.8 0.2 0.3 12.9 0.7 1.3 14.0 0.2 4.0 7.9 0.5 0.9
+Sprott F 5.0 0.3 0.7 9.6 0.5 0.2 9.2 0.5 0.6 12.3 0.3 0.4 12.6 0.4 1.0 9.9 0.7 1.3 10.4 0.7 1.5 10.4 0.6 0.9 12.3 0.4 0.5 12.7 0.5 0.4 11.4 0.8 1.6 13.3 0.3 2.8 13.0 0.6 1.7
+Sprott G 3.8 0.1 0.3 7.0 0.6 0.2 7.7 0.5 0.2 10.0 0.3 0.4 7.2 0.1 0.6 10.9 0.9 1.9 11.4 0.6 1.9 9.2 0.8 1.8 8.0 0.1 0.7 9.2 0.3 0.4 14.1 0.8 1.4 13.1 0.2 1.9 9.5 0.4 1.4
+Sprott I 2.7 0.1 0.5 9.7 0.5 0.7 11.3 0.2 0.6 9.0 0.1 0.5 11.9 0.3 0.6 10.2 0.9 3.5 9.8 0.9 4.9 8.4 0.6 4.6 8.6 0.2 0.4 7.5 0.2 0.2 12.2 0.7 2.0 14.4 0.2 8.8 8.3 0.4 1.0
+Sprott L 3.6 0.1 0.6 8.0 0.5 0.3 6.8 0.1 0.3 8.2 0.2 0.1 11.6 0.3 0.4 10.5 0.9 1.8 10.4 0.8 2.8 11.4 0.9 1.8 8.8 0.2 0.1 11.9 0.1 0.3 10.8 0.8 1.3 14.2 0.3 2.9 8.1 0.4 0.2
+Sprott M 3.6 0.1 0.2 9.8 0.4 0.2 10.5 0.4 0.5 12.4 0.1 0.3 12.9 0.2 0.3 9.8 0.8 3.7 10.3 0.6 3.7 9.9 0.8 3.3 12.5 0.2 0.8 12.8 0.3 0.3 11.6 0.7 2.9 14.3 0.2 6.5 13.4 0.4 1.6
+Sprott O 2.0 0.1 0.4 3.6 0.1 0.4 4.6 0.0 0.4 9.7 0.1 0.5 12.4 0.4 0.6 10.8 1.0 1.6 12.6 0.9 2.6 10.4 0.8 1.8 9.3 0.2 0.7 7.2 0.1 0.2 10.7 0.8 1.6 14.0 0.4 3.7 9.5 0.5 0.7
+Sprott P 3.8 0.2 1.4 10.8 0.3 3.9 11.9 0.3 2.6 12.7 0.3 2.6 12.7 0.3 2.1 7.3 0.5 5.7 10.6 0.6 9.6 8.3 0.6 8.7 10.5 0.3 1.1 10.9 0.2 0.9 11.1 0.7 6.3 12.2 0.3 10.2 11.8 0.3 4.4
+Sprott Q 4.5 0.1 0.2 10.1 0.2 0.3 7.6 0.1 0.4 7.3 0.1 0.5 11.6 0.2 0.4 11.8 0.8 3.6 11.6 0.7 1.9 11.9 0.6 1.0 9.8 0.1 0.5 11.5 0.3 0.4 10.9 0.7 1.1 13.1 0.4 7.5 8.8 0.4 0.4
+Sprott R 4.3 0.3 0.2 9.3 0.8 0.2 6.7 0.5 0.3 11.0 0.5 0.5 12.6 0.5 0.6 10.6 0.8 0.5 10.4 0.5 1.2 9.6 0.7 2.1 8.1 0.4 1.2 8.0 0.3 0.3 11.8 0.8 1.3 13.8 0.3 2.5 11.3 0.3 0.9
+Sprott S 4.2 0.1 0.3 10.1 0.9 2.2 9.9 0.6 1.1 12.3 0.2 1.2 12.6 0.4 0.6 10.6 0.7 1.2 9.7 0.8 6.1 11.0 0.6 3.1 11.9 0.3 2.3 12.8 0.1 0.5 11.8 0.8 4.1 11.6 0.3 10.2 12.8 0.2 3.7
+Sprott Torus 2.2 0.2 0.2 5.4 0.7 0.4 10.4 0.7 0.3 14.3 0.6 0.3 14.5 0.6 0.6 6.2 0.8 4.2 8.3 0.6 8.3 5.3 0.8 6.3 13.5 0.3 0.3 6.3 0.3 0.4 13.6 0.8 2.4 15.5 0.3 4.8 14.3 0.5 1.0
+Thomas 1.6 0.3 4.8 10.6 0.5 6.0 10.1 0.8 6.0 9.9 0.3 5.7 13.9 0.5 4.7 12.6 0.8 3.4 12.6 0.8 1.9 12.8 0.8 2.8 13.2 0.3 4.8 13.7 0.3 3.5 10.7 0.5 6.1 13.3 0.2 3.1 13.8 0.6 3.3
+Thomas Labyrinth 4.1 0.1 2.4 12.5 0.2 2.6 12.5 0.3 2.8 12.4 0.4 3.0 12.5 0.4 2.8 11.4 0.5 3.2 12.5 0.5 2.6 11.9 0.5 3.0 12.1 0.3 3.4 12.5 0.2 3.0 12.3 0.5 3.6 12.4 0.2 2.0 12.6 0.5 2.5
+Torus 11.9 0.9 1.1 9.8 0.3 0.1 9.4 0.2 0.1 10.9 0.3 0.1 12.3 0.3 0.2 12.1 0.9 0.8 10.5 1.0 0.8 11.6 0.9 0.7 11.5 0.3 0.1 12.6 0.2 0.3 9.1 0.8 0.5 12.6 0.5 1.1 10.9 0.5 0.3
+Tsucs 2 3.3 0.1 0.6 11.6 0.3 0.4 10.3 0.5 0.4 10.6 0.3 0.3 11.6 0.6 0.4 10.6 0.7 1.0 11.2 0.8 0.9 8.2 0.7 1.0 4.8 0.2 0.3 7.8 0.2 0.7 11.1 0.7 0.8 8.4 0.4 1.2 10.3 0.5 0.4
+Vallise El Nino 1.1 0.1 0.5 8.4 0.3 0.7 11.2 0.5 1.0 12.5 0.3 1.8 12.7 0.5 1.9 12.3 0.7 1.4 13.4 0.5 1.8 12.3 0.7 2.8 13.2 0.3 1.3 13.1 0.2 0.4 13.9 0.8 3.0 13.7 0.3 3.8 13.2 0.3 2.3
+Wang Sun 4.9 nan 0.1 13.2 0.5 0.5 12.2 0.5 0.5 11.8 0.5 0.6 11.1 0.4 0.8 13.9 0.7 1.5 10.9 0.7 1.7 10.8 0.6 1.8 12.7 0.4 0.4 12.1 0.5 0.5 13.3 0.7 1.6 11.1 0.4 1.0 12.3 0.5 0.9
+Winsmi Reduced 8.3 0.5 0.6 11.4 0.4 0.1 10.7 0.5 0.1 13.5 0.6 0.4 13.7 0.6 0.2 11.9 0.9 3.5 12.2 0.8 3.8 10.5 0.6 3.3 10.7 0.2 0.6 12.3 0.2 0.7 12.0 0.8 3.3 14.2 0.2 4.7 13.6 0.2 1.2
+Yu Wang 4.9 0.3 0.6 9.8 0.7 0.6 8.4 0.3 0.3 12.4 0.3 0.3 12.4 0.5 0.7 12.4 0.9 3.7 12.5 0.7 4.4 12.8 0.8 4.4 13.1 0.6 1.6 13.1 0.2 0.7 12.5 0.8 3.4 13.5 0.4 3.4 13.2 0.5 3.0
+Yu Wang 2 2.7 0.1 0.3 9.6 0.5 0.2 10.0 0.6 0.3 12.4 0.3 0.5 12.2 0.3 0.4 9.9 0.2 0.5 9.5 0.7 1.5 9.9 0.6 2.3 9.0 0.2 0.9 11.5 0.3 0.4 10.6 0.7 2.9 12.6 0.2 2.6 10.6 0.2 1.7
+Zhou Chen 4.6 nan 0.4 11.1 0.3 1.0 5.9 0.4 0.5 11.6 0.3 1.8 14.2 0.4 1.5 12.8 0.4 1.5 12.7 0.5 2.1 13.5 0.4 1.9 12.8 0.3 1.4 9.1 0.3 0.3 13.1 0.5 3.7 13.2 0.2 5.1 14.9 0.4 1.3
+Table 10: Performance comparison for zero-shot inference across all DS in the dataset from Gilpin [2022] using $T_{C}=512$. Best value in red, second best in blue.
+System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-2 Panda Mamba4Cast TTM TimesFM
 $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
-Atmospheric Regime 8.4 0.6 2.3 9.7 0.5 0.9 11.1 0.4 0.5 10.5 0.6 0.7 10.1 0.5 1.0 12.0 0.4 0.5 11.9 0.9 2.2 12.0 0.9 1.9 13.5 0.9 1.8 12.7 0.8 2.7 11.3 0.9 1.3 14.2 0.6 3.1 13.9 0.4 1.2
-BeerRNN 11.7 0.8 1.2 8.0 0.5 0.2 3.3 0.2 0.2 9.5 0.2 0.2 7.3 0.4 0.2 11.9 0.4 0.3 12.3 0.9 1.2 12.5 0.9 1.2 12.1 0.9 1.2 12.6 0.8 1.1 13.0 0.9 1.4 14.1 0.6 1.3 14.0 0.4 0.8
-Bouali 4.1 0.1 0.5 8.3 0.5 0.1 7.3 0.3 0.2 8.6 0.1 0.2 11.1 0.3 0.5 13.2 0.5 0.3 10.6 0.8 1.0 9.6 0.7 1.1 12.2 0.7 2.0 12.3 0.8 1.2 12.9 0.8 2.3 13.4 0.4 2.3 12.7 0.3 0.7
-Burke Shaw 4.1 0.2 0.5 10.8 0.5 1.2 9.0 0.5 1.5 13.0 0.4 1.1 13.2 0.4 1.5 13.3 0.6 0.9 11.5 0.8 1.3 12.1 0.6 2.4 12.2 0.7 2.2 10.5 0.7 4.8 12.2 0.9 3.7 13.7 0.5 3.6 13.9 0.4 1.5
-Ca Two Plus Quasip. 2.3 0.1 0.7 7.2 0.1 0.2 8.3 0.1 0.3 9.2 0.4 0.3 13.1 0.4 0.2 13.2 0.4 0.6 11.3 0.9 1.0 11.6 0.8 1.1 12.8 0.6 1.2 8.9 0.6 1.6 11.1 0.8 3.0 13.8 0.7 2.6 13.4 0.3 1.9
-Chen Lee 3.9 0.4 0.3 4.9 0.3 0.3 4.0 0.5 0.5 4.8 0.3 0.1 4.6 0.3 0.3 8.9 0.7 0.4 7.2 0.9 1.6 7.9 0.7 1.4 8.0 0.7 2.0 4.8 0.5 1.7 11.0 0.9 0.9 6.4 0.5 3.4 7.3 0.4 1.8
-Chua 5.9 0.1 0.3 7.4 0.2 0.6 5.9 0.2 0.7 7.6 0.1 0.7 11.7 0.2 0.4 13.1 0.3 1.4 11.8 0.6 1.9 12.2 0.5 2.2 11.7 0.6 1.7 12.0 0.6 2.7 11.3 0.5 3.2 13.4 0.4 2.7 10.4 0.2 0.7
-Coullet 1.2 0.0 0.2 4.9 0.0 0.3 0.7 0.0 0.1 2.8 0.0 0.1 7.5 0.2 0.1 14.6 0.4 0.6 13.0 0.8 2.8 11.2 0.8 3.8 11.8 0.8 3.4 11.6 0.7 4.9 12.3 0.8 2.4 14.9 0.5 3.3 10.7 0.2 0.5
-Dadras 4.1 0.2 1.1 7.7 0.6 0.5 4.7 0.5 0.5 13.1 0.5 0.3 12.1 0.4 0.2 10.2 0.7 0.5 12.7 0.8 1.7 12.9 0.7 2.0 13.0 0.8 2.2 8.8 0.8 3.8 11.9 0.8 1.7 12.8 0.4 2.9 11.8 0.4 1.2
-Dequan Li 12.7 0.5 2.1 11.1 0.3 1.1 11.0 0.4 0.7 12.1 0.3 0.9 12.1 0.2 1.9 13.0 0.3 3.1 11.7 0.4 3.1 11.3 0.4 2.8 9.6 0.4 2.4 11.0 0.4 1.8 10.7 0.4 3.1 13.5 0.4 4.5 12.8 0.3 2.0
-Finance 3.2 0.1 0.6 8.5 0.6 1.6 8.6 0.4 0.9 9.1 0.4 0.4 12.5 0.3 0.4 13.8 0.4 1.5 11.6 0.9 1.8 10.7 0.8 1.7 13.4 0.8 2.0 11.8 0.8 1.7 11.1 0.9 2.3 13.7 0.5 4.2 13.6 0.4 2.6
-Genesio Tesi 8.2 0.2 0.9 7.9 0.4 1.0 8.5 0.2 1.1 7.7 0.2 1.0 9.4 0.2 1.0 13.0 0.3 2.0 10.4 0.7 1.4 10.6 0.7 1.1 10.3 0.8 2.2 9.6 0.8 2.1 11.0 0.9 1.4 12.6 0.6 5.4 10.6 0.4 1.0
-Guckenheimer Holmes 9.6 0.1 0.5 11.8 0.1 0.8 12.0 0.1 0.4 12.6 0.3 0.5 12.5 0.4 1.1 12.1 0.4 0.6 9.6 0.7 0.6 9.6 0.7 2.5 9.4 0.6 1.0 11.8 0.6 1.4 10.2 0.7 1.0 12.7 0.5 3.5 11.5 0.3 0.9
-Hadley 11.3 0.4 1.3 10.7 0.5 2.3 9.7 0.5 1.8 12.5 0.3 1.5 12.6 0.6 1.8 12.6 0.4 2.8 11.4 0.8 5.1 11.9 0.7 4.7 11.9 0.6 3.9 11.5 0.8 5.7 11.4 0.8 6.3 12.5 0.5 7.1 12.4 0.5 3.8
-Halvorsen 4.3 0.3 0.6 9.7 0.6 0.4 8.2 0.2 0.7 7.5 0.2 0.7 13.1 0.2 0.6 13.3 0.4 1.9 10.0 0.8 5.0 10.3 0.7 4.3 9.7 0.8 4.6 10.9 0.7 4.1 11.3 0.9 3.4 13.4 0.5 4.2 13.3 0.3 1.6
-Hindmarsh Rose 6.5 0.4 0.7 9.8 0.2 0.2 4.9 0.1 0.1 9.6 0.2 0.1 12.7 0.6 0.3 13.7 0.6 0.9 12.9 0.7 1.6 14.3 0.8 2.0 12.9 0.8 1.7 14.0 0.8 1.9 13.2 0.8 2.5 14.6 0.4 1.4 14.3 0.5 1.1
-Isothermal Chemical 2.0 0.2 0.2 2.5 0.1 0.6 5.4 0.1 0.3 8.3 0.2 0.2 7.0 0.4 0.2 11.6 0.5 0.5 11.7 0.9 1.1 11.7 0.9 1.5 10.8 0.8 1.6 12.5 0.9 2.0 10.1 0.9 0.7 15.6 0.6 3.5 7.5 0.3 1.3
-Itik Banks tumor 5.7 0.2 0.5 8.2 0.5 0.5 10.8 0.6 0.4 9.8 0.5 0.3 11.8 0.6 0.4 13.3 0.6 0.3 10.8 0.6 3.4 11.2 0.8 3.8 11.9 0.9 2.9 12.0 0.8 5.0 12.0 0.9 2.9 12.9 0.7 4.1 13.4 0.3 1.1
-Jerk Circuit 4.3 0.3 0.8 8.8 0.6 0.4 9.1 0.4 0.3 10.6 0.3 0.2 11.7 0.3 0.2 9.8 0.6 0.3 12.5 0.9 1.0 13.5 0.9 0.8 12.9 0.9 0.8 13.1 1.0 0.9 10.3 0.8 0.6 14.1 0.5 1.9 13.7 0.4 1.1
-Laser 4.5 0.2 0.2 7.5 0.5 3.0 8.2 0.3 1.2 11.5 0.2 0.5 12.8 0.2 0.5 13.9 0.6 0.8 12.6 0.7 2.0 12.8 0.5 2.8 13.0 0.7 1.7 12.8 0.7 5.1 11.8 0.8 2.5 13.3 0.5 2.6 13.8 0.4 1.0
-Lorenz 3.4 0.2 0.5 9.8 0.6 1.5 8.5 0.3 1.6 13.0 0.4 0.7 12.6 0.7 2.0 13.3 0.6 0.8 12.8 0.8 2.6 13.0 0.7 3.4 13.2 0.6 3.0 13.0 0.6 2.3 11.8 0.8 2.7 13.4 0.5 3.1 13.1 0.3 2.1
-Lorenz 84 7.9 0.2 0.5 10.7 0.5 1.5 11.0 0.4 1.3 12.6 0.4 1.0 11.0 0.4 0.4 13.2 0.6 0.4 11.8 0.9 1.3 12.2 0.9 1.4 11.5 0.9 1.3 12.3 0.9 1.3 11.2 0.9 1.2 13.4 0.6 1.2 13.4 0.5 1.4
-Lu Chen 9.5 0.8 0.7 5.3 0.2 0.5 9.0 0.3 0.7 8.9 0.2 0.2 9.7 0.3 0.5 8.3 0.3 0.6 12.2 1.0 1.7 13.8 0.9 1.8 12.2 0.9 1.6 13.1 1.0 2.1 10.1 0.8 1.1 14.2 0.5 1.9 7.4 0.5 0.9
-Moore Spiegel 9.2 0.6 0.4 9.1 0.6 0.4 10.9 0.7 0.4 12.5 0.5 0.3 11.5 0.6 0.3 11.3 0.4 0.3 11.4 0.9 2.0 11.7 0.8 2.7 12.7 0.9 2.0 12.0 0.8 2.4 10.5 0.9 2.2 13.1 0.7 1.5 12.6 0.5 2.1
-Multi Chua 12.5 0.5 0.4 12.0 0.6 0.8 11.9 0.6 0.4 13.4 0.4 1.0 13.0 0.4 0.8 13.9 0.4 1.3 12.5 0.5 1.5 12.6 0.5 1.1 13.6 0.5 1.1 13.0 0.5 1.2 10.7 0.5 0.7 13.5 0.5 2.1 12.7 0.3 1.9
-Nose Hoover 11.4 0.5 0.9 10.2 0.6 0.5 10.8 0.7 0.5 10.8 0.6 0.5 12.9 0.8 0.6 13.3 0.5 1.0 11.2 0.9 2.0 11.3 0.9 1.4 12.7 0.7 2.0 11.0 0.9 2.2 10.5 0.9 2.9 13.3 0.5 3.3 12.8 0.4 0.5
-Pehlivan Wei 4.8 0.2 1.9 4.6 0.1 0.7 6.8 0.4 0.9 9.0 0.3 1.0 10.7 0.3 1.3 12.4 0.5 2.1 11.1 0.9 5.6 12.0 0.6 6.8 12.6 0.8 6.5 12.3 0.8 6.3 10.0 0.9 4.1 12.7 0.4 7.7 9.1 0.3 3.3
-Sakarya 3.1 0.3 0.4 2.6 0.2 0.2 5.0 0.3 0.3 8.2 0.3 0.3 10.2 0.7 0.6 12.9 0.8 0.9 10.7 0.9 0.5 12.7 0.9 2.5 11.8 0.8 0.9 12.8 0.8 1.7 12.4 0.7 0.8 14.0 0.5 2.5 13.8 0.5 1.5
-San Um Srisuchinwong 5.5 0.1 0.6 5.5 0.2 1.2 10.5 0.8 1.3 12.6 0.4 1.2 13.0 0.5 0.8 13.1 0.6 0.6 11.8 0.8 1.1 13.0 0.7 0.5 12.0 0.6 1.2 11.9 0.7 0.9 10.7 0.5 1.8 12.6 0.5 0.9 12.6 0.4 1.4
-Sprott A 4.9 0.1 1.0 9.3 0.5 0.8 8.5 0.4 0.8 7.5 0.3 0.4 11.7 0.3 0.9 12.6 0.5 2.2 10.7 0.9 4.1 11.8 0.7 4.5 11.5 0.8 4.4 11.1 0.7 5.0 9.8 0.9 2.1 12.4 0.6 6.0 11.2 0.2 1.1
-Sprott B 2.6 0.2 0.9 9.2 0.5 1.7 11.0 0.6 1.8 13.2 0.5 2.5 13.4 0.6 1.5 13.4 0.7 1.9 12.6 0.8 2.7 13.2 0.8 2.7 12.5 0.8 2.2 12.6 0.7 1.9 13.9 0.8 3.9 12.9 0.6 2.9 13.7 0.3 2.5
-Sprott C 2.9 0.2 0.4 8.5 0.6 1.1 12.0 0.3 0.5 13.3 0.4 1.1 12.6 0.4 0.8 11.8 0.5 2.0 12.4 0.7 2.0 13.0 0.6 2.2 11.9 0.6 2.5 11.8 0.6 1.7 13.3 0.7 2.6 13.5 0.5 2.9 13.5 0.4 1.9
-Sprott D 3.8 0.1 0.6 7.6 0.2 0.2 5.5 0.1 0.2 5.5 0.2 0.1 12.9 0.4 0.7 13.1 0.5 0.5 12.1 0.7 0.8 12.0 0.8 4.6 12.1 0.7 3.1 12.0 0.6 3.7 12.9 0.7 1.9 13.8 0.5 2.3 13.2 0.5 0.6
-Sprott F 4.9 0.3 0.5 10.8 0.4 0.3 8.9 0.5 0.5 12.6 0.3 0.3 13.0 0.5 0.2 13.1 0.4 0.3 12.3 0.9 1.2 12.2 0.6 1.3 13.2 0.8 1.2 12.4 0.7 1.3 11.8 0.9 1.3 13.4 0.6 1.8 13.2 0.5 1.2
-Sprott G 2.4 0.0 0.3 4.2 0.1 0.1 6.2 0.2 0.1 7.1 0.2 0.2 7.7 0.3 0.5 12.9 0.3 0.6 12.3 0.8 3.6 12.4 0.7 4.0 12.9 0.8 3.4 11.2 0.9 4.5 11.0 0.9 2.6 14.0 0.5 2.0 10.6 0.3 1.2
-Sprott I 3.0 0.2 0.3 8.6 0.5 0.1 10.8 0.1 0.2 7.4 0.3 0.3 10.1 0.3 0.3 13.1 0.4 0.2 12.1 0.9 1.2 11.9 0.8 2.5 9.4 0.9 1.0 13.4 0.9 2.4 10.9 0.9 0.9 14.4 0.6 3.2 9.1 0.4 0.3
-Sprott L 2.7 0.2 0.2 9.2 0.3 0.2 8.2 0.1 0.2 10.2 0.2 0.1 11.3 0.2 0.1 13.6 0.4 0.2 12.3 0.9 2.6 12.3 0.9 2.7 13.3 0.9 1.8 13.1 0.9 4.8 12.2 0.8 0.6 14.2 0.5 2.5 11.7 0.3 0.2
-Sprott M 3.4 0.2 0.4 10.9 0.5 0.1 12.4 0.4 0.2 12.8 0.3 0.3 13.5 0.3 0.6 14.0 0.8 0.7 12.1 0.8 1.9 11.1 0.7 2.3 13.0 0.8 1.2 12.3 0.8 1.5 10.7 0.9 1.3 14.3 0.6 3.3 13.5 0.4 0.6
-Sprott O 3.5 0.1 0.4 9.2 0.2 0.4 6.7 0.1 0.4 8.1 0.2 0.4 11.6 0.4 0.3 13.2 0.3 0.5 11.8 1.0 1.2 12.7 0.8 2.1 11.4 1.0 2.3 12.6 1.0 3.1 10.4 0.9 0.8 14.0 0.7 3.2 9.6 0.2 0.5
-Sprott P 5.7 0.2 0.4 9.2 0.6 0.9 10.5 0.4 0.6 11.1 0.8 0.9 13.3 0.4 0.5 13.5 0.6 1.7 9.6 0.8 6.8 7.6 0.6 5.7 8.7 0.8 4.7 9.6 0.8 5.8 10.8 0.8 1.4 10.4 0.5 3.2 9.1 0.5 1.9
-Sprott Q 5.8 0.2 0.8 9.5 0.5 0.5 9.4 0.5 0.5 10.0 0.7 0.4 12.9 0.5 1.3 12.6 0.3 1.5 11.1 0.9 1.7 13.0 0.7 2.2 12.5 0.7 0.7 11.5 0.8 3.8 9.7 0.9 1.1 13.2 0.4 6.0 12.3 0.4 0.8
-Sprott R 5.1 0.3 0.7 6.9 0.4 0.3 6.7 0.3 0.2 6.5 0.2 0.1 12.5 0.3 0.7 14.2 0.5 1.4 11.0 0.9 2.8 8.8 0.7 2.2 10.6 0.9 3.2 9.4 0.9 1.9 11.4 0.9 2.2 13.8 0.6 6.0 8.9 0.3 1.6
-Sprott S 2.9 0.0 0.3 9.6 0.4 0.4 10.7 0.2 0.3 11.7 0.3 0.5 13.0 0.4 0.2 13.3 0.4 0.6 10.2 0.9 0.8 10.2 0.8 3.4 11.2 0.9 3.1 11.0 0.8 3.5 10.7 0.9 2.2 12.8 0.5 6.0 11.3 0.4 1.9
-Sprott Torus 3.6 0.2 0.1 6.8 0.6 0.2 12.0 0.5 0.2 13.0 0.5 0.3 14.3 0.7 0.3 14.7 0.6 0.8 7.0 0.9 1.1 12.1 0.4 1.1 7.9 0.9 0.8 12.5 0.7 0.8 12.2 0.9 1.1 14.9 0.5 1.5 12.8 0.4 0.9
-Thomas 2.3 0.2 3.1 12.4 0.6 2.9 12.1 0.3 3.3 13.4 0.2 2.5 13.4 0.3 2.9 13.8 0.2 2.7 12.3 0.8 4.4 12.0 0.8 5.2 13.7 0.7 4.2 13.1 0.7 4.9 11.9 0.8 4.0 13.6 0.4 4.6 13.7 0.3 4.3
-Thomas Labyrinth 6.5 0.1 3.2 12.5 0.2 3.2 9.8 0.4 3.3 11.8 0.3 2.5 12.2 0.3 3.3 12.5 0.4 2.8 11.5 0.6 4.1 12.3 0.6 3.8 11.5 0.6 3.6 11.5 0.6 3.6 12.5 0.6 2.5 12.5 0.3 2.3 12.6 0.3 2.6
-Torus 10.9 0.9 0.9 12.4 0.4 0.0 11.9 0.2 0.0 12.0 0.2 0.0 12.9 0.5 0.1 12.9 0.7 0.5 12.0 1.0 0.5 12.6 0.8 0.7 12.6 1.0 0.6 12.4 1.0 0.9 8.7 0.9 0.3 13.1 0.6 1.3 12.5 0.3 0.4
-Tsucs 2 7.2 0.6 1.5 10.9 0.7 1.9 10.3 0.8 1.4 10.0 0.2 2.0 12.0 0.6 2.4 13.3 0.7 2.6 8.5 0.8 2.8 12.3 0.8 3.9 7.5 0.7 3.4 9.4 0.7 4.0 11.5 0.8 3.3 9.8 0.5 4.2 13.6 0.4 3.7
-Vallise El Nino 1.8 0.2 0.2 10.1 0.4 0.2 5.5 0.2 0.2 7.4 0.3 0.2 12.7 0.5 0.3 13.4 0.6 1.2 11.6 0.8 1.3 12.5 0.8 1.3 13.6 0.7 1.2 12.4 0.8 1.3 13.6 0.7 1.0 13.7 0.6 1.6 13.2 0.5 1.9
-Wang Sun 5.3 0.2 0.5 11.8 0.5 1.7 10.3 0.4 0.7 13.2 0.5 1.7 10.4 0.5 0.4 10.5 0.7 0.9 9.3 0.7 1.7 11.9 0.7 1.9 11.2 0.7 2.2 9.9 0.6 2.5 13.3 0.7 2.3 12.6 0.5 1.4 10.5 0.4 1.5
-Winsmi Reduced 5.0 0.3 1.8 8.9 0.4 3.4 11.2 0.4 3.9 13.6 0.4 3.4 13.6 0.5 3.7 14.0 0.7 4.0 11.7 0.9 2.7 11.7 0.8 4.4 11.9 0.9 3.0 10.4 0.8 5.3 11.4 0.9 2.1 14.3 0.6 4.5 13.8 0.3 2.4
-Yu Wang 6.4 0.3 0.3 9.8 0.7 0.5 11.3 0.6 0.6 10.3 0.6 0.7 8.1 0.6 0.6 12.8 0.6 0.6 12.3 0.9 2.9 12.0 0.8 1.8 12.8 0.7 2.1 12.8 0.8 2.2 12.0 0.7 2.7 13.7 0.5 1.8 13.4 0.4 1.7
-Yu Wang 2 3.4 0.1 0.2 10.0 0.2 0.2 10.6 0.4 0.4 12.7 0.2 0.1 12.2 0.2 0.7 13.8 0.4 0.6 10.8 0.9 2.6 9.4 0.8 5.0 8.6 0.7 4.0 10.0 0.7 4.8 11.6 0.9 1.9 10.7 0.3 1.7 13.5 0.3 2.0
-Zhou Chen 10.0 0.4 0.8 10.4 0.4 2.6 13.7 0.5 4.9 12.7 0.4 1.2 12.8 0.3 3.1 12.8 0.4 2.1 13.2 0.5 4.1 11.6 0.5 4.4 13.5 0.5 4.3 9.7 0.5 5.4 12.8 0.6 5.1 11.6 0.4 7.0 14.3 0.5 5.4
+Atmospheric Regime 8.4 0.6 2.3 9.7 0.5 0.9 11.1 0.4 0.5 10.5 0.6 0.7 10.1 0.5 1.0 11.9 0.9 2.2 12.0 0.9 1.9 13.5 0.9 1.8 10.2 0.3 0.8 13.2 0.3 1.0 11.3 0.9 1.3 14.2 0.6 3.1 13.9 0.4 1.2
+BeerRNN 11.7 0.8 1.2 8.0 0.5 0.2 3.3 0.2 0.2 9.5 0.2 0.2 7.3 0.4 0.2 12.3 0.9 1.2 12.5 0.9 1.2 12.1 0.9 1.2 12.6 0.4 1.3 13.9 0.3 0.5 13.0 0.9 1.4 14.1 0.6 1.3 14.0 0.4 0.8
+Bouali 4.1 0.1 0.5 8.3 0.5 0.1 7.3 0.3 0.2 8.6 0.1 0.2 11.1 0.3 0.5 10.6 0.8 1.0 9.6 0.7 1.1 12.2 0.7 2.0 7.6 0.4 0.5 13.1 0.4 0.5 12.9 0.8 2.3 13.4 0.4 2.3 12.7 0.3 0.7
+Burke Shaw 4.1 0.2 0.5 10.8 0.5 1.2 9.0 0.5 1.5 13.0 0.4 1.1 13.2 0.4 1.5 11.5 0.8 1.3 12.1 0.6 2.4 12.2 0.7 2.2 11.1 0.4 1.0 13.4 0.4 0.4 12.2 0.9 3.7 13.7 0.5 3.6 13.9 0.4 1.5
+Ca Two Plus Quasip. 2.3 0.1 0.7 7.2 0.1 0.2 8.3 0.1 0.3 9.2 0.4 0.3 13.1 0.4 0.2 11.3 0.9 1.0 11.6 0.8 1.1 12.8 0.6 1.2 7.0 0.2 0.5 12.3 0.2 0.3 11.1 0.8 3.0 13.8 0.7 2.6 13.4 0.3 1.9
+Chen Lee 3.9 0.4 0.3 4.9 0.3 0.3 4.0 0.5 0.5 4.8 0.3 0.1 4.6 0.3 0.3 7.2 0.9 1.6 7.9 0.7 1.4 8.0 0.7 2.0 4.9 0.3 0.2 6.2 0.3 0.3 11.0 0.9 0.9 6.4 0.5 3.4 7.3 0.4 1.8
+Chua 5.9 0.1 0.3 7.4 0.2 0.6 5.9 0.2 0.7 7.6 0.1 0.7 11.7 0.2 0.4 11.8 0.6 1.9 12.2 0.5 2.2 11.7 0.6 1.7 9.4 0.2 0.5 12.2 0.4 0.6 11.3 0.5 3.2 13.4 0.4 2.7 10.4 0.2 0.7
+Coullet 1.2 0.0 0.2 4.9 0.0 0.3 0.7 0.0 0.1 2.8 0.0 0.1 7.5 0.2 0.1 13.0 0.8 2.8 11.2 0.8 3.8 11.8 0.8 3.4 9.8 0.2 0.2 9.7 0.3 0.2 12.3 0.8 2.4 14.9 0.5 3.3 10.7 0.2 0.5
+Dadras 4.1 0.2 1.1 7.7 0.6 0.5 4.7 0.5 0.5 13.1 0.5 0.3 12.1 0.4 0.2 12.7 0.8 1.7 12.9 0.7 2.0 13.0 0.8 2.2 9.2 0.5 0.6 12.0 0.4 0.6 11.9 0.8 1.7 12.8 0.4 2.9 11.8 0.4 1.2
+Dequan Li 12.7 0.5 2.1 11.1 0.3 1.1 11.0 0.4 0.7 12.1 0.3 0.9 12.1 0.2 1.9 11.7 0.4 3.1 11.3 0.4 2.8 9.6 0.4 2.4 11.5 0.4 1.3 13.2 0.2 0.9 10.7 0.4 3.1 13.5 0.4 4.5 12.8 0.3 2.0
+Finance 3.2 0.1 0.6 8.5 0.6 1.6 8.6 0.4 0.9 9.1 0.4 0.4 12.5 0.3 0.4 11.6 0.9 1.8 10.7 0.8 1.7 13.4 0.8 2.0 11.0 0.3 1.0 13.3 0.3 0.8 11.1 0.9 2.3 13.7 0.5 4.2 13.6 0.4 2.6
+Genesio Tesi 8.2 0.2 0.9 7.9 0.4 1.0 8.5 0.2 1.1 7.7 0.2 1.0 9.4 0.2 1.0 10.4 0.7 1.4 10.6 0.7 1.1 10.3 0.8 2.2 7.4 0.4 1.2 12.3 0.4 0.4 11.0 0.9 1.4 12.6 0.6 5.4 10.6 0.4 1.0
+Guckenheimer Holmes 9.6 0.1 0.5 11.8 0.1 0.8 12.0 0.1 0.4 12.6 0.3 0.5 12.5 0.4 1.1 9.6 0.7 0.6 9.6 0.7 2.5 9.4 0.6 1.0 8.2 0.1 0.5 10.8 0.2 0.6 10.2 0.7 1.0 12.7 0.5 3.5 11.5 0.3 0.9
+Hadley 11.3 0.4 1.3 10.7 0.5 2.3 9.7 0.5 1.8 12.5 0.3 1.5 12.6 0.6 1.8 11.4 0.8 5.1 11.9 0.7 4.7 11.9 0.6 3.9 10.9 0.2 4.7 12.6 0.3 0.8 11.4 0.8 6.3 12.5 0.5 7.1 12.4 0.5 3.8
+Halvorsen 4.3 0.3 0.6 9.7 0.6 0.4 8.2 0.2 0.7 7.5 0.2 0.7 13.1 0.2 0.6 10.0 0.8 5.0 10.3 0.7 4.3 9.7 0.8 4.6 11.8 0.3 2.3 13.1 0.2 0.3 11.3 0.9 3.4 13.4 0.5 4.2 13.3 0.3 1.6
+Hindmarsh Rose 6.5 0.4 0.7 9.8 0.2 0.2 4.9 0.1 0.1 9.6 0.2 0.1 12.7 0.6 0.3 12.9 0.7 1.6 14.3 0.8 2.0 12.9 0.8 1.7 8.1 0.3 0.6 9.6 0.3 0.9 13.2 0.8 2.5 14.6 0.4 1.4 14.3 0.5 1.1
+Isothermal Chemical 2.0 0.2 0.2 2.5 0.1 0.6 5.4 0.1 0.3 8.3 0.2 0.2 7.0 0.4 0.2 11.7 0.9 1.1 11.7 0.9 1.5 10.8 0.8 1.6 6.5 0.3 0.4 5.8 0.3 0.2 10.1 0.9 0.7 15.6 0.6 3.5 7.5 0.3 1.3
+Itik Banks tumor 5.7 0.2 0.5 8.2 0.5 0.5 10.8 0.6 0.4 9.8 0.5 0.3 11.8 0.6 0.4 10.8 0.6 3.4 11.2 0.8 3.8 11.9 0.9 2.9 9.8 0.4 0.3 12.5 0.4 0.5 12.0 0.9 2.9 12.9 0.7 4.1 13.4 0.3 1.1
+Jerk Circuit 4.3 0.3 0.8 8.8 0.6 0.4 9.1 0.4 0.3 10.6 0.3 0.2 11.7 0.3 0.2 12.5 0.9 1.0 13.5 0.9 0.8 12.9 0.9 0.8 10.6 0.5 0.7 13.5 0.4 0.4 10.3 0.8 0.6 14.1 0.5 1.9 13.7 0.4 1.1
+Laser 4.5 0.2 0.2 7.5 0.5 3.0 8.2 0.3 1.2 11.5 0.2 0.5 12.8 0.2 0.5 12.6 0.7 2.0 12.8 0.5 2.8 13.0 0.7 1.7 7.4 0.4 4.8 13.1 0.3 0.6 11.8 0.8 2.5 13.3 0.5 2.6 13.8 0.4 1.0
+Lorenz 3.4 0.2 0.5 9.8 0.6 1.5 8.5 0.3 1.6 13.0 0.4 0.7 12.6 0.7 2.0 12.8 0.8 2.6 13.0 0.7 3.4 13.2 0.6 3.0 11.8 0.4 0.8 12.7 0.3 0.6 11.8 0.8 2.7 13.4 0.5 3.1 13.1 0.3 2.1
+Lorenz 84 7.9 0.2 0.5 10.7 0.5 1.5 11.0 0.4 1.3 12.6 0.4 1.0 11.0 0.4 0.4 11.8 0.9 1.3 12.2 0.9 1.4 11.5 0.9 1.3 12.8 0.2 0.4 13.4 0.3 0.4 11.2 0.9 1.2 13.4 0.6 1.2 13.4 0.5 1.4
+Lu Chen 9.5 0.8 0.7 5.3 0.2 0.5 9.0 0.3 0.7 8.9 0.2 0.2 9.7 0.3 0.5 12.2 1.0 1.7 13.8 0.9 1.8 12.2 0.9 1.6 7.0 0.2 0.6 11.7 0.3 0.7 10.1 0.8 1.1 14.2 0.5 1.9 7.4 0.5 0.9
+Moore Spiegel 9.2 0.6 0.4 9.1 0.6 0.4 10.9 0.7 0.4 12.5 0.5 0.3 11.5 0.6 0.3 11.4 0.9 2.0 11.7 0.8 2.7 12.7 0.9 2.0 11.4 0.5 1.2 13.2 0.3 0.5 10.5 0.9 2.2 13.1 0.7 1.5 12.6 0.5 2.1
+Multi Chua 12.5 0.5 0.4 12.0 0.6 0.8 11.9 0.6 0.4 13.4 0.4 1.0 13.0 0.4 0.8 12.5 0.5 1.5 12.6 0.5 1.1 13.6 0.5 1.1 12.9 0.2 0.3 13.0 0.2 0.6 10.7 0.5 0.7 13.5 0.5 2.1 12.7 0.3 1.9
+Nose Hoover 11.4 0.5 0.9 10.2 0.6 0.5 10.8 0.7 0.5 10.8 0.6 0.5 12.9 0.8 0.6 11.2 0.9 2.0 11.3 0.9 1.4 12.7 0.7 2.0 13.1 0.4 0.8 12.9 0.3 0.7 10.5 0.9 2.9 13.3 0.5 3.3 12.8 0.4 0.5
+Pehlivan Wei 4.8 0.2 1.9 4.6 0.1 0.7 6.8 0.4 0.9 9.0 0.3 1.0 10.7 0.3 1.3 11.1 0.9 5.6 12.0 0.6 6.8 12.6 0.8 6.5 8.1 0.2 2.2 8.3 0.4 0.8 10.0 0.9 4.1 12.7 0.4 7.7 9.1 0.3 3.3
+Sakarya 3.1 0.3 0.4 2.6 0.2 0.2 5.0 0.3 0.3 8.2 0.3 0.3 10.2 0.7 0.6 10.7 0.9 0.5 12.7 0.9 2.5 11.8 0.8 0.9 10.8 0.5 0.5 11.5 0.3 0.3 12.4 0.7 0.8 14.0 0.5 2.5 13.8 0.5 1.5
+San Um Srisuchinwong 5.5 0.1 0.6 5.5 0.2 1.2 10.5 0.8 1.3 12.6 0.4 1.2 13.0 0.5 0.8 11.8 0.8 1.1 13.0 0.7 0.5 12.0 0.6 1.2 10.3 0.2 2.2 13.0 0.3 0.2 10.7 0.5 1.8 12.6 0.5 0.9 12.6 0.4 1.4
+Sprott A 4.9 0.1 1.0 9.3 0.5 0.8 8.5 0.4 0.8 7.5 0.3 0.4 11.7 0.3 0.9 10.7 0.9 4.1 11.8 0.7 4.5 11.5 0.8 4.4 10.4 0.2 1.1 11.5 0.4 0.8 9.8 0.9 2.1 12.4 0.6 6.0 11.2 0.2 1.1
+Sprott B 2.6 0.2 0.9 9.2 0.5 1.7 11.0 0.6 1.8 13.2 0.5 2.5 13.4 0.6 1.5 12.6 0.8 2.7 13.2 0.8 2.7 12.5 0.8 2.2 13.7 0.7 2.7 12.9 0.3 0.6 13.9 0.8 3.9 12.9 0.6 2.9 13.7 0.3 2.5
+Sprott C 2.9 0.2 0.4 8.5 0.6 1.1 12.0 0.3 0.5 13.3 0.4 1.1 12.6 0.4 0.8 12.4 0.7 2.0 13.0 0.6 2.2 11.9 0.6 2.5 11.6 0.4 0.7 12.0 0.4 0.5 13.3 0.7 2.6 13.5 0.5 2.9 13.5 0.4 1.9
+Sprott D 3.8 0.1 0.6 7.6 0.2 0.2 5.5 0.1 0.2 5.5 0.2 0.1 12.9 0.4 0.7 12.1 0.7 0.8 12.0 0.8 4.6 12.1 0.7 3.1 10.4 0.3 1.2 12.0 0.4 0.3 12.9 0.7 1.9 13.8 0.5 2.3 13.2 0.5 0.6
+Sprott F 4.9 0.3 0.5 10.8 0.4 0.3 8.9 0.5 0.5 12.6 0.3 0.3 13.0 0.5 0.2 12.3 0.9 1.2 12.2 0.6 1.3 13.2 0.8 1.2 13.2 0.5 0.5 13.4 0.3 0.6 11.8 0.9 1.3 13.4 0.6 1.8 13.2 0.5 1.2
+Sprott G 2.4 0.0 0.3 4.2 0.1 0.1 6.2 0.2 0.1 7.1 0.2 0.2 7.7 0.3 0.5 12.3 0.8 3.6 12.4 0.7 4.0 12.9 0.8 3.4 6.6 0.2 0.6 10.5 0.2 0.3 11.0 0.9 2.6 14.0 0.5 2.0 10.6 0.3 1.2
+Sprott I 3.0 0.2 0.3 8.6 0.5 0.1 10.8 0.1 0.2 7.4 0.3 0.3 10.1 0.3 0.3 12.1 0.9 1.2 11.9 0.8 2.5 9.4 0.9 1.0 8.6 0.3 0.2 8.2 0.2 0.1 10.9 0.9 0.9 14.4 0.6 3.2 9.1 0.4 0.3
+Sprott L 2.7 0.2 0.2 9.2 0.3 0.2 8.2 0.1 0.2 10.2 0.2 0.1 11.3 0.2 0.1 12.3 0.9 2.6 12.3 0.9 2.7 13.3 0.9 1.8 6.7 0.2 0.2 12.0 0.2 0.2 12.2 0.8 0.6 14.2 0.5 2.5 11.7 0.3 0.2
+Sprott M 3.4 0.2 0.4 10.9 0.5 0.1 12.4 0.4 0.2 12.8 0.3 0.3 13.5 0.3 0.6 12.1 0.8 1.9 11.1 0.7 2.3 13.0 0.8 1.2 7.6 0.2 0.7 13.7 0.2 0.4 10.7 0.9 1.3 14.3 0.6 3.3 13.5 0.4 0.6
+Sprott O 3.5 0.1 0.4 9.2 0.2 0.4 6.7 0.1 0.4 8.1 0.2 0.4 11.6 0.4 0.3 11.8 1.0 1.2 12.7 0.8 2.1 11.4 1.0 2.3 9.5 0.3 0.3 10.9 0.1 0.2 10.4 0.9 0.8 14.0 0.7 3.2 9.6 0.2 0.5
+Sprott P 5.7 0.2 0.4 9.2 0.6 0.9 10.5 0.4 0.6 11.1 0.8 0.9 13.3 0.4 0.5 9.6 0.8 6.8 7.6 0.6 5.7 8.7 0.8 4.7 7.6 0.2 0.9 11.4 0.3 0.4 10.8 0.8 1.4 10.4 0.5 3.2 9.1 0.5 1.9
+Sprott Q 5.8 0.2 0.8 9.5 0.5 0.5 9.4 0.5 0.5 10.0 0.7 0.4 12.9 0.5 1.3 11.1 0.9 1.7 13.0 0.7 2.2 12.5 0.7 0.7 11.7 0.4 0.8 13.0 0.4 0.4 9.7 0.9 1.1 13.2 0.4 6.0 12.3 0.4 0.8
+Sprott R 5.1 0.3 0.7 6.9 0.4 0.3 6.7 0.3 0.2 6.5 0.2 0.1 12.5 0.3 0.7 11.0 0.9 2.8 8.8 0.7 2.2 10.6 0.9 3.2 7.4 0.3 1.3 9.8 0.2 0.4 11.4 0.9 2.2 13.8 0.6 6.0 8.9 0.3 1.6
+Sprott S 2.9 0.0 0.3 9.6 0.4 0.4 10.7 0.2 0.3 11.7 0.3 0.5 13.0 0.4 0.2 10.2 0.9 0.8 10.2 0.8 3.4 11.2 0.9 3.1 11.9 0.3 0.7 12.9 0.5 0.4 10.7 0.9 2.2 12.8 0.5 6.0 11.3 0.4 1.9
+Sprott Torus 3.6 0.2 0.1 6.8 0.6 0.2 12.0 0.5 0.2 13.0 0.5 0.3 14.3 0.7 0.3 7.0 0.9 1.1 12.1 0.4 1.1 7.9 0.9 0.8 10.7 0.4 0.6 8.5 0.3 0.2 12.2 0.9 1.1 14.9 0.5 1.5 12.8 0.4 0.9
+Thomas 2.3 0.2 3.1 12.4 0.6 2.9 12.1 0.3 3.3 13.4 0.2 2.5 13.4 0.3 2.9 12.3 0.8 4.4 12.0 0.8 5.2 13.7 0.7 4.2 12.9 0.3 4.3 13.6 0.3 0.9 11.9 0.8 4.0 13.6 0.4 4.6 13.7 0.3 4.3
+Thomas Labyrinth 6.5 0.1 3.2 12.5 0.2 3.2 9.8 0.4 3.3 11.8 0.3 2.5 12.2 0.3 3.3 11.5 0.6 4.1 12.3 0.6 3.8 11.5 0.6 3.6 11.7 0.3 2.8 12.6 0.4 3.3 12.5 0.6 2.5 12.5 0.3 2.3 12.6 0.3 2.6
+Torus 10.9 0.9 0.9 12.4 0.4 0.0 11.9 0.2 0.0 12.0 0.2 0.0 12.9 0.5 0.1 12.0 1.0 0.5 12.6 0.8 0.7 12.6 1.0 0.6 11.5 0.4 0.2 12.1 0.2 0.4 8.7 0.9 0.3 13.1 0.6 1.3 12.5 0.3 0.4
+Tsucs 2 7.2 0.6 1.5 10.9 0.7 1.9 10.3 0.8 1.4 10.0 0.2 2.0 12.0 0.6 2.4 8.5 0.8 2.8 12.3 0.8 3.9 7.5 0.7 3.4 7.2 0.4 1.0 6.2 0.2 1.1 11.5 0.8 3.3 9.8 0.5 4.2 13.6 0.4 3.7
+Vallise El Nino 1.8 0.2 0.2 10.1 0.4 0.2 5.5 0.2 0.2 7.4 0.3 0.2 12.7 0.5 0.3 11.6 0.8 1.3 12.5 0.8 1.3 13.6 0.7 1.2 9.5 0.2 1.5 13.2 0.4 0.3 13.6 0.7 1.0 13.7 0.6 1.6 13.2 0.5 1.9
+Wang Sun 5.3 0.2 0.5 11.8 0.5 1.7 10.3 0.4 0.7 13.2 0.5 1.7 10.4 0.5 0.4 9.3 0.7 1.7 11.9 0.7 1.9 11.2 0.7 2.2 10.8 0.5 0.4 12.6 0.3 0.5 13.3 0.7 2.3 12.6 0.5 1.4 10.5 0.4 1.5
+Winsmi Reduced 5.0 0.3 1.8 8.9 0.4 3.4 11.2 0.4 3.9 13.6 0.4 3.4 13.6 0.5 3.7 11.7 0.9 2.7 11.7 0.8 4.4 11.9 0.9 3.0 6.3 0.2 4.1 13.1 0.1 1.9 11.4 0.9 2.1 14.3 0.6 4.5 13.8 0.3 2.4
+Yu Wang 6.4 0.3 0.3 9.8 0.7 0.5 11.3 0.6 0.6 10.3 0.6 0.7 8.1 0.6 0.6 12.3 0.9 2.9 12.0 0.8 1.8 12.8 0.7 2.1 12.8 0.7 1.3 13.4 0.3 0.3 12.0 0.7 2.7 13.7 0.5 1.8 13.4 0.4 1.7
+Yu Wang 2 3.4 0.1 0.2 10.0 0.2 0.2 10.6 0.4 0.4 12.7 0.2 0.1 12.2 0.2 0.7 10.8 0.9 2.6 9.4 0.8 5.0 8.6 0.7 4.0 9.6 0.4 0.9 10.7 0.4 0.4 11.6 0.9 1.9 10.7 0.3 1.7 13.5 0.3 2.0
+Zhou Chen 10.0 0.4 0.8 10.4 0.4 2.6 13.7 0.5 4.9 12.7 0.4 1.2 12.8 0.3 3.1 13.2 0.5 4.1 11.6 0.5 4.4 13.5 0.5 4.3 13.0 0.3 2.3 11.8 0.4 0.7 12.8 0.6 5.1 11.6 0.4 7.0 14.3 0.5 5.4